by TEI TEI

DDS-Beckman Adult Center Change of Use Remediation Project

5/17/10- GBBN received final Structural Analysis for Beckman Change of Use Code Review.

5/25/10- Arcodect delivered final Code Review for implementation into Permit Set of Construction Documents.

5/26/10- Final Permit Set of Construction Documents printed, signed and stamped. Set to be hand delivered to TEI at the Partnering Meeting for issuance to the Hamilton County Planning Department.

6/4/10- Issuance of ITB#047-10 by Hamilton County Purchasing Department.

6/14/10- Attend Pre-bid Meeting at Beckman Adult Center.

6/15/10- Issuance of Addenda #1 with Pre-bid meeting minutes.

6/23/10- Bid question sent directly to GBBN. Bidder instructed to send question to Purchasing per the bidding documents.

7/1/10- Building Department Response Letter received from TEI with two items concerning occupany separations. Building Department requesting drawing showing Use Group separations, net/gross areas and occupancy locations on floor plan.

7/02/10- Responses for Building Department Response Letter received from Wayne Meyer at Arcodect.

7/19/10- Newly created Occupancy Summary Drawings for Building Department Response Letter delivered to TEI.

9/15/10- Construction meeting was held.

9/27/10- TEI announced the Plans Examination was completed and approved for Building Permit by Hamilton County.

1/27/11- TEI and Kevin Maue confirmed the permit is now closed on this project. The Certificate of Occupancy and the Colerain Township Fire Department Letter of Release have been obtained.

1/27/11- Kevin Maue indicated at the Partnering meeting this project can be archived.

by TEI TEI

DDS-Beckman Adult Center Change of Use Code Investigation

3/19/10-Met with Jim Bowman, Colerain Fire Department, and Kevin Maue at Beckman Adult Center for walk-through on current renovation project.

3/20/10-Kevin Maue requested SOSS be submitted. This is first of two SOSS’s for this overall project.

3/21/10-Preliminary Code Investigation performed and submitted to Kevin Maue by email.

3/23/01-Fee proposal from Arcodect (code consultant) received. GBBN spoke to Cindy Rutherford, Hamilton County Building Department to discuss Change of Use requirements and submissions.

3/31/10-Schedule of submissions to Hamilton County Building Department submitted by email to Kevin Maue.

4/1/10-SOSS to be revised and re-submitted showing THP as a structural consultant to GBBN for any seismic implications per directions of Kevin Maue at Partnering Meeting.

4/7/10- GBBN met with Kevin maue, Terry Cannon (TEI) and Tom Frimming of Communale Fire Protection to review proposed AFSS implementation.

4/14/10- Signed SOSS without Ralph Linne’s signature and approval delivered by Kevin Maue to GBBN.

4/27/10- Signed SOSS with Ralph Linne’s signature and approval delivered by Bert Watts via email to GBBN.

4/27/10- Bryan Korb of GBBN assigned to project. Bryan to meet with Wayne Meyer code consultant at Beckman on Thursday and Friday for site investigation

4/29/10- GBBN to meet with owner, Colerain Fire Department, and Hamilton County Building Department to review building and change of use scope.

5/12/10- Kevin Maue announced ThermalTech Engineering would be the lead design professional on this project and GBBN would be the consultant. GBBN established estimated hours for completion to TEI for SOSS.

by TEI TEI

DDS Signage Refacing

10/01/09 – Kevin Maue of MR/DD contacted GBBN about the pending name change as mandated by the State of Ohio. Sign letters on 1520 Madison to be removed as part of ceremony and will be handled as a purchase order. The signage refacing of all other sites will be on a SOSS through the Partnership contract.

10/05/09- ASI Modulex was approved to handle the purchase order work for 1520 Madison based on a submitted quote to MR/DD.

10/07/09-Existing electronic drawings sent to ASI for signage permit with City of Cincinnati.

10/23/09- Shop drawings for 1520 Madison approved for layout and lettering style.

10/30/09- Signage Refacing RFP sent to four selected bidders.

11/24/09- Adams Signs was selected for Signage Refacing and P.O. 536683 released.

12/08/09-1520 Madison sign permit approved. ASI
to order building sign letters.

1/04/10-Adams Signs re-submittal shop drawings approved.

2/19/10-Adams Signs has completed all installations. Kemper Meadow Road sign has bleed through of old signage.

2/22/10-Email sent to Rick Cain of Adams Signs about corrective measures at Kemper Meadow Road location.

2/25/10-Received signed SOSS from Bert Watts at monthly partnering meeting.

3/15/10-Resent email to Rick Cain with photos of Kemper Meadow. Rick claimed the corrections would be completed soon.

4/3/10- Email sent to Rick Cain of Adams Signs about corrective measures at Kemper Meadow Road location.

4/10/10- Rick Cain called to indicate Adams Signs very busy but will try and expedite completion of remediation measures to Meadow Road location.

5/19/10- Kevin Maue sent an email to Adams Signs indicating Application for Payment non-payment until corrective work is accomplished at Meadow Road location.

6/18/10- Notified Adams Signs to submit warranty information to comply with close-out documentation.

6/23/10- Sent Invoice from Adams Signs to Kevin Maue.

6/24/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

6/29/10- Rick Cain called to indicated work is commencing to author standard warranty details.

7/28/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

7/30/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

8/24/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

9/27/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

10/25/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

11/25/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment.

12/14/10- Contacted Rick Cain (copied Kristal Dadisman) by email to comply again with the warranty requirements for closeout of contract and release payment. Rick Cains email was returned as undeliverable.

by TEI TEI

MRDD Franks Center Renovations

04/15/09 – CR Architecture + Design awaiting final punchlist items to be completed and closeout documents.

01/30/09 – Substantial completion issued.

01/30/09 – Final punchlist completed.

10/29/08 – Phase I punchlist.

09/22/08 – Additional service issued for Kitchen Counters.

by TEI TEI

MRDD – THOMAS KIDD CLASSROOM ADDITION

02/14/08 – Punch list submitted.

02/13/08 – CRA sent additional service signed 2/13/08 to Bert

02/13/08 – TEI attended the const mtg today for MRDD. EDI is almost complete. Additional concrete will be installed around the new air handlers.

12/07/07 – CRA job meeting held on 12/05/07

11/14/07 – CRA received and approved shop drawings and substitution request for automatic door operators.

11/10/07 – The construction has commenced.

11/01/07 – Pre-construction meeting was held. See Meeting Minutes below.

10/26/07 – GC/EC submitted revised insurance certificate. GC/EC contracts have been approved approved. HVAC contract is still pending.

09/28/07 – Awaiting GC/EC to submit revised insurance certificate for final executed Contract.

08/16/07 – Contracts submitted 07/31, CRA received comments, re-issued 08/16.

08/02/07 – C+RA and TE submitted the Building Permit Application to the Village of Evendale Building Department.

06/25/07 – Addendum issued. Bids are due 07/02/07.

06/15/07 – A pre-bid meeting was held on location yesterday, June 14, 2007. This project is set to open for bid on June 26, 2007.

06/08/07 – This project has been sent to Cole & Russell for review. TE is awaiting comments back.

04/27/04 – Final review meeting is scheduled for next week.

04/20/07 – Updated drawings are being sent out today, April 20 2007.

04/13/07 – Project is progressing as expected.

04/06/07 – Preliminary set of drawings were sent to Architect today, April 6, 2007. Awaiting review comments.

03/30/07 – Final field work has been assessed. TE working on design.

03/23/07 – Project is progressing as expected.

03/20/07 – Kickoff meeting complete. Check CRA for updates.

by TEI TEI

MRDD – JACKSON CENTER RENOVATION

02/15/08 – Add service CR-077a for $6,755.00 signed by MRDD on 2/12/08, forwarded to Bert Watts.

01/16/08 – Mtg held re additional service request my MRDD.

01/17/08 – Meeting held at site for scope definition. TJ and MAH attended. A few deletions and additions from original scope. ZZ ran meeting and will produce meeting notes. TJ and MAH will produce notes and forward to ZZ.

12/20/07 – Add service requested by MT

12/07/07 – CRA Told by MT that project will be from 2008 budget.

10/10/07 – Received signed SOSS.

09/06/07 – SOSS submitted.

05/18/07 – Kickoff meeting was held.

by TEI TEI

MRDD – Franks Center Renovations

03/31/08 – Bids due 4/11/08

03/31/08 – Pre-bid meeting

03/21/08 – Addendum 1 issued

03/08/08 – Issued for bid

02/15/08 – CRA Add Service CR-084a $6,999.00 signed by MRDD on 2/12/08, forwarded to Bert Watts.

01/11/08 – SOCC CR-084a modified per WM email 1/11/08

12/20/07 – CRA add service requested by MT 12/20/07

11/09/07 – SD review meeting was held at Franks Adult Center. CRA presented Schematic Design drawings and color and finishes samples.

11/02/07 – Next review meeting is scheduled for 11/08/07 at Franks Adult Center at 3pm.

10/24/07 – Meeting was held at Franks Adult Center to discuss the scope of the project and verify existing conditions.

09/24/07 – CRA received signed and approved SOSS.

08/30/07 – CRA submitted the SOSS.

by TEI TEI

MRDD – Storage Buildings

01/02/08 – Final warranty info received 1/2/08, forwarded to MT

11/26/07 – Received As-Built drawings, forwarded to MT on 11/26/07.

10/01/07 – Awaiting closeout documents. Need As-Built drawings.

08/15/07 – CO #1 and Pay Application #3 sent to Michael Trabue.

07/27/07 – Construction is complete. Pay Application #2 sent to Michael Trabue.

by TEI TEI

MRDD – 1520 MADISON RD RENOVATION

08/29/08 – Triton is still working on warranty issues. Boiler control boards. Humidity issues. RCP problems.

02/25/08 – Triton and JCI have been working to remedy the remaining warranty issues on this project which include boiler flame outs, VAV box control, min airflows through VAV boxes, winter heating. TEI has to look at cold restrooms in the 1956 building for Kevin Maue.

12/17/07 – JCI and Triton had to go back to the site today to continue to work on DDC issues since the RTU-2 is still not running properly. Triton put the unit in had on Friday at 40%. TEI does not have an update on Monday from the troubleshooting field trip.

12/14/07 – TTC was on site today with JCI and Carrier going over warranty issues on the job including the LON talk interface between RTU-2 and DDC. Scott fixed the misnomers on the RTU-1 unit in the points list. Scott fixed the economizer program so RTU-1 would actually run in economizer mode when conditions were right (there was a programming glitch in the system) and he addressed the setback temperatures which are correct as programmed. Training will be schedule the second week of January 2008 with MRDD. There was a cold call in the lobby today from Wayne Mckee and Triton and JCI was dispatched to investigate the problems with the space. The hot water is 156F so there should be plenty of hot water available for the system.

12/04/07 – CDFP is scheduled the FM200 training directly with the MRDD folks and they will keep TTC in the loop. CDFP sent a change order in today for additional services on this project that TTC forwarded to ZZ for approval on this project for the extra sprinkler work, additional fan test.

11/28/07 – Jim Lape emailed to ask about closeout on this job. TEI stated that TJ should be 100% complete on this project. Triton is also done but we don’t have final closeout documents from them on this job. TTC is trying to get Grady to give us all closeout docs for this job.

10/17/07 – TTC visited the site and produced some more items for JCI and Triton to take care of on this job. He emailed pics and instructions to the contractors for completion. Everyone is working on completing all the work this week. Triton says the final balance reports are complete. TTC said today that he would adjust the manual high limit controllers to 210F and let the factory 190F shut the boilers down on auto reset. TTC also said he would caulk up the hole in the concrete hatch patch while he was on site. TTC will also check the installation of the electric heaters in the stairwell today while he is on site.

10/08/07 – TTC answered Jim Lape’s email about door control hardware. On drawing E400 #12 & #13 state that TJ will wire up the devices furnished & installed by the GC.

10/01/07 – TTC was on site last week performing commissioning. The control system was not ready to be back checked by TEI. JCI had not performed the functional test and TTC did not have the CX sheets. Nevertheless the system is running and with another week of actual work, it will be fine. TTC requested trend reports for verification of working systems.

09/19/07 – TTC approved JMW change order for roof drain and sent email to Jim, Zach and Kevin to that effect. TTC reviewed Triton invoice and asked for change order modifications before actual billing to the County. Change Orders were not reflected on actual payapp but they were on the schedule of values.

09/17/07 – Discussed Change Orders # 09,10,11 with Tim and Kevin Maue. I approved them as submitted. The steel issue is still are large issue that doesn’t appear close to resolution. TTC responded by email about controlling the pressure of the system by setting the static residual to 5 psig and letting the building set the pressure from there. If we have to change the relief valves to 65 psig from there, we can. The static of the building on the basement pumps should be 26.64 psig (50 ft bldg ht / .4328 ft h20/ psi + 5 psig to keep air out).

09/12/07 – Triton is on site checking the existing VAV boxes for repairs. They found one bad damper and believe 6 of the 7 boxes that are not operating properly. We are concerned that we have paid for some of this work already at the beginning of the job, but we will pay for the repairs to the dampers now and fix the electric heaters.

09/04/07 – JCI has given training but the controls are not complete and they still need some work. Boiler are not operational at this time and Denny Lammers is still need to program the internals of the boilers so JCI can communicate and give them their firings rates and monitor alarms. TTC dropped off TJ Williams hard copy as-builts drawings for the project that need to be included into the O&M’s that Kevin already has in his office.

08/23/07 – TEI printed out 2 more originals for Change Order #1, signed them, and sent them to T.J. Williams with instruction to sign them and forward them to HCFD.

08/17/07 – TEI received as-built drawings from TJ. Williams today. TEI will update the electronic drawings with the information and forward them back to CRA early next week. TTC delivered the O&M’s for EC, PC and FPC to Jim Lape at his office today.

08/16/07 – TTC and Cintas performed and passed the final server room fire alarm test on Wednesday. TTC has the paperwork in the folder and will include in the O&M. Triton said that the handrail was ordered and will be put back by Pat Murphy.

08/06/07 – TJ will do the security camera and emergency lighting change order work at T&M not to exceed but they will not come down in price. TTC sent this info to Jim Lape and approved on his end. TEI needs direction from the owner. TTC contacted MRDD about installing gaskets behind wall plates in server room.

08/02/07 – TTC talked to Kevin today about TJ outstanding issues at the MRDD. Security Camera change order and Emergency Lighting change order. JMW has O&M’s into TEI and they all closeout paperwork into Kevin Maue. TEI received TJ Williams closeout paperwork on this project as well. NV had TTC give to it to NR for approval.

07/23/07 – TTC spoke to Tim Myers today about project closeout, change orders, punchlist and controls. TTC discussed the allowance usage with Tim which totaled $13459 to date not including the refrigerant changeout which is a different cost. TTC emailed CRA about distributing the correct paperwork for change order pricing from Triton on this job and listed 6 items that need to be priced for this project.

1. Extra Diffuser for 1956 building. 17 in Qty. Linear Slots and 2×2’s to meet job requirements.
2. Refrigerant Check, Removal, Replacement on RTU-1. This work wasn’t included in the allowance cost but we authorized Triton to do the work on a separate ticket.
3. Overtime for IT closets and occupation over one weekend (i believe KM authorized triton to do this work)
4. Replacement cost for handrail? Triton removed it but someone else threw it in the garbage. Hobert is supposedly looking for it today but Triton will have to replace it if it can’t be found.
5. Demolition of all housekeeping pads in boiler room. This was Empire work but Triton had to do it all because Empire refused to do it. If Triton is going to pay the steel cost as a change order to Empire which I now believe they will after speaking with Tim Myers this morning, Empire either has to reimburse them this cost or deduct from the steel change order cost. Tim brought up a good point. A $2029 redesign fee is steep. That is 20 hours of a ThermalTech Principal time or 40 hours of a contractors time. It did not take that long. Redrawing that steel took about 6 to 8 hours. That fee should from the sub-contractor $500-$700. Please try to get them to give us a “real” cost and the housekeeping pad back charge could go away. As a side note, the new designed decreased the amount of steel and grating required, why are they paying an upcharge of $2522.97?
6. Replacement of 2 electric heaters. Provide 2 recessed ceiling heaters in the 1956 stairwell in lieu of the scheduled heaters. Return the surface mounted fin heater and the wall heater.
07/17/07 – TTC has scheduled a backcheck of the plumbing punchlist on Thursday. He will meet Jason on site for the walk through. Triton has committed to Approved Products being on site to finish the balance portion of the work starting on 07/23. TTC sent 2 sketches to CRA for plumbing ASI’s.

07/03/07 – The job is mostly complete aside from outstanding issues. The MRDD people are in and using the building. Items that we are still working on include: balancing HVAC & water, relocating a fp line out of the elevator shaft, final test of FM200, hot water boiler loop, controls, roof drain near unit, and punchlist by all contractors.

06/26/07 – All TEI punchlist were generated last Friday. TEI gave Jim Lape a complete set of extra materials required on this project to included in substantial completion for the project. TTC was onsite with JMW plumbing early Monday morning to discuss punchlist items which they are addressing. TTC spoke with Kevin about roof drain issue on 1956 building and we agreed that changing the scupper might be the best option. We can drain it down to the canopy roof.

06/15/07 – TE will be generating a punch list next week.

06/07/07 – RTU is still surging. we are trying to find out the status of the jci & carrier contractors. JMW plumbing will be on site on Friday to demo the 4″ plumbing vent from the elevator room. we believe fire protection is complete and ready for testing.

06/04/07 – Inspections are scheduled for Tuesday 06/05. MRDD personnel are moving equipment into the building today. ttc did not approve payapp #6 for triton and asked them to back down the % on the ground floor of the 1956 bldg to 90% and controls down a bit lower.

05/30/07 – TEI was on site on Tuesday for our weekly walk through. TEI contacted the vendors with questions, TCO milestones and coordination issues at the MRDD to help get it ready for the MRDD move on June 4. Items include ceiling layout, diffuser locations, ductwork insulation, EF flashing, fire protection work, elevator details, sump work, controls, etc.

05/21/07 – Triton reported that all fire dampers were completed over the weekend and that mason would be finishing up wall penetrations today and tomorrow. Kevin Maue seemed unimpressed with the quality of the work being done over the weekend and hopes they pass inspection. Tim, Terry and Kevin will walk the site on Tuesday AM to check progress and get commitments for completion.

05/17/07 – Triton will not be working this weekend but will have another crew on the MRDD job starting on Monday and they will have masonry sub helping with the patching of the duct shaft. The new RTU will not be ready to be ran this weekend.

05/14/07 – A lot has happened since my last entry. We have attended weekly const mtgs. We have attended several additional field trips for construction issues. The 1992 building has reached occupancy. The 2nd floor move was today. The 1992 RTU-1 unit is up and running but not under full control. The FM200 system is not fully tested yet even though the room is built. The 1956 building has ceiling grid in the entire building and now diffusers and fire protection piping is being installed.

04/17/07 – TTC and TA walked the project with all the contractors in lieu of a const mtg. The RTU is on the roof. The Server room is nearly complete. Cintas is working to complete the testing procedure and set it up for us. Shaft wall is offset form fire dampers. TE approved shifting the wall over to catch the sleeve of the fire dampers. The wall holes in the return shaft are open and Triton asked about filling them in (even the ones that weren’t done by them). TE directed Triton that they could sleeve the chimney with the boiler vents. The elevator room exhaust fan will reuse one of the openings that goes up to the chimney. A gooseneck may have to be added to the top of the chimney to keep the exhaust 10ft away from our unit intake.

04/13/07 – Project progressing as expected.

04/06/07 – Project progressing as expected.

03/30/07 – Project progressing as expected. Triton reported that the Liebert unit can’t be started today because of a breaker size issue (req 40A we have a 20A spec’d). TA resolved with TJ and Triton will have startup technician for unit after power is provided. TE received submittals today from JCI and GD has been looking them over all day. KM & TTC emailed Triton with acceptance on the retro-commissioning items on RTU-1.

Replace the Supply Fan Motor $3566.00 – APPROVED

Replace two bad compressor contactors $408.00 – APPROVED

Repair static pressure control for exhaust $2720.00 – APPROVED

Replace one bad condenser fan motor $1149.00 – APPROVED OR REPLACE WITH ALL 6 DEPENDING ON PRICE TBD

Perform leak check of Ref system, repair and replace refrigerant as needed $2339.00 – HOLD

Cleaning of coils and inside if unit including combing of coils as needed $1633.00 – APPROVED

Repair of Inlet Guide Vane and controls $3931.00 or replace same with VFD control $5108.00. – APPROVED BASED ON AVAILABILITY OF PARTS

03/23/07 – Lobby lights are on order and all issues with this item have been resolved. 480V electrical service has been changed over by Duke Energy, has been inspected and is up and running. 240V service to be removed the last week of March 07, tentatively.

03/20/07 – 02/19/07 – Last week the structural steel modifications began for the Carrier unit changes. Triton provided a dwg to SSA and CRA showing the new steel dims. SSA is sending back a proposal. The return duct shaft is open and must be sealed to provide proper airflow back into the unit. CRA is checking on some mold issues in that shaft from a roof leak. The FPC dwgs have been submitted to the city. The new FM is located and routed through an existing coredrills in the connector. TA release approval on change orders for TJ Williams for emergency lighting battery packs and dishwasher power in the 1992 building. Triton is still behind on the VAV boxes for the 1992 building and will have to work with Empire for ceiling pad out.

02/09/07 – TTC walked the site last Tuesday to check on progress. MEP is moving forward okay. Duct chase holes have not been cut yet. Triton is installing ductwork on the 1992 building

01/17/07 – Triton has coordinated with Empire on existing work. Kevin is worried that we don’t have enough written documentation that we told Triton to get this work coordinated and scheduled with the owner

01/15/07 – Somehow this job has not been tracked. Here’s the catch up data. The MC is Triton, the EC is TJ Williams, the PC JMW plumbing. We have had several job meetings. Triton has been poorly represented at these meetings. Last week after missing a rescheduled job meeting, Triton came on board with some initiative in catching up to the rest of the contractors. Triton has added more people to the job and is beginning new HVAC in the 92 building. Triton has demolished two floors of work in the 56 building. Triton has 2 men finishing the demo of the multi-zone unit and boilers in the basement of the 56 building. Triton has submitted an RFI questions who is cutting the holes in the shaft floor for his ductwork. TTC responded the GC and this was confirmed by CRA. The RFI was replied to but Empire responded they did not agree. The work is shown on the Demo and Arch dwgs for the GC and the summary of work says superstructure work is by the GC. The CWW has denied the underground permit and tap because the CGE pit is in the Right of Way of the street. TEI has negotiated with them and Cin-Day Fire Protection to provide the new line inside the building through the connector and throughout the building. This should save money and remove trenching and such in the yard. Kevin Maue has agreed that he likes this approach.

09/14/06 – Weekly design meetings continue. We reviewed cost estimates with the client and architect. Final estimates and drawings are due on September 25th for review and bidding is scheduled shortly thereafter.

by TEI TEI

MRDD Thomas Kidd Renovation

08/15/08 – KM reported the variance was accepted.

07/24/08 – KM and Tim Jones rehashed unit placement and easement issues at Thomas Kidd. This issue is still not resolved. The unit has never been moved and has always been on all review prints in this location. It does sit on an existing easement the MRDD already has.

by TEI TEI

MRDD Jackson Center Renovation

07/28/09 – Peck dropped off O&M’s today for this project. TEi will review them and give them to kevin for Closeout.

06/23/09 – This project is complete. I am not sure if Kevin has received all his closeout documents, please check and confirm at monthly meeting.

05/22/09 – We are closing out the project with TJ Williams next week. All work is complete and we are down to closeout documents. the Substantial completion for this project is Feb 20, 2009.

04/15/09 – CR Architecture + Design awaiting final punchlist items to be completed and closeout documents.

04/10/09 – Back punchlist issued.

03/09/09 – TTC sent out punchlist for Mechanical, Electrical and Plumbing and finalized Allowance & Change Order Payments with the punchlist so the contractors had everything they needed to close out the project.

02/19/09 – Project was punchlisted yesterday. Punchlist due at the end of the week.

02/02/09 – TTC continues to attend const mtgs for TEI. TTC drafted a letter to TJ Williams outlining the problems with the licensed door operator. The letter states that TJ Williams is responsible for hiring the licensed sub-contractor per the drawings and specs. We are however going to try to minimize the cost to TJ of this extra work that we believe they did not have in their original bid.

12/23/08 – Const Mtg 9 shows we are ahead of scheduled. TEI needs to address the sink in room 47, the trench drain, the curve sink is on site, the junction boxes need to be t&m by tj.

11/24/08 – Phase II is complete. TEI must do a Phase II punchlist. Peck requested a sheave change on HV-1 to get new airflow. TEI approved and sent to Kevin. TEI completed new allowance change orders for TJ Williams and sent them to John Parfitt for approval. TJ Williams revised payapp was received.

10/16/08 – Phase I is complete. Phase II has begun. Trenching for new sink was this week. No reports after TTC clarification was given. We are hoping things go well.

09/10/08 – TTC contacted Carrigan & Grimm about WC-4 for this project. MRDD wants light lens on warehouse lighting. Phase 2 is first. Phase 1 is second. All permits are approved.

08/25/08 – HVAC and Plumbing submittals are complete. TEI sent back to contractors. Building & HVAC permits are ready. $2,769.87 for the building permit $1,045.45 for the HVAC permit.

08/22/08 – TEI sent permit drawings to the project team. TEI sent out all HVAC submittals back to PHB.

08/08/08 – No update on approved IBI permit for this project.

07/29/08 – TTC hand delivered third revision contracts to KM at MRDD. We are awaiting project kickoff.

Bids came in acceptable. Pre Award meetings are complete. Contractors are approved.

by TEI TEI

MRDD Due Diligence Report

09/21/09 – Kevin called with final comments on the final draft. TTC incorporated these and other comments into the final printing. TEI made 3 copies and mailed 2 of them to Kevin. Per Kevin’s request we will assemble the report into pdf format and email as well.

08/25/09 – Reports are due Sept 4 to Kevin Maue.

07/29/09 – All buildings have been visited. The reports are about 50% updated. Cost estimates, write-ups and grade cards remained to be inputted.

06/18/09 – TTC turned over the final Beckman report for Kevin to review. TTC spent a few days proving that the EH rate is the best rate for the County in these buildings. During this investigation, we found that several of the MRDD buildings are already on the EH rate but are also buying gas on the open market from a broker. TEI has visited Beckman, Northside and Franks so far and completed the field work. The Beckman report is complete. The Northside report is going to be sent to Kevin next. TTC will have Folck & Counts write up and estimate the Franks Adult Center report next week. Kevin said that we will be visiting the schools next. Try to get a date and time for Danny and Amy to do the next building.

by TEI TEI

MRDD 1520 Madison Road Renovation

12/31/09 – There are minor control issues with the economizer in question but the building and controls are working as designed. TTC archived on 12/31/09.

10/20/09 – Kevin has JCI coming out to change return air humidity to zone average humidity on RTU-1 on Thursday. TTC may attend. Three items remain on the service punchlist. TTC contacted Brett to get answers about economizer control,

09/22/09 – I believe the final controls punchlist items to be complete is to acquire the software CD’s and backups of the current settings.

08/25/09 – Building is overcool in mornings. JCI is supposed to come out and fix timing and response loops. Many emails and phone calls have been made trying to get this done.

07/29/09 – Today we received updated As-Builts on the controls for the building from Chuck Vedder. I forwarded them to Kevin Maue for his keeping. I am also printing out a full set of 11×17’s to keep in the basement of the MRDD building again for our use. Everything seems to be running really really well.

06/23/09 – Triton, JCI and Carrier have finally been to the building enough to help us change the programming on the unit to get it to coordinate with the JCI system. This project may finally be finished. Things are working well in both buildings.

02/02/09 – Triton is reporting that Chuck Vetter will finalize this project for 1520 Madison. TTC contacted TJ Williams for the smoke detector work on the 1956 unit. TTC sent a summary email for controls to Brett at Triton for final completion items. The unit is finally running well but the building pressurization & the duct heaters are not working quite right yet.

10/16/08 – Grady says Triton will put together a final sequence that incorporate everything that is really working on the project and the way it works.

09/26/08 – Const Mtg 2 is over. TEI responded to TJ Williams RFi about another light in the hallway. TEI is approving the PHB Honeywell controller for the HVAC unit.

08/25/08 – Triton emailed today to discuss remaining closeout items. TTC explained about OA sensor.

08/22/08 – Triton is ordering a new board for Boiler 2. Triton relieved the pressure on the hot water system to prevent relief valve issues. Triton may have to get TJ to troubleshoot the overloading problem on the recirculation pumps. Kevin emailed that the OA sensor was flaky again.

07/29/08 – Brett Harpel of Triton discussed final control sequences with JCI. Modifications to system were made a month ago and again last week. Building is running fine. KM wants TEI to bless the control system as it meets design intent of the specs. JCI replied stating TEI sequences were in fact in place on RTU-2 but local overrides were keeping the system from operating as designed on that day.

Triton asked to switch the disconnects serving the RCP to troubleshoot low voltages problems at Boiler-2.

Triton to test expansion tank pressure to see if it is causing over pressurization.

by TEI TEI

MRDD – Beckman Fire Protection and Fire Alarm Upgrades

10/28/10 – Phase 1 & 2 are nearly complete. Punchlist is scheduled for 10/29/10 at 8am. All lights are installed, we want to change bulb color to cool white. We need price from Quality on isolation valve. QFP has been advised to replace the ceiling tiles in the main Administration area from the allowance monies.

09/27/10 – Next Const Mtg was rescheduled to Oct 6, 2010.

07/28/10 – Contracts have bee approved and signed. Awaiting first const mtg.

07/23/10 – Bids have been opened and approved. Pre Award meetings are complete. QCE is the electrical. Quality is the Fire Prot Contractor. Contracts have been written and approved by the Prosecutor dept. We are submitting a package back to the bldg dept to answer their concerns on permit issuance.

05/26/10 – Drawing stamped and signed. I will pick up the GC dwgs on Thursday and take to the permit office after Partner meeting. Budgets look good. Emergency lighting is an issue to design and may add more detail to for bidding. ITB was assigned 047-10. Bid set due June 4.

05/16/10 – Client review meeting brought up several issues that we have to fix before permitting. We will permit on 05/26.

by TEI TEI

MRDD Beckman Plumbing Renovations

10/12/10 – Driekast has replaced all the fixtures. The new ones are acceptable and exactly what we wanted.

09/29/10 – Driekast has provided prices for fixture replacement. Need a direction from Kevin on replacement work.

08/13/10 – BBN has offered the fixtures. Driekast is working on an install price.

07/23/10 – We offered to pay Driekast to install new fixtures if they can get the mfgr to furnish better hands free operators on Beckman. No updates on that progress.

05/26/10 – Fixture replacements still an issue. QCE closeout documents arrived this week.

05/10/10 – Terry contacted Dave about replacement fixtures at Beckman.

04/26/10 – Driekast has several items on the punchlist to complete, TTC and KM have to decide about their change order request. Hands free faucets wiggle, wiring has to be strapped down, balancing has to be addressed, they need to add balance dampers.

03/23/10 – Final punchlist has been sent to contractors. Final inspection is scheduled for Thursday. We have to resolve the air flow issues with the GC on this job in the new restrooms. We have to back check the work on the final punchlist, check with the owner and wrap up any final conditions for closeout and acceptance.

02/22/10 – Const Mtg on site. DC attended and back checked punchlist and took additional air flow readings to be distributed. Change order for heavy duty toilet submitted by Driekast and given to Kevin. The final phase is almost complete.

01/28/10 – Phase I is complete. Phase II is almost complete and will be ready to punchlist next week. Some change orders are in the works and we have discussed handicap water closets being WC-2 in both the men’s and women’s restrooms.

12/31/09 – Phase I is complete. Contractors are working over the break to cut the hallways for the new sink. We have approved the stainless warp around by Driekast for the new sink.

09/17/09 – All contracts have been typed and given to Kevin for prosecutor review.

08/25/09 – Bids have been received. Driekast was low plumber. QCE was low electrical. Butler County Electric was low General and Mech.

07/29/09 – This project has been sent out for bidding. Bid opening scheduled for 08/25. The prebid is scheduled for 08/11 at 2:30pm.

06/18/09 – TEI and CRA completed the design development review with the client. This project is scheduled be out for review back to the owner in mid July and bidding in late July.

by TEI TEI

Justice Center 2010 Waterproofing and Sidewalk Repair

Current Comments (2/24/2011)
Refer to F12 for waterproofing work at the Courthouse drive.

Prosecutor cancelled sidewalk portion of contractor work.

Past Comments

Amendment 04 issued for F12.

November 16, 2010 SOSS approved.

November 4, 2010 THP provided SOSS 19 for additional services to F12 – Courthouse waterproofing and sidewalk repair.

by TEI TEI

JC – Kitchen Trench Drain Repair Project

04/05/05 – Queen City Final Paperwork arrived at TE for review. TE to pass along to Mark Hardin soon.

01/14/05 – TTC sent out allowance confirmation with final instructions to Brian for project final payapp.

01/11/05 – Change Order Approved. TTC signed today and sent to QCM for signing. $8028.68. QCM removed 2 hours from the change order for miscommunication on the handle removal.

11/18/04 – Tiles are installed, Steam kettles are being installed by Peck, Piping is being finished by Peck, QCM will be down this evening to anchor the remaining leg holes. TE would like County to request second shift of Peck workers to complete steam piping so kettles are ready to go by 5 a.m. on Friday morning. Bernie is to call TTC back when he gets off the phone.

11/09/04 – The trench pour is not smooth enough to meet TE expectations. QCM said they will grind down to make smooth and backfill with sand based concrete to meet specs. Siemering is onsite prepping the floor for new tiles. Peck will be in next week to finish the steam kettles. The project schedule is very tight and Mark Hardin and Terry Cannon are trying to keep on top of items as they arise to meet the November 18th schedule. The kitchen should be ready for use on Friday morning November 19th.

11/08/04 – Epoxy on new concrete issue has been resolved. We have directed QCM to follow a new schedule. TE and County met Siemering tile onsite this morning to review the scope of the work. Floor tiles won’t be onsite until Monday 11-15. To accommodate this change of plans, the following new schedule is being incorporated into this project: QCM will demo forms 11-08, County will process Siemering background checks, QCM will provide new studs for kettle installation to replace the ones damaged during removal on 11-09,11-10. Siemering will prep the floor on 11-11 and 11-12, QCM will apply first coat of epoxy on 11-11, no QCM work on Friday, second coat on 11-13, attach stainless steel angles on 11-13, Siemering will begin tile installation on 11-15, finish on 11-16, QCM will install steam kettles on 11-17, Peck will finalizing steam kettle piping on 11-18. Project will be punchlisted by TE. This is our drop dead timetable.

11/04/04 – TTC visited the site. Things are going well. New schedule looks like this: 11-4 pour new trench let cure, 11-5 acid etch new conc, 11-8 apply first coat of epoxy, 11-9 apply second coat of epoxy, 11-10 start tile work, 11-11 complete tile work, 11-12 reinstall steam kettles. PHB will have to finish steam piping by 11-11 and we confirmed today with Mark Hardin that only condensate would be stainless piping. PHB can continue to work during the day while QCM works at night.

11/03/04 – PHB is on site doing steam work. QCM has found and removed the valve handle. The new concrete will be poured on 11/04/04. TE has received and is approving the Pitt-Guard Epoxy Coating (USDA and FSIS approved) for use on the trench surface. QCM is to follow the recommended installation procedures for this epoxy for proper installation including concrete preparation (acid etched or brush blasted) and pour the concrete so that it is self-priming as listed on the cut sheet. QCM is also to provide 3 vandal resistant floor drain covers on the existing trench floor drains and two stainless steel angles to support the stainless grating being put back in place.

11/01/04 – Peck reported that they are all set to begin the steam piping work and have coordinated with Tony about this project. They should be on site and ready to work in the next couple of days.

11/01/04 – Karen Draper from the City Health Department returned my call and told me that our plan to install smooth concrete with an epoxy coating in the Justice Center trench with vandal resistant flood drain grates. After my latest phone call we actually discussed the existing condition and the Health Department is very much inclined to approve our new design with the concrete and epoxy coating where we remove all instances of standing water under and around the SS trench in lieu of the concrete/epoxy hybrid. Karen said that any FDA or USDA approved epoxy adhesive and sealant will suffice so QCM can submit what they plan to do to TE and we will look it over and give it our blessing as soon as we receive it. QCM please also provide the stainless steel angle irons along each side to support the grating, attach with adhesive and i would recommend that anchors be installed in the vertical sides about every 6 ft o.c. to hold the angles in place while the grate will hold down the top. PDF

10/27/04 – QCM is on site and beginning work. TTC showed up at 8:00 to coordinate final details as all the important people took care of business from 7:30 – 8:00 p.m. QCM must submit security releases for the coring company TODAY!

10/25/04 – QCM is ready to begin work on 10/25, they need approved security checks but are planning on being on site at 7:30 p.m. as scheduled.

10/11/04 – QCM delivered the project schedule and it was forwarded to MWH and HS. TE approved the timetable for mid November completion.

10/11/04 – TTC & HS are asking contractors to give a price for replacing the insulation on the steam piping behind the steam kettles while the system is down in October/November. TE is exploring a insulation base paint that can be applied. Hy-Teach Thermal Solutions www.hytechsales.com is one site for this paint that TE found. We have calls into their engineering department for information.

09/24/04 – Notice to Proceed was given by MWH via a phone call. TE to setup up pre construction mtg.

Bidding is complete.

by TEI TEI

JC – Battery Room Fan Coil Upgrade

07/12/05 – Final paperwork arrived today for this project. TTC copied the documents, signed the final pay application and dropped it into the mail for the County.

06/30/05 – Filter rack is fixed. Final paperwork has not been received. O&M’s have been delivered to the County via a hand delivery by TTC. TE contacted PHB today about final paperwork status. PHB will get out soon.

06/15/05 – O&M’s Received today. They were approved and will be held for final payapp paperwork. PHB should be down in the next day or two to fix the filter rack and then they will send in all final paperwork.

06/09/05 – PHB called to questions the $2500 allowance on this job because it was not in the contract. TTC checked the bid documents and the contract and PHB is right. The original bid of $24,400 did not leave enough room to add the allowance alternate without going over the $25k limit by ORC. We will now have to write a change order to increase the budget slightly to accommodate the newly added automatic air vents which are a necessity in a chilled water system like the one we are using at the JC. Final paperwork should be down to you in about a week.

06/07/05 – PHB reported that they would be down to finalize the punchlist about June 20 – 27 or so.

06/01/05 – TTC sent an email reminder to PHB to finalized punchlist and finish the filter box.

05/05/05 – Punchlist distributed to PHB. PDF

04/05/05 – TE stopped at site today to punchlist the FCU job. SBO did the site work. Punchlist will be distributed very soon.

03/15/05 – PHB requested a change order for the new automatic air vents. TE asked for the cost and will suggest the County accept this allowance usage as part of the project that wasn’t included in the original scope. We are awaiting a cost associated with this additional work. I can see now that the last payapp was approved for too much money because we have a $2500 allowance that hasn’t been used at all, but that is okay now because we need to pay for this additional work and that will come out of the allowance. I guess PHB didn’t break out the allowance as a separate line item on the schedule of values.

03/15/05 – TTC coordinated with Howard and Bernie since payapp only left $2300, TE wanted to make sure that controls and electric and vents were all installed and working by the end of the week so that TE can do a punchlist soon. TTC left a messages on Bernie’s cell phone.

02/28/05 – IBI Permit set in PHB hands. Fill up and startup of new FCU should happen within a week of electric hookup and IBI approval. Tentatively we could be looking at startup at the end of next week.

02/25/05 – TTC notified PHB today that the return transition is not necessary on this project as it would prohibit access to the return filters. The supply transition is being fabricated at this time and PHB should be out to install it in the next week or so.

02/16/05 – TTC checked on the status of this project today. Piping looks good. Insulation needs to be installed. Thermostat has to be installed. Unit has to be started. Pressure test has to be removed. Final piping must be filled and checked.

02/07/05 – Doug Turner was onsite making the tie-in connections over the last weekend. They drained the system and refilled it and made the tie-ins

02/02/05 – Permit has been applied for by Peck for this project.

01/25/04 – Peck was onsite to review project. TE is to provide stamped signed dwgs for permit and to discuss supply diffuser and return grille on fan coil unit.

09/23/04 – Bids Received and Posted. TE awaiting County’s approval to proceed with contract. See bid results below

09/13/04 – Bidding Underway

by TEI TEI

JC – Task Force Room Upgrade

09/18/05 – The city has approved the whole permit. This job is complete.

07/22/05 – Steve Ossege called and said the City does not know of any approved independently support handrail system that meets code. He will however approve the platform arrangement discussed on site.

07/18/05 – Ed Enderle of DK called today and quoted $2450 to build a platform and railing for the Task Force Unit. TTC asked them to look into temporary handrails that were eluded to by Steve Ossege. More information to follow.

06/27/05 – TTC and DK have been working on getting the permit issued and closed for a week. Amit Gosh called and verbally told me that he was signing off on the permit today. TTC relayed this information to DK for finalization.

03/24/05 – JCI reported that wiring and status condition has been wired from Task Force unit to DDC. Programming is complete and task is complete.

03/22/05 – DK delivered the final payapp today along with all the final paperwork. TTC will sign and send to Diedra in the Sheriff’s Dept for project completion.

02/24/05 – Debra Kuempel (DK) deliver payapp #2 today. It was approved by TE. All final paperwork should be complete next week.

02/22/05 – DK stated that all final paperwork was sent over last week to TE. TE is going to finalize and send to Diedra Burns this week.

01/25/05 – TE contacted Ed Enderle as all work is complete and signed. TE request final paperwork and signed punchlist!

01/24/05 – TEI called DK and spoke to Michelle Combs in regards to when we may expect Pay Application #2 and final documents. Michelle was not aware that the project was finished and said that she would check with DK personnel and will call back, I told her that Ed Enderle may be of assistance.

12/15/04 – Debra-Kuempel Pay Application #1 not found at Hamilton County. Transmittal and ThermalTech records indicates that the application was mailed September 24, 2004. Doris called and left a voice message for Michelle Combs informing her that the application was lost and asking her to resubmit an original application.

10/12/04 – DK has finished project. The allowance has been used to get a new power source for the computer room unit that doesn’t shutdown the UPS system. TE has instructed DK to finalize paperwork and punchlist to complete this project. Howard reports that things are working okay to date (4 days runtime at this point)

09/29/04 – TE emailed DK for an update and schedule.

09/23/04 – TE directed Debra-Kuempel to redo their Change Order/Allowance price to re-use the existing MCC starter that use to be for a now abandoned pump to the get the cost under $5000 so we can use the allowance money without writing a change order for this project. We have $5000 in allowance money to use towards re-doing the power to the computer room unit. Once this is complete all work on this project and the punchlist items should be taken care of at the same time. Look for completion by September 30, 2004. The will coordinate all field visits with Howard Seal before starting on site.

TE has to finalize power change work with Debra Kuempel.

TE has distributed punchlist.

Compressor short cycling has been eliminated by utilizing the hot gas bypass feature to change the performance of the system.

TE to work on a new power source must be found and the unit hooked to emergency power panel instead of the current UPS connection.

by TEI TEI

JC – Exhaust Fan Upgrade – Project Status

06/15/05 – TE received final paperwork and change order signed for this project. TTC will be mailing it down to Diana with final balance report for HC records. This project is complete. The County is getting $3256 back on this job! O&M’s received and forwarded to County.

05/25/05 – PHB Jerry Govert called today to organize final paperwork. TE sent out a change order to recouple the $3256 in remaining allowance money. All final paperwork should be received in a couple of weeks. PHB should include the CO Credit with final paperwork.

04/22/05 – TE does not or has not seen final paperwork on this job. The balance paperwork was marked as received but no pay applications have been received. Please submit with all final paperwork for payment. Please provide all final paperwork and allowance usage details because we did us the allowance on this job.

02/25/05 – TTC ask Ken Shad to give me final paperwork and allowance change orders to wrap this project up today. He said that he would get right on it.

02/16/05 -TTC has requested all final paperwork on this project including balance reports. PHB sent to TE office on Tuesday. TTC to review.

01/18/05 – PHB stated that all work is complete including the 2 new 5 HP motors. The exhaust % increased by 18-20%. TE is awaiting all final paperwork on this job for closeout. Rick Ahrens will stop by to take a look at a motor squeal on Thursday.

12/27/04 – Howard reported today that EF-C1 has high pitches squeal. TE called Peck, but they appear to be out of office today. TTC will try first thing Tuesday morning again.

11/26/04 – PHB has acquired all existing cfm data from the fans. TE is reviewing to provide a direction to the PHB construction team. It appears as though the allowance money should be used to replace the motor on EF-D2 from 3 to 5 HP. The wire size is sufficient. The existing 3 hp bearings will be taken out and put on EF-D1 as that fan needs a bearing replacement. PHB reported that even with the duct cleaning there are still some security grilles that are clogged with dirt and the secure design makes them virtually impossible to clean out from inside the cell area. TE to advise how to handle this situation. Sam Boehringer will be reviewing submitted data on Monday November 29 at TE and then we will setup and onsite meeting to go over the project with the owner in the coming week. We have a $5000 allowance to use for these repairs.

10/15/04 – TE has contacted PHB about and updated schedule for this project.

10/05/04 – Pre construction meeting was held and PHB plans on starting in 2 weeks. Schedule is basically one EF per month.

09/20/04 – SOSS revised for additional Champion monies and will deliver on 09/21/04.

Champion provided duct cleaning on existing systems. Hired through ThermalTech.

Peck Hannaford & Briggs won the bid ($26,050). Base Bid, Alt 1, Alt 2 and $5000 Allowance. EF-C1, EF-C2, EF-D1

Contract written and being reviewed.

by TEI TEI

JC – DDC Upgrade

03/03/06 – JCI finished the upgrades on Wednesday. The original CD’s were delivered. Some work need to be done to fix what the upgrade busted and then a new backup needs to be done. Then Tony will pay for project. TEI told the contractors this by email today for JCI.

02/22/06 – TTC received final paperwork and approved for this job. TTC is going to approve and give to County even though final CD’s and backups has not been received yet. This is the only JCI project that remains to be backed up and CD delivered to the Owner. The County may at their prerogative continue to hold final payment on this invoice as they do not have the licensed software in hand yet. TTC has emailed Al Bell and has not heard back about getting final CD’s to Howard Seal for this job so I doubt Tony is going to release this money until the material is in his hands. Please let ThermalTech, Howard and Tony know when it arrives on site and please do this ASAP.

01/18/06 – Wayne Jackson reported that a printer with a USB connection would have to be used for the JC ADS server since no printer ports were available. TTC informed JCI that the County has already taking care of this item. Wayne said that he would have Mike Asbrock call Howard Seal directly and prescribe a safe and easy method for deleting the 93,000 alarms in the alarm container. The licensed CD and Backup still have to be provided to the County on this job. No remaining punchlist items remain.

01/10/06 – TTC emailed JCI about installing a report printer on this project and about closeout CD’s for this project.

12/14/05 – TTC emailed Tammy back at JCI to inform her that the project is not complete and cannot be paid until everything is done.

10/24/05 – TTC contacted JCI again about punchlist by email.

10/14/05 – TTC and JC have not seen any final punchlist from JCI yet. TTC contacted main office to get closure on this project!

10/03/05 – JCI reported that they will finalize the punchlist in the next two to three weeks. TTC told Howard about this schedule.

09/16/05 – TTC contacted JCI about received punchlist and trying to get a schedule for completion today.

09/08/05 – TTC contacted JCI and approved their final two pay applications but requested all the final paperwork and TTC has to send out punchlist immediately on this job. Punchlist has been sent to JCI and County.

08/08/05 – JCI Reported that VFD’s have been reprogrammed and are 100% operational. TTC to provide final punchlist on Tuesday, 08/09. When complete this will finalize this project.

07/21/05 – Tony called to say that escrow on the JC DDC project has voided App #14. JCI should resubmit App#14 for $20,921 and App#15 for $23096.00. TTC instructed JCI to send in all final paperwork as well as this project is complete except for the punchlist.

07/05/05 – JCI is re-sending payapp # 14 which appears to be correct. TTC reviewed and asked for 1 original and 3 copies. JCI and its electrical contractor are installing the bypass around the VFD’s. They should be done this week. At that point the project should be complete and TE will do one final punchlist on it.

06/14/05 – JCI prematurely sent a final bill into TE. B&J and JCI aren’t scheduled back onsite until 06/27/05 for the final phase or work on this job. B&J called today to say that all the sheetmetal is prepped for installation. B&J also said that two of the boxes were ordered as 6″ inlet and 10″ inlet but actual field measurements showed the boxes they are replacing are 8″ and 8″. TE approved the new sizes as they can be balanced to what we need in the space. JCI is on vacation

05/02/05 – B&J confirmed that all VAV boxes should be shipped the first week of June. TE advised the JC FM staff.

03/24/05 – JCI has automatic reports working for Howard. The new cooling tower sequence is the only project punchlist item remaining. All commissioning reports (hardware verification) sheets are in the as-builts. JCI will provide new as-builts to insert into set based on the change order work scope currently going on.

03/16/05 – New quote from A&M came in for DDC work at $12,091.53 for B&J Jacobs to do the actual installation. Since they were the original contractor, this would be great. TE can supervise the installation and just turn it over to Ed Braemer when we are complete or still have Ed work with B&J during the installation. I need approval from someone before we let them go. JCI reported that wiring and status condition has been wired from Task Force unit to DDC. Programming is complete and task is complete.

02/28/05 – TTC met with Howard, Peck, Debra and Langdon to solicit additional prices for the Dual Duct Box replacement at the JC. Quotes are due this Friday at Noon. I will coordinate with Ed Braemer on moving forward after we receive prices and we talk to Tony and Howard about authorization. In the meantime, I have asked Ed to call Howard to discuss reporting problems with the new software.

02/16/05 – TTC is collecting a packet to get Peck and Debra to provide a follow up price on the Dual Duct Boxes in the Sheriff’s Area. Howard reports that JCI is almost to the point where they are out of work.

02/15/05 – JCI gave us a price of $3557.40 to hardwire the VFD’s to the Fire Alarm Panels. Should or can we approve this work?

02/03/05 – Construction meeting complete. JCI to install final version of software next monday. TTC to gather additional mixing box estimates for AKM. Fire alarm issue should be hard wired through VFD’s so they can’t be overridden from the front end software (JCI to work on).

01/25/05 – JCI came onsite last night to work on 2 dual duct boxes in Sheriff’s area and to work out some issues with EDB. TE to follow up with HS today on status. Existing boxes don’t have accessible actuator shafts for DDC upgrade. We may have to improvise to modify them for retrofit. EDB, TTC and HS to meet and discuss if needed.

01/24/05 – Ed Braemer or Johnson Controls, calls TEI on required app doc and will follow-up with Tammy of JC to see if she has everything needed to submit the apps.

01/21/05 – Received JC Invoice for Application #13. The request is a 4 page Original Invoice form. Tammy of JC contacted and asked to submit Application on the standard AIA forms and that a completed Affidavit form should accompany the AIA forms.

01/19/05 – TE stopped at the JC to investigate the “howling” on the Stairwell Pressurization Fan.

11/09/04- Submittals on the VFD fans returned. Pay Application #11 returned to Ed at JCI. Both returned to JCI via UPS.

11/08/04 – Voicemail left for Ed Braemer regarding incorrect Pay Application #11.

11/08/04 – TE is reviewing the submittals on the VFD fans this week. They should be complete and returned this week.

09/23/04 – JCI needs County approval to begin. What is the status of the project. They are waiting on a go ahead from you.

New change order approved. Waiting to begin new work.

JCI has completed all phases of the project except for the cooling tower programming.

A new round of change orders has been accepted by the County and the paperwork is en route to JCI for signing.

Steve Steinkuhl to issued a final punchlist for the project by September 3, 2004.

by TEI TEI

JC – Aramark Office VAV Upgrade

05/23/06 – TTC moved this project to the archive as all work on this job will be done under the fingerprinting ID project that is still open.

11/08/05 – TE heard back by email form Jill Williams and we will set up a quotewire job for this.

11/04/05 – TTC and Tony discussed including the purchase of the Dual Duct boxes today and decided to buy this box with the other 8 for Fingerprinting and ID. Trades will install these boxes.

10/28/05 – Prebid meetings were help on other JC projects but this job will be held for Trades to do work.

10/14/05 -Design is complete. TE will have trades do the work on this job to replace the existing VAV box with a larger box and redo the controls. No permit is necessary on this job.

10/04/05 – TTC & SRB have visited the site and discussed VAV options with Howard. Since this space is not part of the Dual Duct air handler, there will most likely be overheating and over cooling of the office during low load periods as the minimum setting on the VAV box is 145 CFM and the air handler supply temperature is reset based on OA and averaged room temperatures. TTC emailed Tony about just having the work done in house and having JCI do the controls directly.

by TEI TEI

JC – Boiler Room Exhaust Fan Upgrade

05/23/06 – TTC archived this project because there is no way to make this work without a lot more make up air into the room.

10/28/05 – TE still has not found a viable solution for the heating problems in the space and will look at other usages.

10/24/05 – TTC has not found a usable scheme for boiler exhaust to date. There is not enough cold air to make a dent in boiler room temperature without additional cooling.

10/14/05 – TE cannot find an exhaust fan design that works without additional cooling.

10/04/05 – TTC & SBO have visited the site and are designed the new exhaust fan design to help remove heat from the boiler room.

by TEI TEI

Justice Center – Backup Chiller Project

09/01/06 – TTC worked out final billing details with Bridgette today by phone. DK is sending over final invoices for payment.

08/30/06 – Steve and Terry finalized the number required on the final payapps. Laurie is typing up the final change order for the project. TTC has to coordinate the change order payment with Tony Matre.

08/29/06 – TTC reviewed the final pay application by Debra Kuempel. Change orders amount were added to the contract amounts and this is not correct so TTC has to contact DK to figure out the correct billing method.

04/03/06 – JC asked DK to insulated the chilled water lines to prevent pooling under the chillers.

03/24/06 – DK has not been on site doing any work. They may wait until it gets warmer outside to make the boiler changeover work. The chiller scope is nearly 100% complete.

02/23/06 – TTC was on site today to look over chiller project. New 12″ mains look great and are installed and cored drilled and anchored.

02/22/06 – We approved some small allowance work today described as: *Cut (2) 11/2″ nipples off existing thred-o-lets on Chilled Water Supply and Return piping. Weld on B/W caps on to existing thred-o-lets to cap lines off. *Install 2″ thred-o-lets with ball valves and caps at the bottom of 12″ Chilled Water Riser for drains on both supply and return piping. Total cost for the work outlined above is $642.00

02/16/06 – Howard asked about replacing valves if they don’t hold and about removing a 1-1/2″ nipple in piping when drained down. These are two agenda items for Friday Const Mtg.

02/10/06 – TTC release this email after talking to Howard this morning:

Steve, When you go to the drain the Chiller plant at the JC, it was our belief in the bid documents that the butterfly valves near the chiller will hold the water in the system and prevent a complete drain down of the mechanical plant. If this turns out not to be the case and we have to drain down the entire building, there is a concern about who is replacing the nitrate in the CHW. It is my belief that this is a condition that will be covered by the allowance we have on this project if it proves necessary. I don’t foresee this being more than $500 or so if in fact we need this. Please let me know how the valves hold next week when you begin to tie into the existing system. If you have any questions, please let me know. thx, ttc

02/02/06 – TTC contacted DK about working on the chiller project and they said they would get started.

01/17/06 – TTC held the pre construction meeting today and distributed the meeting notes. The boiler project will start 03/15/06 and last for 60 days. The chiller project will begin immediately and last for 60 days.

01/10/06 – Pre Const Mtg scheduled for 01/17. Everyone has confirmed for meeting.

12/13/05 – Contracts were returned for corrections. TE corrected, printed and returned to the County for submission to the commissioners.

11/08/05 – Tony gave TE the okay to type the contract for Debra Kuempel on this job. We are taking the base bid and $2000 Allowance.

11/07/05 – TTC opened and tabulated bids for this project. See link below.

10/28/05 – Prebid meeting was held with all contractors. Bids due Friday to TE.

10/24/05 – Drawings are complete up to TTC’s review. Prebid is set for this Friday.

10/14/05 – TTC promised final design complete on Wed 10/19.

10/03/05 – Tony approved our recommendations and we are going with the plan to put in the chilled water tie

09/29/05 – TE recommended to the County that no electrical xfer switches be included in the temporary job because the expense of putting in electric switches just isn’t cost effective to providing emergency generators on site at the time it is needed. Feeding the chillers is problematic. TE will bid the mechanical work out as a small project.

09/27/05 – TE pulled all the old designs from the archive and setup the new project. Electrical is a big issue as this will put the project over $25,000 unless we use generators.

by TEI TEI

JUSTICE CENTER – Backup Boiler Project

09/01/06 – TTC worked out final billing details with Bridgette today by phone. DK is sending over final invoices for payment. TTC sent out the change order for the job to Steve V.

09/01/06 – Tony approved the $1597 change order by phone today. TTC to advise DK to wrap up the project.

08/30/06 – Steve and Terry finalized the number required on the final payapps. Laurie is typing up the final change order for the project. TTC has to coordinate the change order payment with Tony Matre.

08/29/06 – TTC reviewed the final pay application by Debra Kuempel. Change orders amount were added to the contract amounts and this is not correct so TTC has to contact DK to figure out the correct billing method.

08/24/06 – Al Bell reported that the high pressure sensor was re-ordered and will be installed very soon.

06/26/06 – Howard reported that DK will be onsite tonight and begin the work on the final changes to the job. JCI and DK have not been in contact with each other yet on the new sensor.

06/06/06 – TTC contacted both JCI and DK again about progress. JCI came in last week but needed a different part to make it work. DK is supposed to be down this week but TTC will find out more details from Steve Vater and let everyone know. As of Monday, Howard had not heard from DK about coming back.

05/23/06 – TTC emailed both Debra Kuempel and Johnson Controls to try to get them scheduled to finish this project up. TEI will pay the JCI cost to the job and DK will bill towards the allowance to wrap this up.

04/03/0 – DK is going to be on site the week of 04/10 to finish up the boiler work. JC asked for insulation on chilled water system to prevent condensation. TTC relayed this information to Steve Vater and asked DK to coordinate their time and visits directly with Howard Seal.

03/24/06 – DK has not been on site doing any work. They may wait until it gets warmer outside to make the boiler changeover work. The chiller scope is to b nearly 100% complete.

02/23/06 – TTC visited the site today with the contractor and Howard. TTC wants to remove a check valve from the main steam supply from the courthouse because the pressure drop will be to great to keep this device in line. An existing chain operated gate valve on boiler #3 will suffice in place of the one noted on note #3 on the dwg. The non-return valve can come out as planned. At this temperature outside the steam plant from the courthouse can’t be shutdown until 7pm and it must be back online at 3am. DK will have to coordinate this with Howard and the CH staff.

02/22/06 – We approved this allowance usage today:*Install (2) 3″x21/2″ weld reducer on new 3″ pump condensate piping at tie-in point to existing piping. Existing pump condensate piping is 21/2″ in size. Debra-Kuempel quoted total cost for the work outlined above is $261.00

01/17/06 – TTC held the pre construction meeting today and distributed the meeting notes. The boiler project will start 03/15/06 and last for 60 days. The chiller project will begin immediately and last for 60 days.

01/10/06 – Pre Const Mtg scheduled for 01/17. Everyone has confirmed for meeting.

11/08/05 – Tony gave TE the okay to type the contract for Debra Kuempel on this job. We are taking the base bid and $2000 Allowance.

11/07/05 – TTC opened and tabulated bids for this project. See link below.

10/28/05 – Prebid meeting was held with all contractors. Bids due Friday to TE.

10/24/05 – TTC & SRB visited the site to pickup some last minute details about returning condensate the DA system when the Justice Center is on its own steam. Drawings are complete up to TTC’s review. Prebid is set for this Friday.

10/14/05 – TTC promised final design complete on Wed 10/19.

10/03/05 – Final design is being complete. Job is ready to bid out for small projects.

08/23/05 – TTC visited the site and found that bypass valve is already in place so that piping work does not need to be done. At this point there are a couple of viable options: 1) put in a new valve near the jc side of the bridge or automate an existing pneumatic valve in the courthouse. TTC sent an email to Tony, Howard and Mike looking to explore our options.

by TEI TEI

Justice Center North Electrical Device Identification

01/10/07 – Project completed in 2006.

08/24/06 – Drawings were delivered to Hamilton County Facilities. There are no open issues on this project.

05/17/06 – Close-out meeting; no open issues. All work by Queen City has been completed. I will be working on getting the drawings completed within the next several weeks and send (3) full sized sets to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Howard Seal). Queen City presented (2) binders, each with copies of the old and new panel schedules. On the new panel schedules the items listed in bold have been verified; the items listed in normal type are transferred over from the old schedules. The binder also includes diagrams of the MCC’s and Switchgear. A CD was also provided which contains all of the new panel schedules in Excel format.

11/15/05 – Contracts have been submitted.

by TEI TEI

JC – Replace Sallyport Sliding Gates

02/19/07 – TTC is approving and passing along all the closeout documents for this project. TTC is archiving this project today.

01/02/07 – TJ Williams submitted a letter expressing their extreme disappointment with the County not approving the Change Order for the extra work incurred during the Sallyport project. TTC forward to Mark Donnelly for the County to file with the records are address separately if they deem it necessary. PDF

12/26/06 – TTC archived the project today. Legacy has the smallest bit of paperwork to turn in on this project. Once Donnelly has that, this project is complete.

12/14/06 – Mark called today about closeout and Legacy reported: The guides at the justice center are painted. I am not sure the status on TJ Williams. Kenny is in a meeting until about 5. As soon as he gets in I will get an update and get you the information. I have been working on the close out documents and we are about 80% done. We still need to finalize the certified payroll from two of our sub-contractors.

12/04/06 – Close out meeting was held today. 3 items on punchlist were discussed. Pay App 01 was discussed. Legacy to provide a revised Payapp 01 with allowance shown. TTC signed payapp 01 for Mark D to use as soon as a revised schedule was provided by Legacy. Legacy provided one by email but TTC edited it and sent it back out to the contractor with 8% retainage per the project manual. TTC also sent Allowance Clarification 01 to the contractor for signing.

12/01/06 – TTC sent out formalized punchlist items today and reminded everyone about Monday’s meeting. PDF

11/20/06 – The project is complete. There are two punchlist items which remain: touch up paint on door frame and the control panel box that TJ damaged during construction. The doors are working, the fencing and hardware are all installed properly and TEI visited the site for a punchlist walk through. The TJ change order was denied but it was agreed that the allowance money could be used to pay for the extra expenses. TEI has asked for closeout documents on the whole job to wrap it up. We are having a closeout meeting on December 4, 2006 as a final meeting.

10/10/06 – The structures are installed the concrete is finished. We need to let it cure for about 2 days before we install the overhead doors. TJ Williams has some temporary power and control patches installed in the building that have to be replaced. Overhead door is supposed to bring down the doors on Friday of this week. TJ Williams has not started the change order work on the loop control. We are hoping they will do that this week when they come down to do the rest of the work ready on the project. Legacy has a new PM on the job. Ken has taken ill. Howard was able to get the new guy into the building for the concrete today. In a week or so we should have operation of the new doors. The Sheriff may re-stripe the parking lot and move the dumpster outside the scope of our job.

09/15/06 – Legacy contacted us today to report the steel has been ordered. 09/25 looks like a delivery date on that. The doors are in town and waiting to be installed. Legacy may be down next week to do some anchor work on the new structure and Howard confirms that all security cameras are still in operation after TJ Williams completed their partial demo. Saw cutting and more general work scheduled for mid week of next week.

09/04/06 – Change has left our office to the contractor. Project should move forward immediately. Contractor planning on being on site this week.

09/01/06 – Change Order for $3775 was approved today by Mark at meeting in his office. Sheriff will do demo of existing fencing. Legacy will sink new polls for fencing company. TTC approved stainless steel anchors for this project since galvanized were hard to order on time. TEI to write change order asap.

08/30/06 – Legacy reported an error in the change order pricing and a misunderstanding on ThermalTech’s email. New pricing was sent to TTC by fax today. TTC will forward the information to Mark Donnelly on this project for a decision.

08/29/06 – Change Order prices came back today as $1800 for fences with Sheriff demolition and $7195.00 if Legacy does everything on the job. TTC passed along to County. HS was checking with Sheriff’s for their willingness to still do their part.

08/24/06 – Today we approved the electrical submittals and asked for an immediate change order cost for the new door location on this project from Legacy. We have been waiting at least one week for this amount.

08/16/06 – We have had two kickoff meetings for construction. The first issue was the location of the new doors. The Sheriff has now put forth an edict that the busses that enter the site should be closed between the doors. This was not the case during the original design. TEI has issued a change order to collect costing information for this change order. It shouldn’t be too much money as the a lot of the scope is already included. The Sheriff is willing to provide additional demolition of the existing fencing to make this work and Howard stated that the existing dumpster fencing must remain as is, TEI said that is fine. TTC called Ken Jones today and found out that the CO cost is not ready yet and the security checks have not been sent in yet. TTC advised Legacy to move forward immediately with providing this information and the job is ready to begin. Spencer approved the submittals on his and TEI will distribute them as approved to the vendors.

07/25/06 – We received today that the contract was approved. Pre Construction meeting scheduled for 08/02 at 2pm. TEI to contact Legacy Construction for meeting.

06/19/06 – TEI revised the contract and mailed 3 originals to Mark Donnelly for submission for approval. After approval we will convene a kickoff meeting for the owner and contractor.

06/01/06 – Contract was submitted to prosecutor for review.

05/25/06 – Contract was written for Legacy and mailed to Mark for approval.

05/23/06 – Bidding is complete. Addendums were issued. Bid were opening on schedule. TEI drafted our first contract and will give it to Mark for submissions to the prosecutor department. We luckily increased the project budget because the job came in over the estimate but within the 10% margin but still under the actual Archibus estimate.

04/11/06 – Final Client Review sent was sent to Howard and Mark including the work from CRA and THP. There are still some issues to resolve before bid issuance. Mark and Terry increased the budget to $80,000 to cover the additional work added to the job by a more extensive structural drawing set and permits being added back the project. Security Cameras and lights had to be moved to allow for new overhead door placement on project. TTC coordinate with Howard Seal on this point. TEI located the South camera and light on the dwgs and in talking with the Howard the North camera and light were also located on the plans accordingly. County comments are due back on Friday. Final set is due to Purchasing on Monday morning.

04/05/06 – A second design meeting lead to some changes in the door placement. TEI developed a new drawing showing the perpendicular arrangements of the new overhead doors and modified the drawings notes accordingly. Power will be brought from existing connections. The razor wire will be removed and replaced by the contractor. The north door shall be the first to be replaced. The old sliding gates will be completely removed from the job. TEI added special notes to allow other manufacturer’s to be submitted on but they must be approved before bid day by the engineer. The specs were updated and the ITB number and dates have been acquired. A review packages was sent to the team for initial review.

03/21/06 – TTC and Mike Haney spoke today about the project’s structural requirements and THP had some great ideas. Since these are overhead doors they can be turned perpendicular to the building and only be 15 ft wide thus making the doors smaller and the support steel smaller. This sounds like a great idea. TEI will approach the owner about this condition.

03/16/06 – Design meeting schedule on site. Sheriff has approved type of door that we will use for this project. Now we just have to finish the structural design and bid the project out.

02/10/06 – TEI release a review set for comments to the County for this project. TTC still has to discuss product with overhead door and finalize ITB. We have not requested an official ITB from purchasing yet.

02/08/06 – TEI and CR met on site with Howard and Tony and reviewed scope of work.

by TEI TEI

JC – ID & Fingerprinted Dual Duct Box Upgrade

03/15/07 – All Closeout is complete. Howard is satisfied but JCI has a bit of tweaking to do on the sensors. This project is ready to archive.

02/28/07 – JCI dropped off as-builts and O&M and final payapps for this job. TTC to approve and send to Tony. TTC approved and gave to Laurie to send to Tony.

08/30/06 – Triton has all schedules for the proejct and are putting the submittal data together right now. They are not sure about having 8 or 9 vav boxes on their PO.

08/23/06 – Construction Meeting Held. Job should be completed in 60 days. Triton to provide equipment, JCI to provide installation power and controls. PDF

05/09/06 – TTC compiled the complete state term contract and will deliver it to Tony’s office in the morning of 05/10/06. TEI is sending 4 copies of the contracts, spec book, and 1 copy of the JCI bid documents.

02/16/06 – Tony asked for final contract documents to be sent over so he can move the project forward. TEI will have these to Tony by 02/17/06.

01/10/06 – JCI has provided all contract documents for a state term contract. Sam compiled everything and TTC needs to review it. TTC will then deliver to Tony for use on this project.

12/13/05 – TTC has contacted JCI numerous times about submitted an actual state term contract. Al Bell will work with Joe Schutte and get one prepared for this job.

11/08/05 – TE heard back by email form Jill Williams and we will set up a quotewire job for this.

11/07/05 – TTC opened and tabulated bids for this project. See link below. The project bids came in over $25,000. Tony and TTC decided that the best way to get the job done at this

10/28/05 – Prebid meeting was held. Bids are due 11/04 to TE.

10/24/05 – Project is ready for TTC review. Prebid is this Friday at the JC.

10/14/05 – Project is ready for review and bidding.

10/04/05 – TTC & SRB have visited the site and are redesigning the dual duct boxes. We have contacted JCI about scope of Sheriff’s upgrade to use on this job. Drawings are done. Specs need to be finished and scope needs to be ironed out.

by TEI TEI

JC – HTX Upgrade

02/26/08 – Final paperwork, O&M’s and As-Builts have been sent to the County for closeout. This job is complete.

02/21/08 – TTC received the final pay application on this project and signed it immediately and sent it to Tony for closeout on this project. ThermalTech agreed to do the as-built drawings for Tony on this project.

12/05/07 – Today we worked on allowance cost and change order submissions by PHB. TTC asked for finalization and closeout paperwork from Bernie on the job. 3 allowance have to be finalized and TTC asked the County if there were any other tasks on this project they needed done with the remaining allowance money.

11/15/07 – PHB reported today that the equipment went in exactly as drawing except for some minor change in piping right above the heat exchanger. TTC said he would accept a markup on this change and incorporate it into the electronic dwgs for the job and give it to the County for final As-Builts. PHB is working on final closeout documents this week and should have them to TEI for final approval very soon, they only items they are waiting on is JCI sequences.

10/17/07 – JCI does not have the right sensors in the HTX system for their controls. We are investigating replacing them or using an old one. this needs to be resolved quickly.

10/01/07 – TTC & GD visited the site last week and reviewed the project installation. The job was very well done by PHB. GD requested that the strainers be installed in the horizontal position to alleviate water hammer. This is an arduous task for PHB but they are going it today. TTC approved payapp #2. JCI said they were almost finished with controls but had to add flex tight to the controls at the valves. Howard was going to run the system in manual to see how it works. TEI still must provide a punchlist for this project.

09/12/07 – Howard called with a questions about isolating the PRV for maintenance. TTC said it was isolated on the PID. TTC will stop down tomorrow to review with contractors and will setup a progress meeting with GD to go over design.

09/03/07 – In the past week we have approved a pad extension onto the front of the existing HTX pad layout to allow the piping to be connected to the existing piping easier. We have approved new butterfly isolation valves from the allowance monies and TTC approved payapp #01 for the project. In the field TTC approved a slight modification to the layout of the steam traps for easier maintenance and TTC and Walt discussed the atmospheric steam vent that is supposed to be connected to a tube inside the trap and came shipped separately.

08/23/07 – PHB asked about new steam control valves vs old. TEI directed them to the Add01 where hot water control valves are reused but not steam valves. Steam are new CV-27 types and 3″ in size. Howard asked about HTX loads and copper tube sheet strength. TTC asked GD and TZ for info. TTC scheduled a field trip on Friday to the building to review the progress of the project.

08/20/07 – PHB did get the new steam isolation valve installed. The Spence valve and 2″ gate valve was also replaced. Howard said that 2 of the hot water isolation valves are corroded and need to be replaced. TTC asked PHB to give us a price to replace this valves. These particular valves are not more than 5 or 6 years old as they were put in during the

08/17/07 – PHB is installing the steam valve, Spence valve and 2″ gate tonight after hours. TTC will check in with JC tomorrow on progress.

08/03/07 – TTC gave PHB requested information on the Spence valve, isolation valve and existing steam pressures. The Spence valve has unions on both sides for easy replacement. The isolation valve shows 6 inch on the dwgs and with the insulation on the piping, there is no easy way to confirm it isn’t 6″. The steam pressure is 65 psig max for normal operating pressure so PHB’s request to buy a 150 psig isolation valve will be fine. TTC & HS picked the location so that the HTX could be isolated from the system for this work.

05/30/07 – TTC contacted PHB about the equipment orders. The HTX vendor asked for more money for sizing the HTX at 10 psig but TEI emailed PHB and showed that 10 psig was called for on the schedule. This may hold up equipment ordering.

05/21/07 – TEI returned submittals via email to Bernie Tobergte today. Now comes the equipment order and scheduling. TEI will deliver hard copies to Tony and Howard.

05/14/07 – TEI has all submittal data and will review on 05/16/07 and return to PHB by week’s end.

03/14/07 – TM & TTC have contract and are ready for kickoff meeting.

02/01/07 – Contracts have been sent to BOCC for confirmation and then we will kick off the project.

01/02/07 – TEI has submitted all contract and bid documents to Facilities for Prosecutor review but as of this date, we have not had any information back on this project.

by TEI TEI

JC – Revolving Door Replacement

12/07/07 – CRA left message with Steve Jones regarding project status.

09/28/07 – CRA approved door shop drawings and sent to Southern Ohio Door. Next meeting to be held prior to start of construction.

09/18/07 – Pre-construction meeting was held. Notice to Proceed issued at pre-construction meeting.

09/04/07 – Signed contract received.

07/26/07 – CRA received minor comments back from Prosecutor. Contracts revised and sent.

06/21/07 – Bids received. CRA directed to accept and issue contracts.

05/29/07 – CRA reissued documents.

05/23/07 – Bids came in higher than estimated. Directed to raise the estimate to $85,000 and re-bid.

02/20/07 – Kick-off meeting was held at 9am in the Courthouse Room B-103. The possibility of installing remote lock down security revolving doors in the south building was discussed.

01/30/07 – CRA received the approved SOSS.

01/16/07 – CRA SOSS was authored and delivered.

by TEI TEI

JUSTICE CENTER ODR Kitchen Upgrade

05/22/09 – TEI has not been paid on this project yet.

03/02/09 – All work completed several months ago.

08/25/08 – Progress is being made on project. No Conf Call today.

08/22/08 – Permit has been issued and give to County (almost 2 weeks ago). Construction is proceeding. MAH is handling the conference calls.

08/07/08 – TEI submitted drawings for permit and copies of permit set to County and Andre for review and their records.

07/30/08 – TEI has been to the site several times getting additional field work for this project. Original load locations do not work for new kitchen equipment loads. 30 days panel load data is required. We will try to get permit on 3 day info.

09/05/08 – Permit obtained without metering. Work is in progress.

by TEI TEI

JUSTICE CENTER Domestic Pump Upgrade

08/24/10 – As-Builts converted and sent to GBBN and Bert.

08/13/10 – Waiting for final payapp and closeout. As-Built dwg is complete.

07/28/10 – Project is complete. TTC to delivery as-built dwgs to Bert.

07/23/10 – TTC is approved O&M’s on these jobs for Bert. Dwgs are being bound. Paperwork looks fine.

05/26/10 – Project is complete. We are awaiting closeout paperwork.

05/13/10 – I approved payapp #2 today and asked Gary for closeout documents.

04/26/10 – Project complete. TEI will release a no item punchlist and begin closeout.

03/19/10 – Corporate came down to investigate a blown fuse situation when opening the cover panel on the new pumps. JCI is looking to hook up the remaining sensors back to the new pump package for Howard. TEI will issue a punchlist on this project.

02/24/10 – QCM pumps have been verified in the shop for payapp. They will start on Monday at 6am and this is confirmed with Howard and Gary.

02/04/10 – QCM scheduled to begin work on 02/22/10. TEI asked for an updated schedule.

01/26/10 – Equipment is ordered and extra pump has been approved.

12/31/09 – Submittals have been approved and sent back to QCM. We still don’t know if Tony has approved the change order for an extra pump.

09/04/09 – Contracts submitted to the County.

08/27/09 – Bid open today. One Addendum was issued.

06/23/09 – Tony said the SOSS has been signed. We are awaiting approval and the PO. The existing Systecon domestic package is variable speed and will most likely be replaced with a new variable speed system that has readily available parts.

06/17/09 – Terry and Mike visited the building to create the SOSS.

by TEI TEI

JUSTICE CENTER – Kitchen Oven Replacement

10/28/10 – Project is complete and ready for closeout. New paperwork has been received.

10/5/10 – CBS will receive $2,150.00 allowance monies instead of the $5,000.00 due to using $2,850.00 for upgrades to both ovens.

09/29/10 – Project complete. Need closeout documents.

08/13/10 – Waiting for closeout docs. Payapps sent in as paid.

07/28/10 – Oven look good. Punchlist items look to be complete. Training should be complete. Ready for closeout.

07/23/10 – Both Ovens are installed. TTC sent an email punchlist to CBS who said they would address these issues on their next visit downtown. Several of the items on the list require adjustment when the ovens are off. They said it would be no problem. Oven #1 had an issue with a door opening switch which CBS fixed and now all seems to be well.

05/26/10 – Kickoff meeting schedule for June.

05/13/10 – We are trying to setup our next const mtg on this job. Equipment should be here soon.

05/05/10 – Ovens are ready for shipping. Project is ready to startup in two weeks.

04/26/10 – Waiting on Oven delivery to begin the project. Contract checked in and is ready when equipment arrives.

02/24/10 – Tony confirmed the submitted oven is approved by Aramark for this project. Other submittal data is acceptable.

01/26/10 – TEI owes CBS submittals on this project.

12/31/09 – Oven submittals have been received. TEI to review and approve, especially the prison option to be finalized.

09/04/09 – Contracts submitted to the County.

08/24/09 – CBS was the low bidder. TEI interviewed CBS and they seem to have everything included in the bid. The bid very aggressive and want to make sure they can sell the ovens when complete.

07/29/09 – This project went out for bid last Monday. Prebid 08/05. Bid Opening set for 08/24.

06/23/09 – Tony said the SOSS has been signed. We are awaiting approval and the PO.

06/17/09 – Terry and Mike made a visit to the building to look over the equipment and the installation. There is some electrical work and mechanical work to perform but the bulk of the cost will definitely be the equipment itself.

by TEI TEI

JUSTICE CENTER Kitchen Hood Suppression Upgrade

10/28/10 – Cintas has not sent in AIA pay applications yet for payment.

09/03/10 – The kitchen passed its final inspection. All contractors were instruction to complete closeout documents and send final payapps in.

08/13/10 – All contractors reported for work. Cintas is compete. DK is complete. QCE did not finish yet. Final fire alarm test will be schedule after QCE is finished.

07/28/10 – Work is ongoing. QCE was on site and hooking up fire alarm contact last night. Cintas was not on site but Ansul system appears to be 90% complete.

07/23/10 – Cintas & QCE electric will begin work next Monday on the systems both Fire & Elec. We are scheduling a fire test pm Friday July 30. We are trying to get DK to fix the Aqua Vent system by then.

05/26/10 – Kickoff meeting schedule for Thursday.

04/26/10 – County reports contracts are ready to be signed. Then we will begin.

03/23/10 – No updates

02/24/10 – Under contract review. QCE called to ask about status. TTC updated them.

02/03/10 – TTC sent BW the prebid mtgs notes for contract approval on this job. We are waiting for construction to begin.

01/23/10 – We continue to wait for approval to proceed with this work.

12/31/09 – All bids have been received. TEI has delivered contracts to Tony for approval. We are waiting on approval to move forward.

09/08/09 – Steve Ossege discussed the hood coverage with Terry Cannon. All hoods that cover grease producing equipment. This means that one hood in the JC kitchen must be covered. The other hoods can have the sprinkler heads removed. All hoods must keep the fusible links in place with the hand pull stations. The Aqua Vent bid is still under review at the prosecutors.

08/24/09/ – Daryl Meadows and Steve Ossege will provide city direction for TEI work. Our design is complete. We just need to bid out the Ansul system replacement.

07/29/09 – Kitchen Hood Fire Suppression Upgrade is almost complete. This project is set for small bidding next week. The cost estimates show the project can be under $25,00 but not by much. The drawings are complete. The specs are complete. We will be inviting fire suppression contact

06/23/09 – Aqua Vent project in contract approval stage. Kitchen Hood project in code review and design stage. Cintas has not emailed or called TTC back about code requirements vs final design of this small project. They have given this information to the County before and I suspect they don’t want to design it for us. TEI has to update the hood suppression system, the interface to the fire alarm panel, and new relays for the remaining equipment for FA signals. This might be after the cost estimates are finished for 2009.

05/22/09 – The Aqua Vent repair has been bid out and we are awarding to Debra Kuempel for about $12k. Contracts are being written. TEI sent out the bid tab and told DK to get hard copy paper work to Tony by 05/29/09.

05/19/09 – Cintas has provided the estimate and a scope for the fire protection work on this project. TTC to use for bidding.

02/02/09 – Terry and Mike are finalizing a small project bid manual for this project this week. We are planning on bidding the 3rd week of February.

10/18/08 – Approved but new hoods not required. What upgrade is left to do? TTC to inquire with Howard.

08/07/08 – SOSS submitted for approval to County.

by TEI TEI

JUSTICE CENTER – Aqua Vent Upgrade

10/28/10 – DK sent final paperwork in for approval yesterday.

09/29/10 – Last 2 change orders were sent to Debra for closeout per Tony’s direction.

08/13/10 – All contractors reported for work. Cintas is compete. DK is complete. QCE did not finish yet. Final fire alarm test will be schedule after QCE is finished.

07/28/10 – DK finished change order work last evening. TTC check it over and was on site. All but one nozzle is working properly. We need a final approval from Howard before testing with City Friday.

07/23/10 – TTC told DK they needed to be done next week with their work. Rick returns Monday for verification. Nozzles have been replaced but 3 still need fixed. TTC will try to have DK fix these when they fix the rest of the equipment on the rinse cycle.

05/20/10 – Sent Tony and Howard final change orders with prices from Debra for wrap up. Awaiting approval.

05/12/10 – Met Debra today on final change order prices. We don’t have everything we need. I will contact Tony with questions on scope of work and change order cost. Terry sent a proposal to Tony and Howard for comment.

04/26/10 – DK sent changes orders for final repair cost to extra items such as the rinse cycle piping and solenoid. TTC and TM to decide next course of action.

03/23/10 – DK reported a list of problems with the Aqua Vent operation. TTC answered several of these answers and asked the County for direction on several others. DK also asked for a sequence of the original Aqua Vents but we don’t have that.

02/24/10 – DK coordinated final work with Howard to be this week and next.

02/03/10 – DK scheduled control panel installation on 02/22/10 when Howard returns. The owner accepted one change order and rejected the extra electrical work change order. DK has not responded to the TEI email about the change orders.

01/26/10 – Aqua vents assemblies have been rebuilt at Debra Kuempel. Control panel parts have been upgraded at DK. Change orders have been submitted for new valves and shock arrestors. We have approved this work. The change order for extra electrical time has not been approved by Tony.

12/31/09 – Rick Kuethe reports the solenoid valves are in and DK will start next week provided their men are approved and Howard Seal gives him the okay to start working.

by TEI TEI

JC – 20 ton Chiller Replacement

07/29/11 – Glen reported there is no turn down on the screw compressors.

06/29/11 – Tonnage looks to be about 10 to 15 tons. Final came in 11.7 tons. TTC sent review to Tony and Glen Wilhoit for their use.

05/05/11 – Waiting on people and load count from County.

04/25/11 – TEI asked Tony and Glen for more information on the space and to mark up the existing dwgs for space and layout, then we can quickly and efficiently run the model for HamCo.

03/30/11 – No further progress yet. Tony said we might write state term contracts for him.

02/22/10 – Tony asked us to do a load calc in the Sheriff area to confirm a new chiller size for the space serving the Sheriff space for additional cooling.

by TEI TEI

Justice Center Kitchen Flooring Study

07/06/10- Attended Kick-off meeting with Bert Watts and Tony Matre.

07/15/10- Submitted SOSS to Bert Watts.

9/27/10- Meeting with GBBN, Fred Orringer (Kitchen Consultant) Jerry Vogel and Andre Tudor to discuss the logistics and operational issues for implementation of the possible flooring replacement.

9/27/10- Transcript of meeting sent to Bert Watts for clarification of scope and summarization of issues discussed in the meeting from the same day.

12/3/10- Bert Watts called to clarify questions on scope. It was determined the report would be issued to the Owner 1/18/11.

1/18/11- Final document was sent to Bert Watts electronically for review.

1/31/11- Bert Watts instructed GBBN to request color and slip resistance texture samples from the vendor for presentation to the Sheriff.

2/3/11- GBBN requested UreCon color samples from Key Resin. Key honored request to send color samples for the epoxy topcoat as well.

2/15/11- Color sample chips arrived along with slip resistance samples.

2/17/11- Color charts arrived.

2/18/11- Bert Watts sent a list of proposed kitchen equipment to be replaced in the flooring project. TEI and GBBN are to investigate possible connection and logistics costs for this equipment for inclusion into the final Study.

2/21/11- Bert Watts called to indicated the final report with modifications is due 2/28/11.

3/3/11- Delivered final copies of Flooring Study to Bert Watts.

6/28/11- Bert Watts requests SOSS from GBBN to include ThermalTech as consultant.

6/29/11- GBBN sent final scope document to ThermalTech for coordination of SOSS scoping.

6/30/11- SOSS completed and delivered to Bert Watts.

by TEI TEI

HC General Structural Services THP Prime

THP completed billing of $5,000 allow for general services. Funds lasted 10 months from July 2010 to April 2011. THP directed to submit new SOSS, refer to F24.

April 15, 2010 THP requested to review possible leak and stone condition of 800 Broadway Bld. THP and County scheduling site visit.

Past Comments

THP completed site review of 800 Broadway wood deck. THP completing analysis and letter report.

THP submitted four names for badges. County review approval policy.

THP reviewed Hillcrest cracked walls. Walls are cracked and further investigation is necessary. Cracks are not a life/safety concern and cracks do not appear to be a warranty issue from settlement.

February 17, 2011 County requested THP review floor condition at Alms Doepke. Work may be part GBBN floor project, if not THP will notify the County on how it thinks we should be proceed.

February 14, 2011 THP reviewed cabin with cracked walls as requested by the County. As a follow-up THP does not think the cracks are a life-safety concern or warranty related to previous work. THP will follow up with a summary letter.

January 27, 2011 County requested partners submit employees names for background check and badges. Future access to County facilities will be limited to people with badges or escorted by badge person.

January 27, 2011 County distributed HC Safety specification section to added to all facility projects. Partners to determine location with front end documents.

January 27, 2011 THP requested to evaluate 800 Broadway interior wood deck structure for loading capacity. THP to evaluate and if work is minimal work will occur within General Service number, if work is more involved THP to discuss separate SOSS.

THP has $3,200 remaining in the hourly allowance.

Septermber 29, 2010 THP to review MDD North side structure for possible work with GBBN.

County approved updated proposal.

County reviewing initial proposal.

THP provided preliminary pricing of the building envelope assessment of Memorial Hall on 8/17/10.

County requested preliminary pricing for building envelope assessment of Memorial Hall building.

County reviewing SOSS.THP Issued SOSS F9A for an additional $5,000 on July 2, 2010.

Roughly $1500 remaining in general fund.

THP used roughly $2,800 of the $5,000 fee.

THP tracking 800 Broadway storage area floor capacity survey to general fund. (2/24/10)

THP tracking additional Masonry work from Hillcrest project to general fund. (2/24/10)

OSS approved 2/3/2010.

THP submitted SOSS on 2/1/2010.

THP needs to submit SOSS.

by TEI TEI

Hillcrest Training School Dining Hall Loading Dock Restoration THP Prime

July 27, ,2011 THP issued AIA Substantial Completion Letter.

July 19, 2011 A&M confirmed (via email) punchlist items completed.

July 12, 2011 A&M submitted final Pay Application for retainage. THP notified A&M that retainage monies will not be approved until all warranty paperwork has been submitted and approved.

July 8, 2011 THP issued (via email) punchlist items to be completed by A&M.

July 8, 2011 loading dock sealant installed.

July 7 & July 27, 2011 THP requested A&M to submit warranty paperwork.

July 7, 2011 new galvanized handrail system installed and THP approved and forwarded to Hamilton County, A&M request for payment, Pay Application No 2 in the amount of $10,070.00.

June 30, 2011 THP issued Change Order Deduct to the contract, in the amount of $3,778.00. The new contract sum is $51,132.00.

June 30, 2011 sealant scheduled to be installed and THP will punch-out the project.

New loading dock handrail system fabrication completed and A&M installed new handrail system week of June 20, 2011.

A&M completed loading dock bumper installation and installed temporary handrail system on May 27, 2011. Hillcrest facility began using loading dock May 30, 2011.

THP issued email on May 24, 2011 for A&M to install temporary handrail system once loading dock bumpers are installed. A&M to be completed with installation of loading dock bumper and temporary hand rail by May 27, 2011, so loading dock may be used by Hillcrest facility on May 30, 2011.

On May 23, 2011, THP approved and forwarded to Hamilton County, A&M request for payment, Pay Application No 1 in the amount of $38,068.24.

Hand Rail steel shop drawings approved (with exceptions noted) for fabrication on May 12, 2011. A&M expects for fabrication and installation of hand rail week of June 13, 2011.

Field Report No 002 issued on May 9, 2011.

Loading dock slab placement took place on May 6, 2011 with THP personnel on site to witness placement. Weather delayed the original April 28, 2011 placement.

New louver and new loading dock bumper approved to be purchased and installed out of project Cash Allowance via Change Order No 1, dated April 20, 2011, in the amount of $1,222.00. Total project budget remains unchanged.

Hamilton County requested price quote from A&M for a new louver and loading dock bumper, instead of reconditioning these existing items as part of the original scope. Pricing came in within the projects original Cash Allowance which Hamilton County chose to accept. THP issued Change Order No 01 on April 20, 2011 reflecting this modification.

The loading dock slab placement is estimated to take place on April 28, 2011. Weather pending.

April 13, 2011 A&M formed and placed concrete for the loading dock steps. THP reviewed the steel and formwork prior to concrete placement.

April 6, 2011 A&M formed and placed concrete for the loading dock walls. THP reviewed steel and formwork prior to concrete placement and issued Field Report No 01 for both concrete placements.

Submittals not complete as of April 25, 2011. Waiting on resubmittal of galvanized steel hand railing for loading dock.

April 1, 2011 A&M completed loading dock demo while maintaining emergency egress.

March 30, 2011 A&M mobilized on site and began project work.

March 23, 2011 THP held a pre-construction meeting on site with Hamilton County and A&M Contracting.

March 23, 2010 preconstruction meeting occurred. Demolition of loading dock to begin March 31 and closure of the dock will last 6 weeks.

February, 22 2011 A&M Contractors passed background checks and work can proceed. Bert has requested a schedule and submittals to be finalized.

January 27, 2011 A&M Contracting had a number of employees not pass the background check. Work cannot proceed until A&M Contracting project team has passed the background check.

Need to determine if project will occur in Spring or this Winter. If work occurs this winter a change order for heating will be required.

Submittals continue to arrive and be returned; last submittal return on December 7, 2010. THP has requested a number of additional items to be submitted.

October 26, 2010 THP returned initial submittals and requested preconstruction meeting.

October 20, 2010 Contract was approved and PO issued.

County reviewing contract for final approval and execution.

September 16, 2010 THP provided updated draft AIA contract per comments from County.

September 2, 2010 THP provided draft AIA contract.

County reviewed project bid.

August 10, 2010 THP recommended A&M Contracting Ink. based on the submitted bid. THP has no experience with A&M Contracting.

Bids received August 5, 2010. A&M Contracting Ink. submitted only bid. Bid was within 10% of published estimate.

Bids are due August 5, 2010.

THP plans to issue final addendum by July 29, 2010.

Pre-bid meeting held July 20, 2010.

THP updated documents and issued documents for bid on July 12, 2010.

County Responded with minor ITB comments THP to update documents.

THP Issued updated ITB documents on June 7, 2010 for review.

THP anticipates finalizing and sending Updated ITB documents to the County on June 1, 2010.

County finalized review of ITB documents and sent comments to THP on May 21, 2010.

THP finalizing bid set; anticipate release 4/5/2010.

THP complete design review meeting with County on 3/12/2010

Project anticipated for summer of 2010. (2/24/10)

Waiting on Owner comments to complete final documents. (2/24/10)

THP submitted owner review documents on 12/21/09

THP provided plans for an emergency deck for loading dock access on October 22, 2009.

THP submitted an SOSS for approval on October 23.

THP completed initial site visit with County and recommended closure of loading dock.

by TEI TEI

Hillcrest Training School Cottage Slab Review THP Prime

Project complete and closed out.
Past Comments

L&T submitted closeout paper work. THP & L&T revising warranty language in submitted warranties; all other paper work acceptable.
L& T completing final project closeout paper work. (2/24/10)
L&T requested a project closeout, so additional grading work will not be accepted.(2/24/10)
All punch list work including additional interior repointing at interior/exterior wall joints reported to be complete. THP to review next week. (2/24/10)
All work complete except punch list. Punch list published and all work except grading except grading expected to be complete by February 19.
Contract was coordinated with the prosecutor’s office and final copies were sent to Hamilton County for review by the commissioners on September 14.
August 17 THP recommended Langenheim.
Bids received August 14 and Langenheim & Thomson was low bidder within the project budget.
Project in bid process with pre-bid meeting completed on July 28, 2009. Bid due August 11th.
Design development documents anticipated to County 6/26.
Partial front end specification sent to County for review on 6/22.
Kick off meeting completed and field survey’s complete.

by TEI TEI

Hillcrest – Security System Upgrade

Project is Complete. All contractors have been paid.

Zeiser:

11/11/04: Zeiser Pay Application #6 to Sharon via Courier.

11/04/04: Zeiser Change Order #4 for Punchlist items 1.19 and 1.20, new door hardware, in the amount of 690.00 typed and courier.

LaForce went to site today to complete punchlist items hardware repairs. However, Chris was unable due to audit, to show them location. Chris could you please coordinate with Zeiser to a date you would be available to show them locations.

Cottages grandmaster key – Contractors are working on resolution to be presented.

Post caps for doors stops are in and ready to be installed.

Indrolect

ICS to complete punchlist including rewiring system. Please update us to the progress.

12-07-04 – New master keys are on site for cottages. Zeiser to coordinate work with Hillcrest

01-04-05 – Zeiser and Laforce to complete cottage level master rekeying. Work to start on 01-05-04.

04-19-05 – Cottage level master rekeying complete. School has accepted suggested solution for fixing the bathroom keys. Laforce has install the proposed fix and is returning to make repairs.

04-2-05 – Bathroom rekey complete however one issue still exists. Dining hall locks have a total of four cylinders and only two where replaced. The second set can be defeted but according to Dave Jones with LaForce this is not possible because of the type of lock. They proposed replacing the sets and provided pricing accordingly. County review cost and felt that it was not their responsibility because this proposed fix was only a compromise that remedy the situation. Upon conference call with Diana, Mike, Jeff, Chris and Dave fix was agreed upon. Dave with LaForce to complete week by 05/06/05.

by TEI TEI

Hillcrest – Modular Building / Addition

08/08/05- C+R sent a preliminary project report to Hillcrest. TE to review and comment and provide updates to Spencer for inclusion.

06/16/05 – C+R, TE and HCDF had a follow up meeting at site to discuss layout issues. Spencer distributed possible floor plans for review. TE went over utilities that were underground and reported that electric, gas and water are all near the expansion area.

01/19/05 – Second Field trip to discuss cost and scope.

by TEI TEI

Hillcrest – Fire Alarm Replacement

12/26/06 – Project has been archived by TTC.

11/09/06 – Final permit was accepted

11/03/06 – TEI is working on closeout documentation. Final devices have been installed. Final approval has been obtained from Hamilton County Building Department.

10/27/06 – Final devices have been installed. Ginter to set up to get final approval

09/29/06 – Waiting on approval of revision from Hamilton County Building Department

09/22/06 – We have passed and approved (signed off and everything!) on 21 buildings out of the 22. Each building was permitted separately. The one building (grounds keeping) is on hold until the inspector sees the 501 model number heat detectors installed in the or we submit an engineering change that corrects the permit drawings with the 503 model number. They are in the process of getting the paper work together and will forward it to you. As far as the grounds building we are not being told that we can not use the building. Also we don’t believe they occupy it all day, mainly just storage and prep for the grounds care equipment.

9/12/06 – 21 buildings passed fire alarm inspection grounds keeping building approval on hold until heat detector issue is resolved.

9/6/06 – IPS adjust strobes in cottages that were coming back with a trouble condition. They were set to the wrong setting. Ginter adjusted devices.
8/30/06 – Hillcrest performed a drill test and a few troubles came back on the system. Chris and I contact IPS and Ginter to get trouble condition resolved.
8/25/06 – Hillcrest will be unable to have the cottage door delay egress maglocks repaired prior to the 8/29 test date. Test is rescheduled for 9/12
8/21/06 – Complete rewiring of audible devices in cottages so the do not sound when fire alarm system is in alarm.

04/06/06 – Mike Stiles has requested a change in the meeting schedule for 04/10 as he is out of the office on that day.

03/06/06 – Ginter email TEI that the county was requiring separate permits for each building. TEI call the Building department and clarified the situation. The county required 22 permit applications and 6 drawings. The minimum fee per permit is 257 dollars. Ginter indicated that they did not have an amount this large covered in there bid. TEI requested that Ginter submit a total for all cost so far in the project: Some of the additional fire alarm cable, additional permit fees and deletion of 10 in-ground boxes.

03/03/06 – Ginter Electric provided submittals, TEI approved submittals.

02/28/06 – Construction meeting #1, Please see attached.

02/03/06 – Ginter submitted pay application #1, TEI approved.

01/20/06 – Preconstruction at Hillcrest.

12/16/05 – County reported that this project has been approved. Wayne Trent will be project manager on this job and therefore the project is being moved to his task sheet.

02/01/06 – Ginter Electric started field investigation.

12/01/05 – Contract was approved.

11/01/05 – TEI prepared contract for Ginter Electric, Alternate #5 was accepted.

10/31/05 – Bid opening held, apparent low bidder was Ginter Electric.

10/24/05 – Issue addendum which delayed bid opening to October 31st. Addendum was the same form Re-bid.

10/14/05 – The rebid II prebid was held. No contractors showed up for the rebid. This might be expected because everyone has already been on site for 2 prebids already on this job. There can still be and will most likely be bidders on this project.

10/13/05 – The rebid has been rejected due to bond issues again and estimate tolerances. The second rebid is on the street with a revised bid opening of 10/25/05. MS has all the revised documents on the street. The prebid is 10/14/05 at 10a.m. MS will attend for TE.

10/11/05 – Project was issued for rebid.

10/05/05 – Bid was rejected by the prosecutors for basebid plus alternate because the awarding contractor complete the bid bond for only the base bid amount. TEI advise that alternate 7, which was cabling in the ground in conduit plus a spare, was very important. Since the contractor filled in the bid bond purchasing would only allow base bid. The next lowest contractor was reviewed and but there base bid plus the alternate was greater than 10% of the advertise amount. The county decided to rebid the project.

9/02/05 – Contract completed for AC Electric, base bid plus alternate 7. Submitted to Diana for approval.

08/25/05 – Bid opening held, apparent low bidder was AC Electric.

08/10/05 – Project prebid was conducted.

08/3/05 – Project was issued for Re-Bid.

07/01/05 – Diana, Gerald and TEI decided to hold off on rebiding project to gain more interest. MS contacted several contractors to find out marked conditions. The consensus was that the school at released a lot of work to be complete by the beginning of school and the contractors where very busy. It was also determined that breaking the project into several alternates would be required to make sure the project can be complete within budget.

06/28/05 – TTC and MS discussed the bids. A clarification on the addendum split up the labor versus the equipment. The bid is not over the original estimate of 298,000 however it is over the revised estimate. TE has to discuss with Diana when she returns our call. MS believes that this project should be rebid as the bids are simply too high for this type of project.

06/24/05 – Diana emailed TE today to say that the bids came in over 10% of the budget. TTC and MS to discuss with DES on Monday in a phone call.

05/19/05 – I spoke with Andy Boyd at IPS about the project. I discussed the three state term options with him which are: A) Purchasing material only and bidding out the installation to an electrical. IPS (or competitor) would be subcontractor.B) Purchasing material and IPS to set up the system and bid out electrical labor for the installation.C) Purchasing the entire project from IPS.Option C has changed because Purchasing indicated that they can only do the project if they do not use any subcontractors. I presented this to IPS and they gracefully turn this option down because they do not have the resources to complete a project of this size on there own. Option B was discussed. I ask the question, what advantage is it to the county to proceed. IPS indicated the main advantage was for them and that the cost under state term and bid day would be about the same. IPS and I agreed that this was not the best option for the county. Option A was discussed and believed to be the best solution. This makes the bidding process cleaner and only has one contractor to manage. The above is a recommendation but if you have any comments or suggestions please let me know. I would be happy to proceed the manner you prefer.

05/12/05 – TEI contact JoAnn with purchasing to discuss doing the project under state term. She indicated that IPS could not have an electrical subcontractor because the amount of the contract was to much and must be bid out publicly. Prepurchasing material was acceptable.

05/10/05 – TEI presented final drawings and Edwards material pricing. TEI indicated that the material price provided by Edwards was in budget with all the options requested but not with alternates. TEI indicated that Alternates should remain to insure the budget. TEI indicated that there where 3 way to bid the project: 1) Prepurchase material and bid out the installation. 2) Prepurchase material and ESI setup of system, bid out electrical labor to install system. 3) Purchase entire project under state term. Hillcrest indicated that they where interested in purchasing the entire project though state term because of the difficulty of trenching on the site and the responsibility to the electrical contractor.04/18/05 – TEI present near complete drawings written around a full system replacement of any manufacturer. Hillcrest review vendor options and indicated that the two favorites where Simplex and Edwards. Overall Edwards offered a security solution that could handle the need of the fire alarm system and replace the antiquated access control system. State term options was discussed and Hillcrest indicated the we receive compare the price for Edwards with almost all requested options and compare to original budget. Items to be included is: Voice System, Access control panel for each cottage, Access control software, Data base software, Computer based graphics package. Alternates for the project shall be: Voice system in lue of Tone system, Additional access control module for each cottage, Access control module for gates and upgrade spare conduit in ground with honeycomb type.3/30/05 – Vendor presentations.

03/29/05 – Vendor presentations.

03/09/05 – TEI complete field take off and composed 4 options with associated cost estimates to proceed with project. Cost estimate #1 was to completely remove existing system and replace with a new tone system of any manufacturer. Cost estimate #2 was to completely remove existing system and replace with new voice system of any manufacturer. Cost estimate #3 was to upgrade existing system Fire Alarm Control Panel with the same manufacturer, add additional devices as required per NFPA and replace all existing wiring. Cost estimate #4 was to upgrade the Fire Alarm Control Panel with the same manufacturer, add additional devices as required per NFPA and replace some of the existing wiring. Hillcrest to review options and requested vendor presentations be conducted to aid the school in there decision. TEI to continue with drawings as far as possible until decision is complete.

02/22/05 – Project Kick Off meeting. TEI to start field take off and review options. Original cost estimate presented was for a complete system replacement. TEI suggested reviewing the existing system and look at possible options.

by TEI TEI

HILLCREST – Electrical Upgrade Phase II

03/29/07 – TTE received final pay application and the ‘CONSENT OF SURETY COMPANY TO FINAL PAYMENT’ form from QC Electric. All of the final documentation will be forwarded to Hamilton County Facilities the second week in April of 2007. There are no open issues on this project.

03/23/07 – Closeout meeting was held on site today. There are no open issues for this project. Final documentation (O&M’s, As-Built, Final Pay App., Close Out Documentation) have been turned over to TTE. An Allowance Clarification letter needs to be written and signed before the final pay application can be processed. This letter will be sent to Queen City on Monday, March 26, 2007. When TE receives the signed letter back from QC all of

these materials will be turned over to Hamilton County Facilities. Dick Kuhn, Mark Donnelly, Bert Watts, Chris Hohmeister, Phil Lambing and Nick Richmond attended this meeting.

03/20/07 – Closeout meeting is scheduled for this Friday. Project is paid out except for retainage.

03/15/07 – I believe this project is ready for closeout. TTC to confer with PL and wrap this project up.

02/19/07 – Project is complete. TEI is starting closeout phase on this project. TEI will coordinate with Mike Jackson on closeout on this project.

11/10/06 – Concrete pads have been poured. ID studies and receptacle installations have been started in cottages. Project is still on schedule for December 1 shutdown.

10/27/06 – Contracts released, and preconstruction meeting was held. The schedule was approved. Coordination meeting with Duke Energy occurred 10/25/06.

09/22/06 – Still waiting for contracts.

09/15/06 – Bert called back to say contracts have not been received yet. Contracts are in review stage. QCE will not be on site next week but TE will coordinate a const mtg as soon as the paperwork is in place.

09/15/06 – Construction to start next week.

09/15/06 – Bert Watts said Contracts were released, ThermalTech has not yet received them.

09/08/06 – County reviewing contracts and insurance.

by TEI TEI

HILLCREST – SECURITY BUILDING

01/28/08 – This project was added late because it started before the website. After 4 years of futility, the County and ThermalTech finally agreed on a solution to the cooling and heating problems in the Security Building. Tony bought the new computer room ductless split system and County trades installed the unit. It is currently up and running and the space is doing great! EDI came to the site on Monday, Tuesday & Wednesday to install the AC unit. We had some unit problems as we received an AC unit and not a heat pump. Jeff did a great job returning the unit and getting the heat pump for us which amazingly was in stock at Habegger. We installed a 4-ton fan unit and 3.5 ton heat pump so that I could ensure full flow into the building since the ductwork sizing has always been a problem. The new unit and heat pack are up and running from the Carrier thermostat on the wall and the space has been 72, 74 and 77F when called to check on. Mike Jackson and I discussed the cold feet issues today and he is comfortable with the unit providing enough heat and said the space was 1000% better. I will recommend some radiant heating solutions to Hillcrest for their feet but I don’t believe that TE should have to pay for those. I believe this project could finally be complete.

by TEI TEI

HILLCREST – ELECTRIC PANEL UPGRADE PHASE III

04/23/08 – QCE had to fix the same AHU unit again when they cleaned up some of the school panel wiring. Project is nearing completion.

04/11/08 – MAH released the Const Mtg 05 Notes to the project partners.

03/13/08 – (3) full size sets of drawings and a letter stating that no loads were being added and that the feeders and panels are of sufficient size were delivered to IBI today. QCE should be able to obtain a permit with these actions.

03/06/08 – Today’s construction meeting has been rescheduled to 03/17 at 2pm. We also discussed the revised schedule for the project which I updated in a pseudo meeting minute memo in PDF format. TEI will process the IBI letter for no load increases during this job for QCE to give to IBI to get permit squared away. Bert also asked for a revised scheduled update which we have included in the memo.

02/26/08 – QCE called today to tell us that IBI denied our permits for this job. They now instead want TEI to provide a letter saying no loads have been changed or will change during this project and we are doing a direct panel for panel replacement. A shutdown is going to be scheduled in late March for school panelboard replacement(22 or 23).

01/09/08 – Today we decided that the pre const meeting notes were correct and the project completion date for this project is March 1, 2008. QCE will do and update project schedule with delivery details for TEI to approve and give to the owner very soon.

01/07/08 – MAH released the Const Mtg 02 Notes to the project partners.

11/28/07 – QCE provided an update to Mark Donnelly today with the following information: The ID portion of the project is going well. The (4) cottages are complete. Chapel, John Morrow building, Goundskeeping building, & Pole Barn are completed also with the ID. We still have to change locks and install the new locks on certain panels. Administration building is approximately 75% completed, the School is approximately 50% completed. Some work has been done in the Dining building, Rec building, & Gym. We have not been in Security yet. The transformer pad for the school transformer was poured 11/27/07. All Panels have been measured & verified. The order has been placed, but I don’t have delivery dates yet. As soon as I have the delivery dates I will put a schedule together and ask for a meeting to go thru what has to happen for the shutdowns & installation of the panels. TTC setup a const mtg for 12/10 at 2:30pm to follow up with panel order and schedule.

11/12/07 – TTC released the Pre Const Mtg Notes to the project partners.

08/31/07 – TEI still awaiting word from HCFD. This project remains in the Contract Phase.

08/17/07 – TEI awaiting word from HCFD in order to move forward with this project.

08/10/07 – TEI spoke with QCE about their bid. They were comfortable with the bid so TEI wrote the contract and forwarded to the Hamilton County Facilities Department. TEI awaiting word on signed contract so this project can move into the construction phase.

07/31/07 – TEI has obtained a copy of the bid package and is reviewing the information. Dick Kuhn of Queen City put together the bid; he will be back in his office on Wednesday, August 1. NR will contact him in the morning to discuss the bid.

07/27/07 – Today was the bid opening. NR attended for TEI. QCE and Ginter are the low bidders. QCE was at $138,000 for this project. TEI will check the bid to make sure everything is covered.

06/29/07 – No change.

06/22/07 – TEI obtained an ITB #099-07. This project is set to be issued on June 26, 2007. The pre-bid meeting is scheduled for July 3, 2007.

06/15/07 – TEI conducted a site visit on June 13, 2007. The final design strategies have been discussed and this project is moving into its final stages before being issued for bid. TE plans to issue this project for bid by next Friday, June 22 of 2007.

06/08/07 – This project was reviewed in a meeting between TEI and Mark Donnelly. There will be a set of drawings and a cost estimate delivered to Hamilton County Facilities Department this afternoon (June 8, 2007). A meeting will be held on site next week to review final conditions and then will go out for bid.

06/01/07 – NR will have a review set of drawings, specifications and a cost estimate ready for Mark Donnelly to review by the end of next week (June 8, 2007).

05/25/07 – Project is moving along as expected.

05/22/07 – Project review meeting was held today. Various points were discussed as described in the meeting notes. TEI is moving forward with this project while keeping the points discussed in today’s meeting in mind.

05/21/07 – TE has completed all field work. Drawings are 98% complete. A meeting to discuss the status of this project is scheduled for May 22, 2007 with Mark Donnelly.

04/07/07 – TEI has completed the device identification portion of this project and are currently inputting the data into CAD. The panel replacement portion of this project is moving along as expected.

04/20/07 – TEI has been on site working on the device identification portion of the project. The only areas left to complete are north and south wings of the School Building. TEI will be on site today, April 20 2007, to work on the panel replacement portion of this project.

04/13/07 – TEI will be on site beginning next week to work on the electrical device identification portion of this project. Design for panel replacement is moving along as expected.

04/06/07 – Project moving along as expected.

03/28/07 – Kick-off meeting held on site. Buildings to be included in the Device Identification portion of this phase are Administration, School, Chapel, Security, Maintenance, Storage, Dining & Recreation and Gym & Pool. Efforts are aiming for August of 2007 for the shut-down and transformer pad replacement to coincide with the week that school is out of session. Attendees of this meeting include: Mark Donnelly, Bert Watts, Chris Hohmeister, Phil Lambing, Nick Richmond and Ron from Facilities. The next meeting is scheduled for May 2nd at 10:00 on location.

03/23/07 – Design meeting to be held the last week of March 07.

03/20/07 – TE to assist in deciding scope on this project and settle on which buildings are left to do on this site.

03/15/07 – SOSS received and posted to the website. A kickoff meeting will soon be scheduled.

03/14-07 – NR created project journal and filled in back data from previous meetings. Project awaiting signed SOSS to be returned to TE.

by TEI TEI

HILLCREST – KITCHEN HOOD UPGRADE

03/13/08 – TEI awaiting scheduling of suppression system test. Will witness on request.

03/09/08 – TTC sent closeout paperwork to DK and request for Credit on project.

03/05/08 – DK tested the electrical portion of the system successfully today. All relays operated as required. A test of the fire suppression system will now be scheduled by FES. This will need to be done when no cooking is occurring under the hood. A tank of ordinary air will be connected to simulate the suppressant.

02/26/08 – DK reported that this project is complete and ready for final fire alarm test. This is being coordinated through Hillcrest and MJ. TE may attend this final fire alarm test.

01/28/08 – We have received Payapp 2, O&M’s and an approved permit for this job from DK. MJ requested and TTC asked DK to setup a FA test showing the working components of the system. MH has the O&M’s to approve for this job.

01/08/07 – Earl Moon approved the project yesterday and TTC spoke to Kevin Broxterman today about the allowance usage and remaining allowance monies on this job. TEI will schedule a punchlist visit to the site and close out this project. TTC tried to contact Huntwork to do the punchlist today while he is on site with another job. TTC instructed DK to do as-builts and give them back to TEI for inclusion into the CAFM system. TEI will send an email outlining all the closeout procedure documents for this job. MD has requested closeout before next Monday. TEI finished the punchlist today and distributed this to everyone.

12/12/07 – MAH talked to Kevin Broxterman and the disconnects are scheduled to be installed on Friday. He asked if there had been any discussion about his using some of the allowance for this work. I directed him to Bert Watts while Mark D is out.

11/30/07 – The electrical inspector has an issue with the disconnect location because of access to the them over the equipment. MH went to the site today to review and find a new location based on our discussions with Kevin B at DK and Mark D at the County. We want to find a new location on the first floor for the disconnects. MH located a space behind the wall of the existing equipment. We will have to relocated the existing fire extinguishers, sign and fire blanket so everything can share the space but it appears very do-able. TEI is sketching up the photo right now for changes. The inspectors wanted the disconnects in the basement but the owner does not. DK submitted their first pay app on the job and TTC approved it for $8941.60.

11/13/07 – Debra-Kuempel has reported updates as well as Mike Jackson from Facilities. Debra says they should finish tie-ins this Thursday. Greg will schedule with IBI and IPS for inspections and fire alarm tie-ins for the end of this week or early next week. Greg of DK reported that they had no problems or questions at this time. MJ reported that the tie-in in the basement Rec Ares is complete. Also that the new disconnects for the steamer and the oven are complete. DK scheduled a shutdown last Friday on 11/09/07 and all went well. DK has not been above the ceiling in the crawl space yet nor are they on site on today. To TEI, this project seems to be going well. TTC will check progress with MH at TEI to make sure we are on target for this job and possibly schedule a field trip this week to check on progress. This field trip may be combined with meeting EDI at the site to discuss the AC installation and also for TEI to review the computer room installation that Facilities is currently working on.

11/12/07 – TTC checked in with Kevin Broxterman. Equipment is ordered and DK will be on site this week not last week to do the work per the schedule agreed to with County Facilities.

10/12/07 – Pre Const Mtg held and meeting notes distributed to all attendees. Super timeline crunch has been averted by Facilities passing the accreditation testing with the understanding that this hood system is being upgraded right now. DK will order parts and scheduled work very soon.

08/13/07 – TTC interviewed Debra Kuempel and confirmed their bid price was good. TEI wrote the contact and TTC marked it up. It will be in County hands on Tuesday for approval. The PO can be acquired at the same time. TEI will apply for the IBI permit on Tuesday as well.

08/10/07 – This project was sent out to bid on August 2, 2007. A mandatory pre-bid meeting was held on site on August 3, 2007. This project opens for bid on September 3, 2007 and is to be completed by September 15, 2007.

07/27/07 – This week TTC & MH began working on the final design this week. MH visited the site for fieldwork and created the construction drawings. TTC is hoping that

05/14/07 – SOSS has been received and logged. TEI asked for a kickoff meeting around 05/28-05/31. We will await MD direction on this.

04/27/07 – TEI awaiting signed SOSS.

04/17/07 – MJ called to say that fan belts were replaced and airflow in the hoods is up significantly. One motor has been replaced previously. TTC is writing the SOSS for the report.

04/13/07 – TTC & TA met MD & BW at the site for a walk through and looked at existing conditions of the kitchen hood and electrical requirements. TTC found that the EF did not have any nameplate date on them. Airflow to hoods seemed low. MD requested traversing the hoods to confirm airflow.

by TEI TEI

HILLCREST – Oak Cottage Wall Investigation and Hillcrest Maple Cottage Wall Investigation

3/29/12 – Langenheim & Thomson install block walls per THP details.

County reviewed THP proposed test of replacing block wall with reinforced stud wall system. Due to security concern, wall is to remain a block wall system. THP to complete test installation of block wall system with control joints.

3/29/12 – THP issued preliminary summary letter and will follow-up with detailed letter report.

3/30/12 – THP issued final report with all test repairs complete.

by TEI TEI

HILLCREST – CT Cabinets Upgrade

09/28/10 – SOSS received. Design documents are complete and ready to be sent to County for review. TEI QC complete. Should be bid next week.

10/28/10 – TTC has to release for bidding. The job is ready to bid.

01/24/11 – DeBra Kuempel won the bid and we turned in a contract for approval. I am not sure of the current status. Confirm with Bert when the pre const mtg should be scheduled. DK should have everything turned in for approval.

03/30/11 – The contract is either approved or close to approval so we can start the work very soon.

04/27/11 – Bert sent the contract to Debra Kuempel and let them know that background checks should be gotten and given to Bert to start the project.

05/06/11 – Bert asked about background checks for DK.

06/29/11 – Kickoff meeting was held Monday. TEI is authoring the meeting notes. Work scheduled to be done late July or early August.

07/28/11 – ThermalTech returned all DK submittals on the project. No exceptions.

08/22/11 – Work scheduled for early September. DK sent a preliminary timeline for Bert to share with contractor. Power permit is being coordinated with Duke.

02/28/12 – ThermalTech is in receipt of the closeout documents that you have submitted for this project, along with a request for final payment. Per the requirements of the project manual, there are items that must be submitted prior to processing the final payments. Upon receipt of the following items, closeout documents can be processed: Notarized statements of warrantee, AIA Document G706 contactors affidavit of payment of debts and claims, AIA Document G706 A Contractor’s affidavit of release of leans, and AIA Document G707 consent of surety to final payment.

6/26/12 – Mike Barth spoke with Tom Helmes of DK, indicating that the Notorized warranty is the only item that is holding up the closeout package. Tom indicated that he would drop off the warranty on 6/26 for final processing. Upon receipt, ThermalTech will process the closeout documentation.

by TEI TEI

HILLCREST – Underground Fuel Tank Replacement

09/28/10 – SOSS received and project out for bidding. Pre bid is Oct 4.

01/24/11 – Pre Const Mtg schedule for next week with Valor Construction and Indrolect electrical company. The County asked us to finalize a temporary design for power to the maintenance building that broke during the winter.10/28/10 – Bids are opened. Valor has passed their pre award meeting questionnaire. Contract has been sent to Tony for approval.

02/22/11 – Submittals are approved for the tank and sent back to Valor for construction and ordering. Indrolect submittals are approved and sent back. Indrolect has additional background checks because they are going to bore the new conduit under the driveway on this job. Valor has to come to HamCo to do their backgrounds checks. We notified Valor by email of this requirement.

03/30/11 – Contractors are on site and beginning work the first week of April.

04/27/11 – The old tank is out. The new tank is in. Valor ordered the wrong sensors so they have been reordered. IBI rejected the disconnect install so that has been reworked and TEI will pay for that and submit a change order for the all the engineering changes on this project under the new contract for Tony to approve. If everything goes well Indrolect and Veeder-root rep should have the Veeder-root and TopKat working on Thursday. Valor will arrive on Friday to put the interstitial sensor in. On Monday, Indrolect will arrange the final inspection and by next Tuesday everything should be complete.

05/02/11 – Valor has completed the tank install. All sensors are installed and wired. Indrolect is complete with all wiring and hookups but will be back to backfill the grounds this week. We have IBI permit tag in place. Reliable is on site programming the TopKat with Denton. Theoretically, the system is ready to use if needed. I believe all permits and inspections are complete. Now it is down to getting Veeder-root programming complete and making the TopKat communicate with the pump properly and that has to be coordinated with Denton and Reliable I believe.

05/09/11 – The new software is on it’s way and Gene from Reliable is on site working on comm issues. TTC spoke with Denton this morning.

05/13/11 – Michael Denton asked if any word on the software for the ball vent yet. TTC replied that Valor had been asked how to remove the ball valve with no response, but the software was on it’s way and should be on his computer before the 12 day test runs out IMO.

05/25/11 – Valor must complete punch list items and train on software.

05/31/11 – Valor will finish the punch list on the tank next week and get factory trained representation training for Denton. He promises to get this work done ASAP. Closeout documents are in process and may be at ThermalTech this week.

06/14/11 – Indrolect instructed to use May 17, 2011 for date on final paperwork. Substantial completion will be typed up by LKV.

06/16/11 – TTC asked Liam for an update on the following punch list items: Has the lip at tank fill been fixed? TEI will apply laminated tags for Valor if they can get the rest of the project completed. The ball valve has been removed. Factory training on the TopKat system needs to be done, asked if Gasboy will send someone. Local rep says he will get everything up and running.

06/21/11 – TTC and Liam spoke and the closeout packet is coming to TEI for final closeout. Liam believes the punch list was left with Denton and is complete except for factory training which Liam is currently working on. Denton had said that the only thing they thought Valor did last week was remove the ball valve, but Liam believes that Loren completed everything but training. Some confusion on who has the punch list and how much of it has been checked off completed by vendor and witnessed by Owner. If everything is complete, TTC will make a punch list visit and sign off to completion and send Tony all that documents that Valor is sending to TTC for approval.

06/29/11 – We contacted Valor again this week to try to get them to rectify the punch list issues.

07/05/11 – Valor called back today and said they would be down this week for fix the water leaking into the fill sump issue.

07/11/11 – TTC contacted Valor and Indrolect about closeout documentation.

07/16/11 – TTC Received a new payapp from Indrolect. Signed and sent to Bert Watts with closeout documents.

07/29/11 – Valor has completed punchlist work on the tank. We still need that extra training from the factory. Some water got into a fill sump again Denton noted. ThermalTech has to investigated.

08/08/11 – Valor payapp #2 for 15690 arrived but TEI believes the county will want to put this on hold for now, until water issues and training are complete. Tony Matre stated to hold the invoice until items are resolved.

08/22/11 – Uploaded Closeout Documents for Tony. PDF

08/30/11 – Still having some issues at Hillcrest with water in the fill and the training portion of the TopKat System. TTC would like Valor to check with Mike Denton to discuss these issues and a timetable to rectify. TTC should be copied on all correspondences.

08/31/11 – TTC has sent photos of the water issue to Christ Batch at Valor. Chris will also check into the training that was scheduled for 8/18 and they thought was completed. Chris will get back with TTC soon to discuss options.

09/08/11 – New OPW spill bucket has been ordered by Valor, should solve the water issue.

09/20/11 – Valor has replaced Emco bucket with O.P.W. The O.P.W. bucket lid is much heavier and does have a gasket. It is a nice install. The plastic was removed from the new concrete and the lid opened. The inside of the bucket was damp, but this could be because the work was done in the rain.

09/21/11 – Mike Denton says that the factory authorized training has not happened yet. It was scheduled but was never followed through with. This training is on how to use the TopKat software and reporting. The initial installer had never worked on the system before. TTC has asked Valor to please check on this since it is the only open item remaining. The sump pump was dry as of this morning.

09/27/11 – The supplier e-mailed Valor Contracting and stated that the training issue would be resolved within 24 hours.

10/05/11 – TopKat has set training up for October 6th between 9:00am-10:00am through Chuck Storey.

10/06/11 – Training by TopKat was completed today.

by TEI TEI

Hamilton County EMA

THP work complete. Refer to TTE for project status.

December 3, 2010 THP provided a summary review letter of siren pole wind loading review.

November 18, 2010 Thermal Tech requested siren pole wind load review. Work to proceed hourly.

by TEI TEI

Courthouse 2010 Lighting Repairs

Refer to TTE for final closeout status.

March 7, 2011 THP published review of completed stone work.

Work complete in late February.

Once approved, THP will close out the project.

March 3, 2011 THP completed review of work and sealant work is acceptable. THP did not see the installation, but received written confirmation that backer rod and shims are installed in the joint.

January 27, 2011 County approved contract. Ram reviewing contract with corporate office. Ram insurance paperwork submitted and approved.

December 28, 2010 Preconstruction meeting occurred.

November 4, 2010 bid received. Thermal Tech and County processing bid.

Prebid meeting completed October 28, 2010.

Prebid meeting scheduled for October 28, 2010.

October 7, 2010 THP sent draft documents to TTE for review.

THP anticipates returning the documents October 7, 2010.

September 28, 2010 THP received draft documents for THP review and design of sealant repairs.

September 2, 2010 provide an email proposal to Thermal Tech to provide consulting services.

August 26, 2010 County, Thermal Tech and THP had kick off meeting to review lighting repair project. THP requested to provide assistance with sealant joint repairs. THP to be a consultant to Thermal Tech.

by TEI TEI

Courthouse 2010 Service Drive Water Infiltration Investigation

Current Comments (9/292010)
Project updates included with F12.

August 26, 2010 Tony Matre requested option #2 repair (full repair) be completed. To expedite 2010 repair schedule work to be included with F12 as an add alternate. THP to bill design services to remaining hours on F13; then review additional service proposal.

THP delivered report August 25, 2010.

Past Comments

THP delivered report August 26, 2010

THP to complete report.

SOSS approved June 8, 2010.

THP submitted SOSS for review on May 11, 2010.

THP and County completed kick-off meeting.

by TEI TEI

Courthouse – Exterior Lighting Project

04/11/05 – TE still trying to get final paperwork from Mike Cavanaugh. Once we receive this final paperwork, we will send to Mark Hardin.

04/11/05 – ICS completed As-builts and sent to ThermalTech and TE sent to MWH in middle march.

04/15/05 – Received final pay application and closeout documents. Signed and sent to Hamco.

by TEI TEI

Courhouse – Video Surveillance Project

07/12/05 – Stapleton turned in final pay application, change order and final paperwork. TTC will review this week and get down to the County.

12-20-04 – Construction meeting #2 see meeting minutes.

12-02-04 – Construction meeting #1 see meeting minutes.

11-0-04 – Pre construction meeting took place, reviewed standard county items.

10-05-04 – AIA Contract for Stapleton typed and mailed to Sharon Chamberlin.

Project bid open had taken place. Mike Stiles has contacted Stapleton Rep. and confirmed the package and Sharon has authorized the contract to be prepared. Mike Stiles requested a copy of Stapleton’s bid from JoAnn in purchasing. Waiting for her to let us know when its available for pick up.

01-20-05 – Construction Meeting – see meeting notes below.

03-07-05 – Construction Meeting – see meeting notes below.

04-08-05 – Completed punch list and sent to contractors to complete – see list below.

04-29-05 – Received call from Paul Wilson with CSI. He indicated that the items on the punch list are complete. He also indicated that their time exceeded the bid amount for setting up the NAS device and switch. He request a change order for the additional time. TEI indicated to submit the additional hours and money requested for review by HC and TEI.

by TEI TEI

Courthouse Archive Room Renovation

08/08/05 – TTC received and approved the final paperwork for CH Archive from Ginter. Warranty and other paperwork received. TTC will hand deliver on 085/09.

08/01/05 – Diana called today because only 2 of the closeout documents have been received. TTC is checking with Ginter.

07/11/05 – TE received Ginter’s final paperwork for this job. TTC will review and sent to County.

07/07/05 – TE and PHB worked out the payapp details today by phone. TE had payapp #4 in house but wasn’t logged into the website. Issues are now resolved. Final paperwork will be sent to the county on Friday.

07/05/05 – TE received final pay application for project. There are some details that TE has to discuss with PHB before TTC will submit to County. Owner’s address is incorrect and has been on previous apps – TTC missed that. The previous payments don’t add up. Line 7 (previous payapps) doesn’t match Line 6 on previous payapp. TTC will stop in at PHB in morning and go over with Lee Anne.

06/21/05 – PHB (LeAnn) called to say they are sending a revised final payapp to closeout the project. TTC coordinate final numbers and instructed the to apply for final.

06/15/05 – Dorn Sprinkler Invoice arrived for $220.00 to clean the sprinkler head covered with insulation. TE will mail to Diana today.

06/10/05 – Peck sent a Word document confirming final allowance usages. TE instructed them to sign the change order sent and send back all paperwork. Bernie said he is on it. TTC emailed Dorn Sprinkler about invoice for cleaning the sprinkler head.

05/24/05 – TTC and PHB discussed final paperwork and O&M’s. PHB is finalizing project for closeout. Paperwork should be received in a week or so.

05/11/05 – DES advised to closeout the project and have PHB return the balance of the allowance. TE will write the credit change order for everyone to sign of $1358 and send for everyone to sign. TE also updated the allowance usages for this project below.

04/26/05 – Ginter and Honeywell ran the FA test this morning. Everything passed with flying colors. Larry Keifer asked us to replace one sprinkler head that had some spray foam insulation on it from the ceiling work by JB Schmitt. TTC called Dorn and instructed them to get with Mike Bartlett and take care of the issue ASAP. TTC told Ginter to send in all their final paperwork for the job for approval.

04/22/05 – TTC contacted Bernie about final paperwork and O&M’s for this job and finalizing allowance work too. Bernie said that the would work on it right away.

04/18/05 – Ginter called today to inform us that the city has finalized our final inspection date of April 26 at 6:00 a.m. at the site. JC Schmitt does not believe they need to be there so TE, Ginter, Honeywell and Mike Bartlett will be required to run this test for approval. All the work is already complete and the pretest already passed. Please make arrangements for this test and I believe Jim Ginter already called to let you know about this test.

04/11/05 – TTC called city directly today to inquire about status of permit. It has been two months since we submitted engineering change to this project and TE pointed out that we need get this project finalized. TTC is waiting to hear back from city. City mailed to Jim Ginter on 04/08/05. We need this permit to call for final inspection according to Amit. Jim did you receive the permit in the mail yet?

04/05/05 – Permit Number for Engineering Change to Fire Protection: 2004P07997.

04/0405 – Ginter called and got approval on the permit from Amit Gosh. Once more approval or actually waiting for the city to call us to tell us to pick up the permit – we will be ready to go.

03/24/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/16/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

03/03/05 – TTC checked with Ginter today on status. No new information to report – City has permit and is reviewing it.

02/22/05 – Ginter reported that all new fire alarm devices are installed in the Archive Room. Joe Dale of Honeywell is working on the new Fire Alarm permit drawings which have to be submitted to the city as an engineering change to the existing permit drawings. Mike Stiles will assist Ginter if needed to submit these plans. Honeywell was supposed to turn in these drawings on Mon or Tues of this week to the city. They are not turned in as of this email and Jim Ginter is calling Honeywell with word that we want them turned in ASAP so that we can pay out the GC in two weeks and have final inspections complete. TE received signed Change Order #002 from Ginter today.

02/16/05 – Ginter reported that new fire alarm devices are being installed today. They will coordinate with building officials on final inspection. MS wrote a change order for Ginter to have Copp provide the power supply, ~$700, that was left off of the door hardware schedule.

02/10/05 – TE has coordinated with the City and Inspector. Ginter has submitted the new fire alarm permit drawings to the city. TE will attempt to keep everyone apprised on the new status.

02/09/05 – GBBN issued punchlist to JB Schmidt for this project.

02/01/05 – Change order #01 for Ginter was issued by ThermalTech to Contractor and County.

01/20/05 – Final Inspection scheduled for February 3, 2005. Building Inspector and Fire Marshal should be present. 6:00 a.m. Larry Keifer requested additional strobes. Te to coordinate with Honeywell to get this work done.

01/18/05 – The Archive room passed its pretest this morning no problem according to Mr. Bob Ginter. Honeywell has to hook up the outdoor air actuator still and that won’t happen until next week according to joe dale (via b ginter). then we can set final inspection. Peck has to modify the outdoor air duct so that it is not connected directly to the unit and blank off penetration at unit as well.

01/11/04 – TE checked on job progress today. Peck reported that unit could not be completely started up because two receiver tanks were missing from the condenser on the roof. Peck found them in their shop and brought them onsite. They will try to have their crew at the building on Wednesday to install the tanks and the associated piping. Peck also noted that the OA duct is connected to the return plenum at an access door that will make unit maintenance difficult from that side of the unit. TTC discussed with Mike and conferred with other TE engineers and there is no reason to bring in 200 cfm of outdoor air for this space. TE will direct Peck to remove the OA connection at the unit and just terminate the outside air duct at the wall with a grille. This will allow a small amount of OA to enter the space to meet the one person 20 cfm requirement by code. If in the future, we need too, we can add an additional exhaust fan to the OA to bring in small quantities directly into the room. This is a better design and use of the equipment in my opinion and the factory startup technician agrees. Please advise TE that this is acceptable to the County.

01/13/05 – Ginter faxed pencil copy of revised Application #1 which has been approved.

01/12/05 – Called Ginter and Spoke with Michael. Michael remembers the problem with App. #1 and didn’t know that the Application was hold and should be resubmitted. Michael thought that it had been sent to Hamilton County and was upset that he would have to resubmit and wait 30-60 days for payment. Michael asked to revise Application #2 combining the Payment Due amounts of App 1 & 2 and resubmit the Application as #1. We’ll skip submission of Application #2 and use the already submitted application #3 which will be on hold for the final documents.

01/11/05 – Ginter Applications #2 & 3 reflects previous payment for Application #1. Application #1 has been on hold since November due to the incorrect change order amount. Contacted Ginter, spoke to Jim Ginter and explained why I could not submit Application #2, he said that I should talk to Michael who is not in the office and that he would have Michael call me.

01/10/05 – Ginter Pay App. #’s 2 & 3 received.

01/03/04 – JB Schmidt is installing the doors and security hardware on Monday and Tuesday and Ginter will hook up the security right afterwards.

12/30/04 – TTC & MB walked the site and MB pointed out that the Danfoss valves being installed may cause more trouble than they are worth as the piping is sold old and the connection to it could result in steam leaks in the piping below the concrete slab. TE will direct PHB on how to proceed

12/24/04 – TTC talked to Peck for and update. The piping and sheetmetal guys should be back on the job on Monday, 12/27. They are currently under the gun at the Quest building downtown and this is pushing our schedule back. TE emphasized that this project is scheduled to be completed at the end of the year and Doug Turner said a significant portion of the remaining work could be completed this week. Hopefully they do make it on site and a lot of the ductwork and piping gets completed before our next construction meeting. TE has to contact Ginter because communications questioned the width of the cable tray Ginter installed. I don’t see making it bigger now, but let me see what they say. All of the remaining HVAC equipment including controls and steam “donfoss” type valves are on site or Peck has them for installation. Mike Bartlett questioned the rubber pads under the feet of the Liebert units. TTC will check into this item with Peck.

11/29/04 – Liebert unit and crane delivery scheduled for this Saturday. TE forwarded request for CH personnel for this weekend work onto the County. TE requested that TE, PHB and Mike Bartlett get together later this week

11/10/04 – TTC coordinated with Peck on the feet of head available from the condensate pump on the Liebert unit and what is called for on the drawings. Pump has 20 ft of head so overhead piping can be ran about 15ft above the floor and along the water line route to the shaft and into the drain per the drawings. AC unit is still schedule to be delivered the 3rd week of November. Peck will still be coordinated the crane and equipment lifts for that weekend. If anything changes TE will let everyone know.

11/09/04 – Ginter Pay Application #1 received with incorrect Chg. Order Amount Jim of Ginter notified and will resubmit.

11/09/04 – Ginter Change Order #1 sent via courier to Ginter for signature and then to Hamilton County.

10/21/04 – Mitch from HCCD at last meeting requested an EMT conduit or cable tray from existing telephone data closet to archive computer room. TEI to get revised sketch to Ginter Electric for pricing.

10/08/04 – TE visited the site. We noted the walls are up to the deck, the caps are on the ceiling cavities per our last meeting. The Liebert HVAC unit is not expected to ship until late in November. Pet.

09/30/04 – Rough in inspection complete for Electrical, Some supports installed, waiting for GC to install doors for door contact work.

09/15/04 – Construction Meeting complete. Project on schedule. See meeting notes below.

09/13/04 – Submittals received from PHB on Mechanical Equipment.

09/13/04 – Permit ready. JC Schmidt to pick up and pay for with reimbursement from PHB

09/03/04 – JB Schmidt on site. Moving studs up into space.

10/08/04 – Sheetmetal caps were installed on existing openings in ceiling.

by TEI TEI

Courthouse – Mediation Department

07/14/05 – Apparently a change order for the VFD drive was never encumbered and this has prevented Reupert from getting paid. Diana said that Mediation was going to work on getting it taken care of.

03/25/05 – Reupert submitted closeout documents and paperwork for final pay application for project. TTC approved and sent to Ms. Supe.

by TEI TEI

CH – Pumped Condensate Project

04/20/05 – All final paperwork arrived for TE approval and submission on this project.

03/16/05 – Process will have final paperwork and payapps in this week.

03/14/05 – TTC contacted Process about final paperwork and payapps.

02/01/05 – Process contacted County to state that project is complete.

01/25/05 – Project punchlist sent out by TE.

01/18/05 – Process has completed pumped condensate work including the summer boiler relief vent. TE to stop by and punchlist on Thursday.

01/14/05 – TE approved 1-1/2″ insulation on all condensate for this project. AKM approved a $1600 invoice for Process to tie an existing summer relief vent into and existing 1-1/2″ pipe with a ball valve. Process is to submit an invoice for this work to Mike Bartlett today for approval so they will not have to pull of the job to do the work.

01/10/05 – Process came on site and began work on new piping for pumped condensate.

by TEI TEI

Courthouse – VFD Upgrade on CT-3 and CT-4

08/08/05 – Final Pay Application was received. Allowance was straightened out. Ginter will keep $1000 allowance and pay DK $680 for additional programming work at the site to set the condenser water temperature to a constant number.

07/18/05 – O&M’s arrived to from Ginter Electric. TE will deliver to County offices on Tuesday by hand delivery.

07/06/05 – TTC delivered final paperwork and O&M’s to the County today. They were hand delivered to Martin.

06/29/05 – Ginter delivered the O&M’s to ThermalTech today. TE will give to the County.

06/16/05 – MB reported that the new cooling tower sequence is still maintaining a very stable condenser water setpoint. All parties are very happy at this time. TE emailed Ginter about final paperwork and punchlist work.

06/06/05 – TTC emailed punchlist to Ginter for final work and sent closeout documents with it for final pay application.

05/20/05 – TTC completed punchlist.

05/10/05 – MB reported today that VFD startup was completed by DK and Stoermer’s Gary John. The County reset the vibration safety switch and cleaned on of the low level sump alarms and CT-3 and CT-4 respectably. Ginter said they would check the wiring if that didn’t fix the alarm issue at the panel. The new program to control CWS/R temperature is in place and working from all first test. As the County brings on all 4 towers

04/22/05 – TTC and Dennis Faul worked on finalizing the cooling tower programming. MB and Mike Denton approached TTC about concerns about cooling tower control algorithms especially where concerned about the screw chiller. To that end, TTC has revised the cooling tower sequences such that a steady state condenser water will be more achievable. It is my hope that this alleviates many of the concerns expressed about this cooling plant. Here is the revised control sequence for approval and comments PDF. This should make commissioning much easier as this control strategy is much easier than the original.

03/15/05 – Programming is partially complete. Testing of program must wait till warmer weather. Project is in a holding pattern until that time. Ginter confirmed today that this was the

case.

02/01/05 – TE ready to punchlist these projects within the week.

01/25/04 – Dennis Faul returned TE’s call yesterday. The DK technician is onboard running the electric for new controls. We cannot test the system until we reach summer conditions but DK is still going to put program in soon.

01/21/05 – TTC contacted Jim Ginter of Ginter Electric in regards to 11,640.00 controls work which was not completed and should not be included in the payment due amount of the pay application. Michael Ginter delivered a revised Application #1.

01/13/05 – Ginter faxed Apps 1 & 2 to TEI for review prior to mailing. Changes required to both. Marked-up and faxed back to Ginter.

01/12/05 – Called Ginter and spoke to Michael. Asked him to resubmit both Apps 1 & 2 with allowance included. Michael to redo & fax copy for approval.

01/11/05 – Called Ginter, Jim Ginter said I would need to talk to Michael who is not in the office and he would have Michael call TEI.

01/10/05 – App 1 @ 22,024.80 and App 2-final @ 1,915.20 rec’d. w/o allowance included in contract price or on Sched of Values. 01/03/05 – Ginter provided a Change /Order deduct of $458 if we allow them to leave existing disconnects in place. They are worried that new dimensions will force changes to existing PVC conduit and this could result in problems with breakage of brittle material. County to advise TE on how we want to proceed.

11/18/04 – TE approved the design of all 4 cooling towers being routed from the main MCC in the penthouse. TE will provide new stamped design drawings for Ginter to submit to IBI showing the new work.

11/05/04 – Pre Construction meeting.

10/27/04 – Contract approved, TE to setup kickoff meeting.

09/24/04 – Mailed Contract to Tony for approval.

Prebid complete. Addendum 1 distributed.

by TEI TEI

CH – Relocate Basement FCU

03/30/07 – DK called today to inquire why they weren’t paid their final monies on the bsmt FCU project. TTC resent the 07/14 data to MWH.

07/14/05 – Received final invoice from Debra Kuempel today. It will be voided because of billing errors.

04/11/05 – TE talked to Michelle Combs today. DK sent an invoice in December for this project. TE does not have a record of this receipt. Michelle will resend a full invoice today for TE approval. TE will send to MWH immediately afterwards.

11/18/04 – Project complete. No punchlist items to complete. All units are running and trades finished up wall and electric work.

11/10/04 – TTC talked to Ed Enderle today. DK is hoping to be 95% complete by the end of the day on Thursday. TTC will visit the site and check on progress. DK is also going to seal up the walls during this project. HamCo will have to reconnect the electric if they haven’t already done so before units are operations

11/08/04 – TE visited the site on Friday afternoon. Fan coils are relocated and piping is in the midst of being changed over. Ductwork has been removed from through wall condition. Wall penetrations are open. Ceilings are partially disassembled for FCU relocation, ceiling tiles are stored inside the rooms.

10/21/04 – Pre Construction meeting complete. Debra-Kuempel scheduled to begin the first week of November. See notes below.

10/14/04 – TTC messed up the pre construction and it has to be rescheduled. Debra-Kuempel and TE will be onsite on October 21 at 8:30 am for the next meeting.

Bidding is complete.

County: ThermalTech is awaiting Notice to Proceed

by TEI TEI

Courthouse/Justice Center – 2003 Boiler Plant Upgrades

11/02/05 – Triton called to say that the certified payroll will be delivered tomorrow. This should wrap up the project completely.

10/18/05 – Triton’s final payapp came in. TTC approved and will give to Tony Matre on Wed via hand delivery.

10/03/05 – The valves and spool piece have been replaced. Triton is getting an insulator down to the site for final work. The final pay application is en route to ThermalTech for final approval.

09/22/05 – Triton is scheduled to replace the defective valves in the Courthouse this week on Saturday. They will being working on Saturday morning. The CH will turn off the steam on Friday to allow the plant to cool down.

09/13/05 – Triton came down and measured for the spool piece today. Mike really wants it installed by 10/01 and Ed Makin said they would try for two weeks. Mike asked about putting back an isolation valve in the spool. TTC said that TE prefers to keep the main wide open with no restrictions plus the pneumatic valve provides isolation.

09/12/05 – Today we kicked of the low pressure steam flow meter project. Process will do all the work during the first week of October. Triton is coming down tomorrow to measure the spool piece for the Courthouse valve replacement.

07/18/05 – TTC contacted Brian Pastura by both email and voicemail today to discuss resolution on the boiler plant valves and back pressure regulator. No return contact as of this email. TTC will continue to follow up with Triton.

07/06/05 – TTC found a quote from JCI to control the bypass valve on the supply to the JC. Based on the Spirax/Triton quote this may be the best solution to the steam supply issue in the

wintertime because the full flow works as proven last year through the bypass. By automating this valve to try to match what the JC needs, we can control total flow and still have the shutdown of the spirax valve to prevent carryover. PDF. TTC sent out Add01 for the Steam Flow meter project. The new bid opening date is July 12 at NOON. PDF

07/05/05 – TTC emailed Triton and Spirax trying to get the repair price out of them for the steam valve and to find out how to independently control the valve without DDC as TTC requested and Mike Bartlett questions (TTC did not have an answer for).

06/15/05 – TTC emailed Alpha Controls about new bid project: Mike, We spoke about 3 or 4 months ago about replacing the existing flowmeter at the Hamilton County Courthouse. I now have to bid out the replacement of the flow meter at the Courthouse. I will need the specs on the valve that you believe we should install. If you remember our final decision was to abandon the valve in the vertical and go with a steam flow valve downstream of the PRV in the low pressure supply to the building. This will allow us access to a large header from the floor with significant straight length. We will provide all the bid documents for three to five mechanical contractors to give us pricing on the job. I will ask them to hire you to provide the meter and do the control hook up. I will need a budget price from you now for your portion and I will need your help on spec’ing the right model number and getting the right flanges and connections in to accomplish what we are trying to do. You will have take the existing control leads and reconnect them to the new flow meter we are providing. As you already know – we don’t need exact lb/hr reading for some sort of process control but rather a flow meter that can give reasonable accurate readings from 0 – 10,000 lbs/hr or so.

06/15/05 – TTC and Pete Rorison talked by phone today. The existing valve will not and can not work under the existing conditions. Spirax can provide a Globe Control valve that can work in this situation. They cannot however, provide this valve at no cost to the County. Pete said the original valve was designed at what they believed was the correct design conditions at the time. Spirax will provide a new control valve, actuator and positioner and re-use the existing pilot positioner tap for control. Either the County or Triton will need to run a new pneumatic line over to the controller/actuator for the motive force to operate the new control valve. TTC instructed Pete to coordinate with Triton and get a best and lowest price together to do this work. There will be a credit for returning the old valve in lieu of the one we may install. When TE receives this new price, we will pass it along with a recommnedation.

06/14/05 – Pete Rorison emailed TE back today. Many of the issues in his memo were not correct. TTC sent a response email to rebut what was said and to try to clear up exactly what Spirax/Sarco is ready to do to make things right. TTC will follow up with Triton tomorrow for final resolution.

06/10/05 – TTC emailed Brian and Pete with a final direction on the Steam Control Valve. TE told them to have a solution in place by June 30th. PDF

06/09/05 – Triton called today to inquire about status of project. Per TTC’s conversation with Tony Matre on 6-7-05 there needs to be some resolution on the Spirax/Sarco valve issue before final payment is received. TTC will call Pete Rorison to see if another valve will work. TE sent out an email to discuss with County PDF.

05/20/05 – TTC and MB discussed the findings of Randy from PHB. The final word is that the steam valve is working correctly. TE next will make a recommendation about increasing the steam pressure in the boiler plant up to compensate for the low flow through the Spirax Valve. Another option is to set the max flow with the bypass open a certain % (field verified) to meet the max demand of the JC in the winter time. PHB stated that the 20:1 spring is the correct spring for the pilot and that the pressures seen on the inlet and outlet of the valve are correct for this valve.

05/10/05 – Triton’s revised invoice arrived. TTC discussed with Jim Drye the steam valve and decided to talk to Rich DeBat, a former Armstrong rep, about this control valve. TTC also contacted PHB about troubleshooting the existing valve. TE had Bernie provide a name of the their steam control valve tech (Randy) who is scheduled to meet Mike Bartlett on Friday at 6:30 a.m. to look at the valve and maybe do some troubleshooting for us. I told him to bill ThermalTech for now (as a reimbursable expense for now) We can judge how much of this time and work we need to help us out. After he gives us a synopsis, TE will provide a permanent fix for the boiler plant and we understand this has to happen this spring/summer.

04/20/05 – Triton and TE discussed in great deal when Triton is going to get paid for their remaining outstanding balance on this project. TE sent an email to Tony about what Triton wants to do – bill the whole contract now at this point. We are waiting on the County to decide about this.

04/19/05 – AKM & TTC talked about how to resolve PRV issue and Tony said that he has $10,000 to get these final two issues resolved through Triton Services. They will charge to the existing PO’s that Tony has available for this job. TE promised to have these issues taken care of this summer before winter ever gets here.

03/25/05 – TTC and County discussed solutions for the PRV valve since Spirax cannot provide a valve big enough to handle our load. TTC wants to bring in another engineer to discuss placing a second control valve in line to meet the maximum demand and still protect the Courthouse from steam spikes of the Justice Center.

03/14/05 – Triton delivered all final paperwork to ThermalTech for this project. TE to review and deliver to County.

03/11/05 – ThermalTech, County, Triton & Spirax/Sarco met to discuss the regulator valve issues at the CH. Several problems were discussed and solutions were also discussed. The bottom line is that the valve is not sized large enough to handle the full load required at the JC and still maintain the steam pressure required at the JC. Per Mike Bartlett’s excellent suggestions, TTC is pursuing JCI about automating the bypass valve to make up for this discrepancy. More info is coming in the near future to further discuss this issue.

03/08/05 – TTC has spoken to Brian Pastura about getting chuck on site to make these changes and they will do what they can but would like to have a meeting between TEI, Mike, Tony and Triton to decide on how to handle the workings of this valve for now and in the future. i agreed with brian that this would be a good idea and asked him to set up this meeting.

03/07/05 – Mike reported that Alpha’s service called yielded some design and equipment problems with the existing flow meter: Mike Maloney from Alpha Controls came in Thursday to check the Courthouse Yokogawa Steam Flow Meter. He found the sensor not working and recommended replacement. He estimated a replacement cost of just the sensor at $1500.00. After looking at the meters specifications, I found that this meter does not read steam flow under 3000 lbs/hr. An 8 inch valve is set up for reading between 3000 lbs/hr to 40000 lbs/hr. We believe that the Courthouse steam flow would be no more than 10000 lbs/hr. He recommended downsizing the valve to 4 inches and redoing the piping to allow for more straight piping that is recommended. He says that right now the 8 inch meter does not have enough straight piping for correct flow readings. He estimated a complete 4 inch meter at $2000.00. TE recommends doing this work to get this additional flow added to the Master flow readings on the front end for use and troubleshooting reasons.

02/24/05 – Pete Rorison called today to basically say that ThermalTech gave him the wrong design conditions for the valve and that it will meet the design conditions of 65 psig inlet and 50 psig outlet at 16,000 pph. The problem is that the JC suffers and 8 lb pressure drop at full flow so the valve should be sized at 65 psig inlet and 58 psig outlet. ThermalTech sent their rebuttal and the Submittals data stating the the valve was to operate between 50 – 80 psig.

02/15/05 – TTC contacted Triton and Spirax about pressure regulating valve and its proper workings. We will advise when we have more information.

02/01/05 – TE contacting the manufacturer about valve failures. Plants scheduled to be shutdown and worked on when OAT is above 50 degree F.

01/25/05 – Triton tried to fix the pilot positioners but the Powell Steam Valves did not hold. TE requested that Triton contact the mfgr and get new valves to replace the defective valves. The spec and the contractor wanted positive shutoff valves in this system. MB may talk to JC about a steam plant shutdown to replace the valves and do the pilot control work on the steam pressure regulator/backpressure control valve.

01/14/05 – Tim Cancilla is supposed to be onsite today to readjust the controller valve for correct steam flow and to rectify pilot tubing installations.

01/12/05 – Called Tim @ Triton to verify final doc list rec’d. Tim said finals will take awhile to get together and that he needs a Affidavit of Contractor Prevailing Wage form (2 Affidavits mailed), App 5 sent via courier to AKM.

01/11/05 – Revised Triton Apps 5 & 6 rec’d, 6 on hold for final docs, 5 signed.

01/10/05 – TTC request I notify Triton of incorrect contract amount on Apps. Tim w/Triton notified that Apps 5 & 6 should be resubmitted w/correct contract amount & that 6 will be held for final docs, refaxed final doc list.

01/04/05 – Tim with Triton, faxed a pencil copy (Application No. 6) requesting $27,206.76 with zero balance. Pencil copy approved.

01/04/05 – Email AKM regarding status of Triton App 5 pymt.

01/04/05 – 010405 App 7 received w/zero balance ( Requested 9,571.00 Change Order amount). App 7 voided due to incorrect $’s on G702 and G703, Called & faxed mark-up of 7, requested a revised App be sent as App 6.

12/10/04 – Follow-up with Triton Services, Timothy Herzog regarding final documents. Mr. Herzog requested that ThermalTech re-fax the list of required documents (which was done in the am of this date).

120605. T. Herzog contacted & fax sent informing that App 6 on hold for final docs.

11/22/04 – Triton Services, Timothy Herzog contacted regarding final documents and a list of documents required faxed to Triton.

11/16/04 – Process has replaced the traps and they are insulating them. Kerry approved contract changes and TE will deliver the documents to Tony for final work, Triton will bring diamond plate down for installation and have their insulator finish up the piping on the boiler room. Mark Neary returned my phone call and I have to follow up with him on flow meter recalibration.

11/04/04 – TE contacted Triton about finalizing the punchlist on this project including the diamond grate, insulation, flow meter, summer boiler vent line. Process has been told to replace the steam traps asap even pending the outcome of the purchase order submitted. TE instructed Doug Doxsey that in addition to the traps and the pumped condensate lines that a steam trap to drain off the summer boiler may also be a change order the County is looking to get prices on.

10/27/04 – TE sent change order request to Process for steam trap replacement. PDF

10/22/04 – TE scheduled to perform punchlist on project. Insulator has wrapped up majority of work. Steel grating not installed yet.

10/21/04 – TE has to get a steam trap decision to Mike Bartlett this week. TTC has to send data to Michigan to get answer from our special Armstrong Engineering.

10/21/04 – TE and MB discussed having process add the summer boiler vent to the new venting scheme instead of floor connected.

10/20/04 – Three Original AIA 101 Process Construction Contract document and Original plus two copies of submitted contract documents, from Process, sent via UPS to Hamilton County Dept. of Facilities.

10/19/04 – Original Contract documents received from Process Construction.

10/14/04 – TE approved Change Order #2 and sent 4 copies to Brian Pastura for Triton signing and then onto Tony Matre. Also the insulator is making steady progress on finishing the rest of the boiler plant insulation work. Once complete TE will do final punchlist.

09/29/04 – Triton is starting on concrete work, Neary provided controls programming earlier this week and appear to have done a nice job, Mike Bartlett reported that a recent Justice Center steam spike of 7000# reduced JC pressure from 55 to 14psig but the CH plant went from 55 to 51psig thus the new regulating valve appears to have worked as the plant maintained nominal working conditions. TE to schedule their punchlist as soon as insulators, concrete and controls are complete by Triton.

09/24/04 – Tony approved bid for Process at $9680 for pumped condensate work. TE will write contract.

09/24/04 – Water hammer is occurring at the CR-2. Stoermer-Anderson and TE will have to study this scenario and advise the County as to what might be occurring and how to eliminate this effect. Theoretically with no “live steam” going to the receiver, we should not have steam hammer at this piece of equipment.

09/24/04 – Triton submitted the NO BID paperwork so the bid tab is correct. How should we proceed. Please advise.

09/23/04 – Triton updated TE on project status today. Concrete contractor should be done in the next couple of days. PCI should be onsite today starting insulation. Ed Makin will be back down to put the bypass around the city water to the CR-2 on Friday maybe.

09/23/04 – Bids are in for pumped condensate work. Process is low with $9680 bid. Bid Tab – PDF

09/13/04 – Triton has completed work inside the boiler room and removed all tools. Room has been cleaned in general. Outstanding items include: Controls work, final inspection, concrete work, lead/lag switch on CR-2, plumbers must fix softeners so that we can start boiler plant, air control regulator must be replaced, insulators must do their work, steel grating must be installed at entrance to CR-2 room.

09/08/04 – Inspectors questions flash tank drain going directly to floor drain. TE to coordinate in morning on 10th. Code states there must be an air gap of 1″. Existing condition did not have it installed that way. TE to resolve.

09/03/04 – SA instructed us to take the new 3″ pumped condensate into all 3 of the water inlets in the top of the tank. This is a changes from the design documents. Triton will have to submit a change order for this work.

09/02/04 – MB reported today that Triton installed the 3″ make-up from CR-2 to DA into a 3-inch opening without a spray head. After much discussion, it was decided that Triton has to install the piping such that the 3″ CR-2 feed must tie into (2) existing city water connections both using spray heads. The remaining opening can stay in place for an emergency city water feed control to the DA.

09/02/04 – TTC and MB decided that after the plant is complete, we will begin to run the plant at 65 psig just like it has always been designed to operator and therefore we won’t need the PRV’s installed at this time.

09/02/04 – Dorn contacted TE today. TE gave them a PO that will be reimbursed during subsequent invoice from TE to do work on moving sprinkler line on T&M basis in CH walk-in cooler.

09/02/04 – Mike Bartlett report that Spirax stated that strainers must go in front of the PRV valves. I checked with Todd and Gary and the HTX are rated for 125 lb service and they should be able to take the higher pressure in the boiler plant. I am recommending that instead of incurring more cost, we just return the valves and take the credit to the job to offset some other change order cost on blowdown lines.

09/01/04 – TE talked to Spirax engineers. Both PRV’s must be installed for HTX’s, flow meter can go in vertical, they will provide flow differentials based on varying steam pressure.

09/01/04 – Stoermer Anderson confirmed that summer boiler relief valve is already set at either 80 psig or 100 psig, so no additional work is required.

Triton Services will be back on site August 16 to begin finalizing the project work

CR-2 and DA piping will require a FULL plant shutdown scheduled for August 20th.

Summer boiler must be online independent of the main plant by August 20.

TE will coordinate with Stoermer Anderson for final Summer boiler startup details

Summer boiler will use treated city water during 3 week shutdown as work is completed.

TE and Triton to coordinate schedule with Mark Neary for new control sequences and SA for boiler work and equipment changeout and use.

County will provide an airline close to top of CR-2 so that project doesn’t incur additional change orders

Boiler plant work scheduled to be finished and back online by September 10, 2004.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: There are (2) 3″ inlet to the DA tank, can we remove the 2″x3″ bushings and run the pumped condensate directly into the DA tank. Response: The bushing and utilize one 3″ inlet for the 3″ pumped condensate from CR-2. Since all material are now steel, the existing dielectric flange for the brass fitting will not be required. Keep one of the existing connections in place and utilize the existing city water connection and controller as backup or emergency fill if ever needed.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Triton asked TE to locate the circuit setters shown on the drawings. Response: TE advised Triton to NOT install the circuit setters shown on the drawings and save the money to use on the blowdown piping.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: Can 6-inch piping to steam regulator on catwalk be maintained in size. Response: TE and Triton met onsite to discuss field issues and many small items were resolved including the piping layout for the new back pressure regulator valve on the catwalk service the JC. 6 inch piping can feed the JC and increase to 8-inch immediately after the final isolation valve just like existing piping is installed. Existing Justice Center main will now be bypass. Per the contract documents Triton will replace regulator on wall with new equipment to operate bypass valve.

Vendor: Triton Services | Date Issued: 08/23/04 | Date Answered: 08/23/04 Question: How do the new control valves get place and installed on the project? Response: The new control valves for the CR-2 and DA-1 shall be installed per the direction of Stoermer Anderson. The existing city water feed to the DA can remain in place as on 3″ outlet can be used for the feed from the CR-2 unit and the other 3″ (bushing down to 2″) can remain as emergency feed to DA.

Vendor: Triton Services | Date Issued: 08/24/04 | Date Answered: 08/24/04 Question: The original separator location was in the Justice Center. Now the separator is being installed in the Courthouse. This is a change in the contracts. Please advise. Response: Brain, you are correct the original location was in the Justice Center but it was to be installed over 15ft high in the air requiring lifts and extra manpower to reach it and install it at the location shown on the drawings. All of this work is included in your base bid. It is my belief that the new location provides a significant reduction in manpower and installation time. Therefore I see no reason to incur additional cost to the project for this project change as we can just use that savings to apply it to the additional work to reroute the condensate line which should only include 2 90 degree elbows and their associated welds. The rest of the scope remains the same.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: There is one 3″ flange on top the DA that has significant pitting. Should this be replaced? Response: Please replace the flange at the weld neck. Make it suitable for the blind flange required in the construction documents.

Vendor: Triton Services | Date Issued: 08/25/04 | Date Answered: 08/25/04 Question: The separator has a 1-1/2″ outlet on it. The TE drawings show a 1″ outlet with a 3/4″ condensate drain. How should this be remedied? Response: Please route the condensate drain as 1″ to the condensate system.

Vendor: Triton Services | Date Issued: 08/26/04 | Date Answered: 08/26/04 Question: What pressures should the new flow meters be calibrated too? Response: Set the Courthouse flow meters to 65 psig and the Justice Center flow meters to 55 psig. Name them CH-FM3 and CH-FM4 and JC-FM1 and JC-FM2

by TEI TEI

Courthouse/Justice Center Separator Installations

05/04/06 – Tony called today because Geiler’s final payapps arrived unsigned by ThermalTech. TTC drove down picked them up, signed them, made copies and sent them back down to the County for payment. Geiler did not sign the copies so TTC just copied the original after he signed it. Once paid this project should be complete.

04/10/05 – Indrolect inspection at the Justice Center is finally complete and approved. We are now ready to finalize and archive this project.

02/17/06 – I don’t think the inspection will be next tuesday because indrolect just got to the bottom of Geiler’s permit applications. apparently geiler took out two permits even though they didn’t know it. the city doesn’t even have a copy of the jc one in their file but on the computer it says it was approved. ibi will accept the previous plans exam as valid for the work but now has to schedule another field trip to sign off on the jc wiring install. when ibi decides they can come out, indrolect will call us and let us know and I will call you and let you know.

02/114/06 – Dave Memory of IBI approved the CH install with TTC, MB and Don Ritter in attendance.

01/11/06 – Geiler and Indrolect have submitted for all permits. As soon as plans exam is finished, Indrolect will call for final inspection. Should be within a week or so.

01/01/06 – The job cannot be closed out because of permitting issues from the contractor. Indrolect and Geiler have been given the preconstruction meeting minutes where we state that permits are required on this project. TTC sent Geiler an email today saying that the additional $1024 in allowance money that is being paid to Geiler should be used to acquire all the permits on this job. TE will provide additional stamped sets of drawings as necessary for permits.

11/28/05 – TTC received all final paperwork and will deliver to Tony Matre on 11/30/05. This job is complete.

11/03/05 – Lisa Huber said that all paperwork has been sent to ThermalTech for this project. Certified Payroll will go directly to Tony after they get it on Friday.

11/01/05 – Finalized final billing strategy for final bill. Tony said they don’t have any certified payroll and they need final pay applications. TTC emailed everyone to let them know.

10/20/05 – Lisa from Geiler returned TTC’s phone calls about project closeout. TTC returned on 10/24 at 5:30 p.m.

10/18/05 – TTC emailed and faxed the change to Michelle Wells again in the hopes the final payapp will soon be processed.

10/14/05 – TTC contacted Geiler again for the correct final pay app. No work yet.

10/03/05 – Still no word back from Geiler on corrected pay application.

09/21/05 – We received the final pay application for this project as application #2 this time. This payapp is not right because of the change order amount is wrong versus the allowance. TTC contacted Geiler to advise them on how to fix it.

08/08/05 – TTC reviewed and will deliver the final paperwork to Tony on 08/09. We have not received a final pay request nor the signed punchlist. TTC emailed Barry Bowman to inquire about final pay application.

07/25/05 – TTC received in his mailbox the final paperwork.

07/18/05 – Contacted Barry Bowman about finalizing this project and sending all paperwork in.

06/10/05 – Relisted Punchlist PDF.

05/20/05 – After cleaning the system several times and removing large amounts of rust from the piping inlet to the separator and disassembling and reassembling the 4″ supply line to the separator, the team was finally able to start the separator and see significant flow at the cooling towers. The system when new provide 40 psig of gauge pressure at the outlet. TTC stated that with this much pressure and flow out of the nozzles that 2 towers could be scrubbed at once and the normal operation would be to run the separator with CT-1, 2 for a week then switch it to CT-3,4 for the next week. The punchlist will be sent but Geiler took care of issues at the site during startup. The strainer’s and filters should be changed a couple of times a day for the week or so. Monitor how much cleaning the SEP-1 unit provides and modify the cleaning schedule as needed.

05/17/05 – Startup for CH separator is scheduled for 05/20/05 at 8:30 a.m. with TE, Geiler and Trane.

05/16/05 – MB emailed today to say that all four CH cooling towers are online and ready for separator startup. TTC emailed the contractors and requested the dates. TTC approved application #01 for payment and us mailed to Tony Matre.

05/02/05 – Geiler promised that Bob or his associate would be at the JC today working on the punchlist items for the JC separator.

04/22/05 – TE contacted Geiler for a third time about completing punchlist work at JC. HS confirmed again that no punchlist work has been done. Geiler wants paid badly and keeps inquiring about payment and TE keeps pointing out deficiencies in payapp procedure. TTC believes that we have it resolved now. CH staff believes that their towers will be online next week for startup and testing.

02/11/05 – Punchlist distributed to Barry, Tony, Howard and Mike.

02/10/05 – Geiler asked to see the original meeting minutes where we told them to put a drain and valve into the Courthouse Separator system so it could be used in wintertime operation. TE to send out from original pre construction meeting.

02/01/05 – TE ready to punchlist these projects within one week. Punchlist is complete. TTC to distribute.

01/25/05 – Indrolect made the tie-in into the bus bar on the elevator equipment room yesterday without incident. We powered up the disconnect and then removed the fuses awaiting startup by Geiler. TE contacted Barry about status of project and if it was ready for punchlisting.

0/1/21/05 – No response to the previous day conversation. Followed-up with a fax to Geiler about Application #1.

01/20/05 – Pay Application #1 placed on hold. The Application was not signed or notarized and also did not include a completed, notarized, Affidavit of Original or Subcontractor. Michelle Wells contacted and told of the reason the pay application would be put on hold. She asked for a voice message to be left for Lisa explaining why the application is being put on hold (which was done).

12/03/04 – County approved Geiler change order allowance usage for redesign of PVC system for separator. $3976 will come out of the allowance on this job.

11/22/04 – TTC stopped at the Justice Center this morning and reviewed the existing piping layout with Joe Phillips, Howard Seal, Barry Bowman. What we are proposing to modify the design so that each tower can be isolated from the system and cleaned to prevent any of the “dirt” from the system from getting entrained into the condenser water supply to the chillers. Each separator could be piped separate with PVC and ran an determinant amount of time to clean the tower daily and then return to normal operation. TTC will provide a revised drawing to the Geiler company for this work today.

11/12/04 – TE contacted Geiler about completing the Courthouse and redesigning the Justice Center to meet actual design request. TE will follow up with Geiler.

11/04/04 – TE contacted Barry about field progress on this job.

10/20/04 – Geiler needs to install a drain and valve on CT-2,3,4 mains for winter draining. Install a drain on the main unit. Possibly relocate the existing control panel 90 degrees rotated to meet NEC requirements.

09/29/04 – CH CT 1 & 2 piping is complete. CT -3 & 4 and JC work to be scheduled in the upcoming weeks.

09/24/04 – Geiler estimates that the courthouse should be done is 1 – 1.1/2 weeks.

09/23/04 – TE emailed Geiler to inquire about project schedule update since progress is slow.

09/13/04 – Piping is being installed in Courthouse cooling tower #1. Looks good.

08/01/04 – Geiler asked how do they fit piping under fill? on 08/10/05, TE advised to decrease manifold size to 2″ and slide piping under existing fill. This will maintain integrity of cooling tower. TE provided a new highly detailed dwg of the tower piping for Geiler to use on all towers for this project.

Geiler has been onsite doing the work.

ThermalTech received written approval for sweeper design.

2″ pipes will be slid under existing fill.

Contract should be back on site soon.

Cliff Wingerter is on medical leave so all questions should be routed to Barry Bowmen.

by TEI TEI

Courthouse – Steam Flow Meter Project

01/03/06 – TTC received final paperwork on this project. TE will process it and turn it into Tony for closeout on this project.

12/02/05 – TTC contacted Process again for final paperwork.

11/28/05 – TTC was at the CH today and discussed with Mike Bartlett final job status and any punchlist items. MB did not report any further punchlist items and TTC did not record any incomplete from original walk through that asked for the insulation to be complete and the programming and scaling of the steam flow from the new meter to be re-adjusted. These issues are complete and therefore no further punchlist items are required at this time. TTC approved allowance usages to repair a defective gasket on an isolation valve during the project for $1000 which included some miscellaneous other travel charges and such to get parts for this replacement and a start/stop scenario while ThermalTech gathered information to allow replacement to take place. TTC will contact process for final payment.

11/08/05 – Mark Neary reprogrammed the Teletrol system, Mike Bartlett confirmed the flow readings were correct. Insulation on the old steam flow meter spool piece is not complete yet but everything seems to be working fine.

11/07/05 – Process sent in pencil copy payapp. TTC contacted Mark Neary for online programming of the Teletrol trying to save paying a trip from Columbus for work.

11/01/05 – Project is complete aside from reprogramming the Teletrol system to the new range of inputs from the low flow meter. TTC is working with Mark Neary to get him to dial in and change the expected input to match the actual flow. Mike Bartlett reported today that the partial flow was around 1900 lbs/hr.

10/24/05 – TTC emailed Mike & Clem to get project update.

10/21/05 – Mike reported that the main steam flow was back up and running for the JC. The flow meter eccentric reducers were being welded in. One was already installed.

10/18/05 – Process is all set to begin and finish work this week on the project. The steam plant is down and cooling off. TTC will stop at the site on Wed morning to coordinate and then go to a prebid on a different project. Process will attempt to do the work on the existing steam meter first and get the plant back up and running on Wednesday.

10/14/05 – Personnel issues are preventing Process from starting but Clem reported that they can start on Wed 10/19 at the CH.

10/04/05 – We were trying to be done with the Steam Flow meter the first week of October. TTC has contacted Process about availability of parts and getting this work completed.

by TEI TEI

Courthouse – LIP Schematic

10/14/05 – TTC has inquired about status of project. If the new door is being handled within, then this project is complete. Diana and Mark please advise as to the status of this project.

05/12/05 – We all discussed alternate schemes for a new door to meet state and local requirements and TTC sent this color scheme with 2 options to GBBN for review and for GBBN to forward to the County for approval. The County shall make the final decision as to which way they want to proceed. PDF

04/26/05 – TE recommended provide a new door in the space to meet code requirements from state and city per this PDF. TE must wait for County to approve a way to go on this issue before project is complete.

02/2/05 – TTC talked to Mike Bartlett and found that complaints have subsided since Greg greatly reduced the total amount of air being fed into the space. This makes sense since the space was being overcooled because the thermostat was on the second floor in the courtroom. Mike and I agree that this should work until warmer weather where the airflow into the space will have to be increased and further substantiates the need for a revised VAV system in this space. As it right now, Mike and Greg are acting as a manual VAV system by increasing and decreasing air into the space as needed. This also shows us that if you are really crunched on money, you may be able to only put a couple of VAV boxes into the first floor area and begin to achieve some control with the hopes that a full redesign can be done later.

02/25/05 – Client reported that space is still not conditioned during the evening time properly. The daytime loads can still never be fixed without a new VAV system and HW reheat (see SOSS TE-115)

02/10/05 – TE has not seen a balance report or payapp for the additional work.

02/01/05 – MB reported today that Room 169 reported no comfort problems with the space except when the main ahu turned on. PHB still has to balance the new air device. TE to punchlist this job shortly.

01/31/05 – MB reported that PHB finished the job last week. TE has to confirm. We will not be able to assess whether this project was a success as of yet because the units are being operated 24/7 in that area due to other heating and cooling considerations.

01/24/05 – Peck started the job. They need a different transition and will be fabricating that in their shop and finishing the job in one or two days.

01/18/05 – Peck said they should start on Wed or Thur, Jan 19, 20 per Ken Shad.

01/03/05 – Peck should be onsite at the end of the week or beginning of next week to start after hours work.

12/24/04 – TE has given Peck a scheme to price up. Assuming they have the best price, they are the odds on favorite to do the work since they have an unused $5000 allowance at the CH in the Archive Room.

12/10/04 – Price came back from B&J at $6000, TE is now inviting Peck to give us a price since they have allowance money left on 2 County jobs. TTC will bring down Peck for a price quote next week. We will coordinate the visit with Mike Bartlett.

11/05/04 – TE met with Mike Bartlett, Bill Scholl & Jim (Bailiff’s) and reviewed the general layout of the 5-ton split system into the Bailiff’s area but the actual unit would be in the Room #167 space (this would have to be approved by LIP), ductwork would have to routed above the 167 space and over to the 169 area, 6 diffusers at 250 cfm would have to be installed to meet the 2000 cfm requirement of the 5-ton unit with a return ductwork. the final design is still under review. TE has asked B&J Jacobs to take a look at the design to provide their opinion about what is easiest and most cost effective. TTC will provide prices associated with each of these schemes so that the County can make an informed decision.

11/04/04 – TE discussed a ductless split system installation with Mike Bartlett, Bill Scholl, Diana Supe and left a message for Tony Matre in the 24 hour bailiff’s area today. TE will follow up with Tony about funding this project ASAP.

10/25/04 – Bill Scholl called today. A full size duct won’t fit through above the ceiling into the bailiff’s area due to a water pipe on the other side of the wall but a duct sized at 12″ x something will. I told them to find out and maximize the size of this duct and proceed with the design. We are trying to get as much air into the space as possible without spending much money. This is the first stage of the “fix”. Other things may have to be done later to enhance the space control.

10/14/04 – Mike Bartlett showed TTC B-12 which is actually the unit that serves the Concessions area. This unit is served by both heating hot water and chilled water. There is a temperature sensor in the supply duct outlet that monitors the leaving supply air temperature. TE believes that this is the sensor that Bill Scholl was talking about in our 10-12-04 meeting. When there is no building chilled water, the DX unit provides cooling for B-12. Our drawing and scheme will still work as the timeclock will divert air to the room #169 space starting at 7:00 p.m. – 6:00 a.m. and regardless of how B-12 is providing cool air it will get into 169 to assist in cooling the space. The system should still be modified as design on the ThermalTech packet.

10/12/04 – TTC met with DES, MB and Bill Scholl to discuss the reworking of the ductwork on unit B-9 to included Room #169 during the evening time. Trades and HVAC will do the work in the next couple of weeks. The plan will have B-9 provide cooling to the space after hours between 7pm and 6am. TE also stated that the ceiling work required should be done after hours as this will be a messy job especially in the kitchen area above the sink. We will monitor the temperature during the first few months and see if the County needs to add additional controls to the new air damper later. Bill reported that a Paragon timeclock will be used as they have one of those instock and the damper and actuator have been ordered. TE is fine with this information.

10/08/04 – TEI received a letter asking a question from the plan examiner. MS needs to talk with him because a letter will answer his question. However, he is out of the country till the 18th. They directed me to his supervisor Bill Spurling. TE have been playing phone tag with him for the past couple of days.

10/07/04 – TE visited the site and TTC has sketched up a possible solution to the Room #169 heating issue. TTC will deliver to the County and look at making the changes utilizing existing equipment for the work on 10-08-04.

09/13/04 – Fire Test at 6:00 a.m. Space did not pass. Egress into LIP has been questions by Larry Keifer. Security drawing was not included in permit set to City. Mike Stiles to rectify ASAP.

09/03/04 – TE delivered a set of permit prints to Brian Meyer for use on this project.

Security issues are ongoing between LIP and Recorders.

Trades doing all work on this projects.

10/18/04 – TEI received question from plan examiner and TEI responded. Plan examiner was out of the country until Monday October 18th. TE waiting for response.

4/19/05 – TEI recieved call from Mark Harden indicating that the State Agency that governs the computer system information will not accept the security level provided by the rear door in the computer room. However, the motion sensor that was added was per the the state building code. Mark and TEI to discuss options. TEI contacted Ben with GBBN to discuss the option of another 2nd path of egress. Ben to respond.

by TEI TEI

Courthouse – EMU Relocation Renovation

10/14/05 – Wayne Trent advised that some portion of this project is moving forward, TTC and Spencer are trying to find out what part of scope is confirmed.

10/03/05 – Spencer is still working with County to get project scope finalized.

08/23/05 – TTC and MS met Spencer Johnson and Mike Bartlett at the site to get field work to prepare for the permit drawing due at the end of the week.

by TEI TEI

COURTHOUSE – BREAKROOM REMODEL

12/26/06 – The project is complete and the work was completed by EDI. The only problem is getting the city buy into the final design. A problem exists between Steve Ossege and ThermalTech about what engineering dwgs were submitted on the project. TEI gave a engineer dwg to EDI to put the work in but the city permit doesn’t reflect it. Once this resolved the project can be closed out and archived.

12/04/06 – EDI has installed the drain pan under the FCU, TTC needs to resolve the dwg issue with the city examination department and get Steve Ossege approval on the closeout.

11/20/06 – TEI has been working with EDI to get Steve Ossege to approve the building permit for this space and rectify the deficiencies in this space according to Steve Ossege phone call to TTC at TEI about the space and the permit. The ductwork has been modified so the return is fed back tot he unit, Mark Donnelly’s office has an OA duct and a supply duct and the FCU will get a pan underneath it as soon as the water switch arrives.

05/26/06 – Dorn received building permit number from Spencer today and is applying for actual fire sprinkler permit work.

05/23/06 – TTC contacted Dorn sprinkler about the progress of the fire permit for the breakroom area. Dorn sent a dwg and a revised dwg to TE showing the changes. Bert Watts will inform us when the ceiling grid is in place and then Dorn will put the new sprinkler heads in the new space. This should work fine.

03/06/06 – TTC got a late start on the project that was supposed to be done last Wednesday but today finished up both the mechanical and plumbing but realized a big problem with the HVAC and asked for guidance from HamCo. TTC sent an preliminary HVAC plot to the County for review. The plumbing is finished and ready for permit.

by TEI TEI

Courthouse Electrical Device Identification

01/10/07 – Completed in 2006

1/08/07 – NR picked up the panel index binder from Hamilton County Facilities. There are no open issues left on this project. All documents have been turned over to the appropriate parties.

1/03/07 – NR met with Jim Eby of Hamilton County Facilities to review the as-built drawings for this project. The open items were resolved. (3) full sized sets of drawings have been delivered to Hamilton County. The binder containing each panel index is in the possession of Bert Watts. A copy of this binder shall be sent by Bert to ThermalTech Engineering.

1/9/06 – 2nd Floor through Roof is complete. Presently working on the Basement & 1st Floor

12/15/05 – 4th Floor through Roof is complete. Presently working on the 2nd & 3rd Floor

11/15/05 – Roof is complete. Presently working on the 5th Floor

by TEI TEI

CH – 2nd & 3rd Floor VAV Renovation Project

03/15/07 – TTC has received and printed out as-built drawings. The entire packet will be delivered to the Courthouse on Friday.

02/05/07 – EDI sent a revised payapp #7 in that looks fine but they also are asking for additional monies for ceiling work.

02/02/07 – EDI’s final O&M’s checked over and are fine. EDI stated that As Builts are included with them. TTC to confirm. TTC rejected pay ap 08 because it shows a negative balance due on this project.

01/22/07 – All Closeout documents received. TE set to work on as-built dwgs to turn over to County. O&M’s and closeout paperwork and pay applications to be sent directly to Tony Matre on 01/23/07. Project was archived by TTC on this day.

08/25/06 – O&M’s have been received. TTC will drop them off to Tony today by hand at the Elevator bid opening. TTC will schedule a punchlist for the job next week

08/15/06 – Payapp # 5 was received by TEI.

07/25/06 – Mike emailed today to report that balancing is supposedly complete. Honeywell needs to finish up because tenants are complaining about space conditions. TTC forwarded to EDI and will get the project completed.

05/22/06 – TTC talked to Mike and John both of whom like the proposed changes to the HVAC system for the courtrooms on the second floor of the new space. TTC will advise the EDI team that we can move forward with the revised layout. Here is a copy of the revised changes PDF

05/16/06 – TTC contacted EDI about change order on pay apps and about next const mtg with john and mike at the courthouse.

05/02/06 – TEI set a construction meeting between TTC, MB and John from EDI for tomorrow to go over project status.

04/11/06 – Const Mtg is tomorrow, TTC to have new change order drawing ready for EDI to go over changes in the space per TTC’s discussions with Mark Donnelly today at the CH site.

04/06/06 – TTC typed and sent out Const Mtg #2 notes. Mike called about fire damper locations. TTC responded that none should be required but if EDI found a certain fire wall we are penetrating that TEI will answer those conditions on a case by case basis.

03/16/05 – TTC wrote change order for EDI to do FCU relocation in basement.

3/15/06 – TTC talked to the Controls Contractor today to review the submittals data. Honeywell will field locate power supplies and provide quantities required for complete installation. They will coordinate with Mike Bartlett for power supply locations and update EDI as-built drawings for future record. Final questions was how many VAV boxes. We confirmed there are 17 new boxes. 9 on one floor and 8 on the other. The Computer room VAV is existing.

03/09/06 – Const Mtg #2 was held yesterday and the meeting notes distributed today. EDI emailed the employee list to Tony today.

02/27/06 – Pre construction minutes distributed. Submittals have been returned.

02/16/06 – TTC approved using NEBB approved balance reports on this project and in particular the sub-contractor Approved Products. TTC will send out the meeting notes later today to the contractors and owner.

12/14/05 – Contracts came back, TE edited, corrected, printed and sent back down to County for commissioner approval.

10/24/05 – Final Addendum went out on job. Other control contractors were added by Addendum. TTC answered MB’s questions about reheat temperatures in the coil locations in saying that coming off the end of the HHW loop is okay, we just need to have the coils selected properly at the time of purchase. 10/14/05 – Sam Boehringer will be handling the prebid mtg for this job as TTC has a traffic court date with his wife at the same time and day. TTC will join the party after court. Habegger called and will be added to the acceptable manufacturer’s list for this job.

10/13/05 – The project has bee publicly advertised. The ITB is 160-05. Bid Opening scheduled for 10/28/05. This is a strictly mechanical only project with Honeywell or Habegger as the controls contractor.

06/16/05 – TTC & SBO met with Diana and Mike to discuss possible scopes for this project. TE designed the full system with all VAV boxes and reheat boxes. The new cost is significantly higher with a full zoning and reheating plan ($161,000). An alternate plan that may work (only 8 boxes) was also given to Facilities ($51,000). It was decided that to half do the design would only waste money and inconvenience the tenants twice. Diana will approach Tony about the budget and doing the full job.

04/04/05 – TE and CH Facilities had first design meeting discussing scope of work and project goals. PDF

by TEI TEI

Courthouse – Blowdown Project 2006

07/17/07 – TTC emailed Tom Heyl today to tell him that TEI originally signed the payapp on 05/16 but resigned a blank copy on 07/13 at Tony’s office for payment. TEI also asked Process to do some piping work with the remaining $1000 on the PO if Tony says it is ok to do.

01/22/07 – Project complete. TTC not sure final paperwork has been received at this point.

06/27/06 – Bids have started to come in on the job and TTC is tabulated them. TTC will send tabulation to Mike and Tony today when complete.

06/23/06 – TTC sent out prebid meeting notes for project. PDF

06/12/06 – TTC met with Tony and Mike. Prebid on this project is set for June 20. TTC contacted Steve Hutton at GE about attending the prebid. TTC will contact the contractors to schedule the prebid.

by TEI TEI

COURTHOUSE – ROOM 468 JURY COMMISSION OVERFLOW

06/11/07 – TE received the final pay application along with the allowance clarification and close-out paperwork. All materials have been sent to HCFD.

04/27/07 – Mark Donnelly has directed TE to let QCE know that they should send their final pay application so that this project can be closed out through TE. Hamilton County Facilities with deal directly with American Sound to complete the remaining PA system work. TTC reviewing EDi Closeout for submissions to County.

04/13/07 – American Sound will directly quote project to revise PA system wiring to Mark Donnelly of Hamilton County Facilities Department. QCE will be closed out of their involvement with this portion of work.

04/05/07 – QCE is working

03/30/07 – Project reported to have received occupancy from City. Bulletin issued for PA system. Final paperwork from QCE after change order is needed.

03/23/07 – Space is occupied and there a couple of items left to finish and closeout. The PA system has to modified for client satisfaction. TE and MD are working with QCE to resolve the issue. TE and MD agreed on which QCE change order will be paid from the allowance (via email) and what will be a change order to the project (extra conduits).

03/20/07 – Punchlist are complete. Everything is working except the paging system.

03/15/07 – We received nice kudos from Frtizz on this job. Closeout meeting went well. County agreed to insulated the steam piping behind the radiator. TTC and Honeywell agreed that some control tweaking still remains to be done before project is completed. TTC received O&M’s from EDI last Monday and must turn them over after review to MD. Once payapps are received and O&M’s approved this project can be closed out for EDI and QC M.

02/23/07 – TEI distributed the mechanical punchlist for this project to EDI and partners. EDI forwarded to Honeywell and requested everything be completed by 02/28/07 for project completion. PDF

02/15/07 – EDI provided an update on the project as they are trying to do their balance report but the unit has no power yet and the controls are not hooked up. QCE was contacted about having power to the unit and was suppose to do that work yesterday as was Honeywell who did not make it down until today. QCE reported that they did have power to the unit but the fusible disconnect on the 5th needed to be turned on except that no one from QCE was there to turn that on. The balancer is on site today waiting for the system to be turned on to get the readings they need to startup the unit and get the balance report complete. EDI sent in a list of change order and allowance request that TTC processed and forwarded to Mark Donnelly for discussion. This list included the condensate pump, relocation of CHWS/R, steam lines, and fire protection supports. MD and TTC discussed by phone and the County is willing to pay at this point for everything but the relocation of the CHWS/R lines.

01/22/07 – Everyone met onsite today to discuss the MER. Space was allocated for the Controls, Data and Mechanical Systems. Electrical coordinated the locations of the additional power outlet for the mechanical sump pump and the receptacle for the controls. Owner stated that fire support piping will be hung from deck and not form interstitial ceiling in Room 468A. TTC showed EDI this space and EDI will adjust the hangers. Owner will not accept ceiling hung from ductwork. Radiators are complete. In reading room TE asked for Architect to check height of new cover to handle the concrete brick under the legs. Steam system is up and running. 3 of the 4 steam control valves are replaced. One could not be removed because of pipe rupture concerns so TE decided to leave as is. Honeywell submitted a 3-way valve for control on the job but that is not required and it has been plugged to a 2-way valve per TTC’s instruction. The new drain and chilled water piping will be installed above the unit at about seven foot AFF and drop down to the unit next to the unit leaving the unit open for access. We will trade the access door work by the GC for a square opening and two backboards for controls and data in the room. TE coordinated these locations and sizes with CRA and the contractors.

01/17/07 – TEI requested that EDI relocate the CHWS/R away from the disconnects and route over top the unit. QCE and CRA decided that the soffit can be lowered 4″. We will issue bulletins for the additional work and get prices for the County. Honeywell reported that their control panel will only require 8A of service which means it can be fed from our AHU unit or electric heater with a separate transformer and given the low IT requirement, it may share the same circuit as it. TEI called for a coordination mtg on MER on 01/22/07 at 1:00 p.m. to coordinated all these issues.

01/16/07 – TEI is meeting QCE to discuss additional conduits under the soffitt in the CH hallway. TEI invited CRA and HamCo to attend on Tuesday at 4pm.

01/10/07 – Sprinkler contractor was on site working on the sprinkler system on 01/09. HVAC work is progressing. TTC instructed EDI to use the condensate pump provided with the unit, that the integral check valve meets our requirements and to use the 1.25″ piping as called for on the dwgs to drain system. EDI asked if the 3-way valve provided by honeywell could be plugged to be a 2-Way valve and TTC stated that this was okay. EDI reported the the GC would be reworking the ceiling grid to match the dwgs. It is currently running in the wrong direction. The sprinkler contractor have the sprinkler mains ready for final arm overs down to the new ceiling and will come back after ceiling is reworked. EDI inquired today about placing the radiator on concrete blocks, I am not sure this is an idea I am going to approve.

01/02/07 – The AHU was delivered into the MER on 12/29/06. EDI will be working nights to hook up the HVAC to the new unit runs. Honeywell will be in to hook up and download the controls soon but an exact timeline has not been set. TTC asked the County to allow us to use $600 for a steam condensate fix for the old air handlers located in Room 468. MD has given TEI approval to do this work under the $2500 allowance for this project.

12/21/06 – Const Mtg # 2 was held today. EDI has VAV boxes installed, all ductwork hung except for MER. EDI stated FD will be installed when unit is installed. Unit is scheduled to be arrive on December 27. QCE needs to communicate better with field visits

12/14/06 – TEI was onsite today to review mechanical and electrical progress. On the Mechanical side EDI has all the main ductwork installed up to the mechanical room. The VAV boxes are installed and the sprinkler work is installed except for the arm overs that will eventually be connected after the ceiling is installed. The exhaust fan is installed but the new louver is not installed yet but will be next on the agenda. We are waiting for the unit to arrive before more work is completed. We discussed the new radiator location in the small reading room and TEI instructed EDI to try to fit the existing radiator in between the piping and use piping crosses to connect before moving the radiator up or more into the room. They will try this and advise. I do not believe that any controls programming has been installed yet. TEI is meeting with QCE and Mitch on site at 4pm to go over data cabling, power routes and the shutdown of power today for our tie-in for the new AHU unit. Mike Bartlett is keeping an electrician late today to assist in case QCE runs into trouble.

12/04/06 – Work is progressing. EDI has submitted a full schedule which I believe will be incorporated by the GC into the master schedule. EDI informs me that the VAV boxes have been delivered to the site. The unit is still scheduled for 12/28. The project schedule is still valid.

11/21/06 – We held a construction and schedule meeting with the contractors. EDI has began work in the space on the mechanical side. Demo has been completed. The old AHU has been removed along with its ductwork. Holes have been cut into the walls for the new ductwork paths. Equipment is ordered and scheduled to be here mid to late December. Electrically, 3/4″ conduits have been snake across the ceiling including the voice and data for HamCo communications. One conduit in the mechanical opening for the new supply ductwork has to be moved and will be by QCE. QCE cut a hole in the soffit to facilitate new conduit paths. The GC will have to patch this hole at the end of the job. It was discussed and no access panel is needed at this location. TEI stated that VAV box has SCR control on electric heat and this may be done through a PID loop by Honeywell only if they can provide enough stages of heating to keep the space a constant temperature. Fire dampers were discussed with the code official and they are not required but an engineering dwg would be required to change them so it isn’t worth the savings. The FD will be installed.

10/27/06 – Construction has started. Light fixture submittals have been approved. TEI approved final AHU submittals with EDI and Trane. The units size will be decreased by 1 size allowing it to fit into the space but this increased the face velocity through the coil to slight above 500 fpm (a normal max for airflow). Trane assured us that there would be no moisture drift through this unit. All equipment is on order.

10/01/06 – Pre Construction meeting today. EDI turned over submittals. ThermalTech reviewed and sent back comments. Trane AHU is revise and resubmit. EF, Louver, Air Devices are all furnish as corrected. PL is talking to QCE about down sizing to two 3/4″ conduits across the hallway. Mitch requested a conduit for the IT but we also discussed leaving room on the wall in the MER for a patch panel and have one cable go back across the hallway for easier access. The county has already remediated the space so it is ready for work. TTC instructed EDI to have sprinkler piping in MER room ran close to the outside wall to stay away from new electrical disconnect switches in the space. EDI is supposed to find out who the FPC is on this job and let TTC know.

09/01/06 – TTC contacted the City Plan Review to discuss outdoor air calcs. They have reported that we did not use the correct tables for our ventilation tab. Actually I doubled check this and it is the EXACT table he listed. Perhaps they are misconstruing our use of the transient occupancy rule in both the code and the ASHRAE guidelines. This is where I said the owner may have to provide a letter stating that the space is actually only used at full capacity for under 3 hours on any given day – it is OVERFLOW not full use. Bob Martin called me back yesterday and I am calling him back right now to resolve.

06/15/06 – Bids were opened. Mechanical and Electrical are at budget. Once GC bid and is acceptable also. Partners are waiting on a formal approved bid tab to write the contracts for this project

05/25/06 – Drawings arrived today for this project.

05/22/06 – Spencer has submitted the drawings for bidding. Spencer and TTC discussed bidding estimates and the cost were entered and were below the original estimate. QCR called to say that 2 of TEI dwgs were mislabeled as mechanical dwgs when they were really electrical.

04/11/06 – TTC & Sam discussed project design today in office for layout. Sam met with Mike Bartlett yesterday to find a suitable HHW source for reheat. HHW was so far away that it doesn’t make sense for the space design we are using. TTC discussed zoning of the space today with Mike and decided the best and easiest way to serve these spaces based on our learning’s in the 450 Mediation space is to provide two zones with averaging thermostats to keep areas about the same temperature. Ease of use, easy to maintain and both zones can be kept in the MER adjacent to the air handler for ease of maintenance as well. TTC has to confirm with Mark Donnelly and Mike Bartlett economizers for this air handler on Wednesday.

04/11/06 – Final design is being put on the drawings. It appears that electric reheat will be used. TEI will break the room up into 2 zones total. One for mediation and one for the Jury commission. We will use averaging thermostats in the different rooms to adjust the air flow and temperature of the unit serving the space. This is a modular unit with electric reheat on the VAV boxes. A cooling coil and steam coil will provide the space and supply temperature to the main ducts. The two VAV boxes can be located in the new MER we are using for this job.

04/05/06 – TEI has been coordinating with CR and the owner extensively on the requirements of the new space. An owner provided sketch has created a higher density floor plan that Spencer sent to everyone. CR and TEI decided that ductwork can be ran in the space if the ceiling is lowered to 8′-0″. This will allow TEI to run supply and return ductwork in the interstitial 18″ of open space. A modular unit with direct outdoor air can be used to provide VAV boxes with supply air for the space. The radiators can be used for heating in the winter time. Mike pointed out that the reheat system is off during the summer so no reheat is available in the space. TEI will take this into consideration when designing the new system.

03/14/06 – Second Design meeting yielded a more formalized plan. CR will provide a mirrored floor plan (MS suggestion) and TEI will have mechanical and electric work done by April 14th for review with the owner. TEI will most likely replace the lights and the HVAC unit on this project.

03/07/06 – First Design Meeting. Scope defined for space but we need direction from owner on mediation’s part of the space and actual tenants requirements.

by TEI TEI

CH – JAX Security Camera Upgrade

12/13/07 – TEI received all the closeout paperwork, O&M’s and final pay applications earlier this week. MH, TTC has reviewed and approved. TEI will have these documents sent to Bert Watts for final closeout on this project. TTC sent an email to Copp asking about the remaining project documents for closeout because the Release of Liens, Surety, PW Affidavit were not included.

10/18/07 – TTC received and sent BW certified payroll but no closeout paperwork has been received.

10/04/07 – Terry and Bert still have not received closeout paperwork and payapps for this project. CSI is working on it.

09/03/07 – Bert sent back the final payapp because schedule of values is not itemized. TTC emailed Paul Wilson at Copp and requested new documents and signed allowance usage. TTC sent the allowance usage again by email and filed the final pay application in the HamCo Hold bin.

08/16/07 – TEI sent all final paperwork to Bert. Copp said they would finish and send their final paperwork to Bert.

08/15/07 – TTC reviewed the final payapp and closeout paperwork. Most of the closeout documents look fine. We have not received any O&M’s for approval on this work. I also don’t believe we have a signed allowance usage on this job. We also do not have a signed Punchlist. TTC sent an email to Paul Wilson outlining all the deficiencies on the project closeout scope and told Copp to send the O&M’s directly to Bert Watts. TEI will send the final signed payapp and paperwork to Bert on 08/16.

07/03/07 – TTC contacted Copp systems about closeout paperwork and job forms to use. TEI gave them until July 13 to get everything to TEI for approval.

05/14/07 – Copp submitted a final payapp on this project but it was rejected by ThermalTech for various reasons. TTC emailed Paul the correct document format and closeout paperwork for resubmission.

03/20/07 – Project ready for punchlist. Only training remains.

03/14/07 – Copp stated the the project is complete including our change order work. TTC has to punchlist this project.

02/19/07 – Paul Wilson has the new control panel in. The countertop is being refinished.

02/02/07 – TEI has not heard anything back from the County on the status of this project.

01/22/07 – Sketch was issued in PDF and DWG format to Bert Watts for Dave Spitznagel.

01/10/07 – TTC will issue a sketch based on the field trip on 01/08. TTC also to get prices to add two cameras to day holding cells on 5th floor from Copp. TTC to coordinate new cabinet dimensions with Facility trades.

01/08/07 – Work has progressed to the point where the monitors are installed on the countertop. They are really large and Mark would like to look at changing the CRT monitors to LCD monitors in this project. Dave Spitznagel has the new countertop built that might be skinnier than the old one. A const mtg will be scheduled for Wednesday.

12/28/06 – Paul Wilson called today to inquire about a quad receptacle behind the DVR recorder for this project. He could not find one on the drawings and we all believe there should be one. TTC instructed them and in talking to MWH he agreed that one should be installed. MWH also question why we didn’t spec LCD screens on this job. TTC could not answer as MS did the design but it does seem kind of odd.

12/26/06 – TTC created project journal and filled in back data from previous meetings.

by TEI TEI

CH – SECURE BASEMENT HALLWAYS

10/18/07 -MD sent back final payapps for record on this project with signed change orders.

09/04/07 – The final close out documents where sent to Bert Watts this morning. These items include final pay application, close out documents, O&M manuals with as-built drawings.

08/31/07 – TEI received signed CO#01 from T.J.Williams. Final pay application was reviewed along with all close out documents. All of these items and the O&M manuals will be forwarded to HCFD on Tuesday, September 4.

08/27/07 – TEI sent final change order for signing to TJ Williams. Bert inquired and we told him they were already in the mail last week.

08/22/07 – TJ dropped off as-built dwgs from Copp on the security system work for the project. TEI will integrate into the system and asked for acad dwgs from Copp to do this, not just the hard copies.

08/14/07 – TEI received the modified final pay application along with the signed Change Order No. 01. NR forwarded pay application No. 03 and Change Order No. 01 to Mark Donnelly. TEI still awaiting the closeout documents from T.J. Williams. An email was sent to Diane Vogelsang of T.J. Williams with a list of the final closeout documents on August 8, 2007. NR will contact her again to make sure it was understood that final payment would not be issued until all of these documents were received and forwarded to HCFD.

08/07/07 – TEI received the final 2 pay applications from T.J. Williams. The close-out paperwork was not included with the final pay application. NR will contact them to let them know what is still needed.

08/02/07 – The system was tested and passed last week. T.J. Williams will be submitting the final pay application today. All of these items are being sent to TEI. The only remaining item for TEI is the as-built drawings. T.J. Williams has been told to send the drawings to TEI so that an electronic version of the drawings can be made turned in the HCFD.

07/30/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/29/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/22/07 – TEI still awaiting word on the installation of the final 2 stair treads.

06/15/07 – TEI is awaiting word on when the stair treads have been installed so that the system test may be performed. The treads are the only thing holding this project from being closed.

06/08/07 – There are two stair treads that are not installed. When these treads arrive and are installed there will be a system test scheduled. When everything is proven to operated as intended this project can be closed out.

06/01/07 – NR performed a punch list today. The items on the list will be issued today. The stairs on the north side of the building are currently being installed and when they are finished a function check of the security system will be performed by T.J. Williams. Once this is complete and the punch list items are finished the project will be ready to close out.

05/25/07 – TEI will be on site the week of May 28th to complete the punch list.

05/14/07 – Project is complete.

04/27/07 – Issues with the lock hardware has been rectified. Work is progressing as expected.

04/20/07 – Work is progressing as expected.

04/13/07 – Work is progressing as expected. Wiremold issues are still being addressed.

04/06/07 – Work is progressing as expected.

03/30/07 – Work is progressing as expected. Wiremold issues are being addressed.

03/23/07 – Project progressing as expected.

03/13/07 – A construction meeting yielded nice progress. The sprinklers are complete. TJ will use angle to support the new lights in the structure. Bulkhead issues were discussed and Legacy was instructed to have Ken Jones contact Spencer directly for project direction. TTC & PL attended this meeting.

12/04/06 – A pre construction meeting was held. PL is handling the scope on the EE side. TTC will approve the FPC dwgs for the job when they arrive. City requested response from the Architect which has been sent. No submittal data has been received yet.

09/29/06 – Waiting on contracts to be returned from County.

09/22/06 – Contracts have been written and forwarded to county. T.J. Williams was the successful contractor

09/08/06 – Prebid held. Contract is in progress.

09/01/06 – Project was rebid. New prebid was last week. New bids are expected on this job.

by TEI TEI

CH – 6TH FLOOR RENOVATIONS – PHASE II

05/11/08 – TTC faxed and emailed the original approved permit to Bert Watts at the County.

12/20/07 – TEI gave Honeywell a PO to do the FA work on the 6th Floor space. All of this work was not shown on the dwgs and Honeywell prepared a separate PO for the work and gave it to Bert. TE took it from there and will coordinate the installation to make sure it matches our drawings.

11/30/07 – This week we ironed out the Fire Alarm permit and drawing scope with Honeywell for this job. In the future, we (TEI) should be more proactive in getting this work coordinated with our work but I admit that with both Allgeyer and Richmond leaving that I let this little permit issue slip through the cracks. I appreciate Bert and Mike B helping me out on this issue with this project – TTC.

10/17/07 – CRA emailed correction for door hardware access controls code.

10/04/07 – The plumbing permit has been given to the County, the HVAC permit is approved and was emailed to the County, TTC to delivery hard copy on Friday. IBI permit is coming on Friday and then County can call for inspection.

10/01/07 – Building permit approved. CRA delivered the permit to the County Facilities Department.

09/19/07 – TTC sent the HVAC permit to the County and CRA. Don Grace called and said the plumbing permit was approved and given to Bo Blanks. He said there was one small change. TTC contacted 3 FPC firms to move the 4 sprinkler heads in the Phase 2 space. J2 called back and is ready to do the work. TTC gave this information to Mark and Mike to use whomever they wanted. The sprinkler work will not require a permit according to the contractors.

09/10/17 – TTC finished the latest drawings with Friday’s comments. TTC sent them to MD and BW and will print out the permit sets for Tuesday.

08/31/07 – TTC dropped of hard copies of the drawings to Mark Donnelly for his review. TEI is awaiting word back from him on how to proceed from this point.

08/23/07 – TTC sent plumbing fixture cuts to CRA and HamCo for approval. Our floor plans are finished and ready for review. TTC still has a visit to the CH for underslab field work with MB.

08/20/07 – CRA & TEI met with MD to discuss finalizing this project. New project direction was given to the partners. Some of the items covered at this meeting were: Paracubes will be used in the new office area, deep parabolic will not fit. The unisex bathroom has the WC and LAV flipped and TEI can coredrill the floor for the WC and new FD. We will use a floor mounted tank type WC for this job. The refrigerator will need a drip pan that will have its tube ran over to a trap primer connection on the FD. On the HVAC system we discussed removing all the VAV types and just ducting to the existing branches.

08/17/07 – TTC and NR conducted a site visit to review the work area in regards the the changes made to the plans. These changes will be discussed next week.

08/10/07 – TEI will have drawings ready for owner review next Friday, August 17, 2007.

07/17/07 – TTC discussed with MD via email today about MEP completion on this project. TTC agreed to be complete in 3 weeks with the project 2008 project use. CRA sent over a furniture plan for TEI to use for design criteria on this project.

06/29/07 – At the monthly Partnering Meeting with Hamilton County it was decided that the money to perform this work would be budgeted for a 2008 project. TE was directed to complete the design and turn over the drawings to the Hamilton County Facilities Department.

06/15/07 – Electrical field work and design on this project has been completed. Awaiting HVAC work for this project to be issued for owner review.

06/15/07 – NR created project journal and filled in back data.

by TEI TEI

CH – GENERATOR REPLACEMENT PHASE I

02/21/08 – ThermalTech approved the generator submittals from DK on this project today. GM reviewed the Generator submittals and approved them “No Exceptions” 2 copies were mailed back to DK, Wally Gore. 1 copy was sent to Bert Watts and 1 copy was retained by TEI. TEI printed out 6 sets of stamped mechanical dwgs for Triton today and sent them to Grady Reed for permit application on this project.

12/11/07 – We have had a pre construction meeting with Debra Kuempel and Triton on this project

11/30/07 – At last report, this project was in contract approval stage with DK getting the contract on this job as the base bid winning contractor. TEI is waiting to schedule a pre const meeting for this job as soon as the contract is in place and ready to go.

09/10/07 – MD gave TEI the OK to write the contract for Debra Kuempel and Triton Services on only the base bid work for this project. TTC & LKV will write the contract on Tuesday.

09/03/07 – Bids have been opened. MD asked for company portfolios to assist in helping him decide on which contractor to use on this project. TTC called and requested project portfolios form both QCE and DK for the electrical. Triton Services appears to have the lowest price for the mechanical work on this project. TTC sent an excel bid tab to Mark to use to compare total prices on the project.

08/22/07 – TEI finished Addendum 01 to the project which constitutes all the Alternate 01 work on the project. New fuel oil piping system and new fuel oil pumps on the project. In emails, MD has directed us to use manual fill for the belly tank on the job, accept the substitute vendors who may be rejected due to specifications, double contained pipe in the trench, keep the manual fill station for the belly tank. All of this has been incorporated into our Addendum document. Bidder will have to pay and pick up drawings from Queen City Repro for the addendum.

08/08/07 – MD called to instruct TEI to issue and addendum with all the Phase 2 generator work included. TTC brought up quite a few points about this being tough to do at this point. TEI wants to discuss in further detail with Mark. Prebid is next Monday.

07/31/07 – TEI is finalizing the specs and the dwgs today. As with most projects, during the last 10% of the design we find 90% of the info we have needed to do the job all along. TTC has to add generator exhaust insulation to the piping spec and reformat all the PDF files and check the table of contents before sending it to purchasing for bidding. The job is complete. TTC answered MD that the belly tank is 435 gallons, the smallest available. 300 gallons is good for 8 hours so the new tank should last for 11.6 hours.

07/27/07 – We have finished the dwgs and specs and had the TEI QC check performed. After some final changes on Monday, TTC will send to purchasing for bidding.

07/23/07 – TTC requested ITB and SBE numbers from the County. TEI to have final documents ready this Friday for bidding. Advertising and documents should be at Purchasing by Monday July 30th according to TTC email to team on 07/23. TTC & CB to finish mechanical design on this project.

07/16/07 – JP and TA had final review mtg with MD today. New generator is being pushed back until later. For now we are going to install a replacement generator of the same size in the same space. Project is schedule to be completed by next Friday. TTC will assist in the design.

06/30/07 – A review meeting has been set up between TEI and Hamilton County Facilities Department for Monday, July 2nd.

06/15/07 – TEI is awaiting answers from the Building Officials as to what equipment needs to be on the new generator.

06/08/07 – TEI has continued investigating what equipment will be on the new generator.

05/30/07 – TEI coordinating with IBI and building department about who is going to decide if the fire pump has to be put on the new generator. No one is claiming coverage for this piece of the puzzle.

04/27/07 – Loads are being reviewed by TEI. Project is moving along as expected. TTC & TA visited the site and did some additional field work on the space. One concern TTC noted is that the existing Halon system will have to be removed and upgrade if ANY changes are made to the room. According to code officials and inspectors that Halon system must be removed if it is touched. This may cost an additional $10,000 to the project..

04/20/07 – Project is moving along as expected.

04/13/07 – Kickoff meeting was held this week. Project has moved to Design Phase.

04/04/07 – SOSS has arrived signed with a PO. TEI will schedule a kickoff meeting for this project.

03/30/07 – Project still awaiting signed SOSS.

03/23/07 – Project awaiting signed SOSS.

by TEI TEI

CH – ROOM 340/360 HVAC UPGRADE

04/10/08 – Triton submitted closeout paperwork on this project. TTC review and approved. TEi has to write a change order deduct for the allowance money not used on this project and will send that to Triton for processing on this project.

12/28/07 – Grady asked about payapp # 2 and TEI reported that punchlist, O&M’s, payapp 1 have all been received and sent. To closeout this project we need closeout documents, final payapp, certified payroll and an as-built dwg.

12/21/07 – O&M’s have been received by TEI. TTC approved them for distribution to the County. We are sending them to Bert Watts. TTC asked for an additional airflow balancing on the main units themselves for maximum flow with the VFD at 100% to included later.

12/19/07 – JS reported today that the final punchlist work is complete. TEI can now backcheck the punchlist and closeout the job. MB asked about allowance money remaining for moving a piece of equipment at the CH.

12/12/07 – John Sanders faxed over a final signed punchlist to TTC for review yesterday and reported today that the flow switch cut outs have been approved by the code officials. They will now get them installed into the units for final inspection and approval by the city.

11/30/07 – Triton and Honeywell have finished all their punchlist items but the project did not pass final inspection because of the new mechanical code requirement about secondary condensate drainage under cooling coils. This requirement is listed under the General Regulations Section 307 of the OMC for 2007. TTC sent scanned copies of the code to Mike B and Grady R for assistance in getting the permit and finishing this job. Triton reports, “We are getting documentation from carrier as to what they recommend as an overflow protection device to shutdown in the event there is water present outside of the drain pan. The inspector is ok with this as long as carrier is ok with the install – Grady Reed.

10/17/07 – John Sanders called today to tell us that Triton is lowering the piping on AHU-360 and will coordinate that with MB, bringing out a sheet metal cover for the insulation touching the door on CR 360, moving the CHWS label to the correct line and that they have finished the painting of the doors which I told him was fine they way they looked now, . The controls are still being finished

10/08/07 – TTC sent a revised punchlist with picture email to Grady based on his conversation with Mike Bartlett and Willie Schadler. Triton is trying to finish this project but balancing and controls remain to be done.

10/0407 – County said AHU piping was not switched on original air handler. They may mean the piping blocking the filter rack on the second air handler. Regardless, Triton has to change the piping to match the design.

10/01/07 – Triton has stated that have finished punchlist work and controls are finishing their. Payapp # 1 has been received. Grady has the two VFD’s that TEI is supposed to purchase for our error on this job. Units are installed and working very well as far as I know.

09/12/07 – Triton and QCE are on site performing punchlist items. Honeywell has final control parts to install so system can begin running the VFD’s automatically. Randy from Honeywell said he should be able to install the components this week. TTC will check back on progress next week. The CHW piping has not been relocated yet.09/10/07 – TTC inquired about the status of this project with contractors today.

09/03/07 – TTC finished and sent the punchlist to Triton for this project. Controls are not finished and VFD’s are still in manual causing some overcooling of the space. I am confident that the systems will work when complete.

08/22/07 – Both units are now installed in the space. There were some problems on the initial installation but they have now been worked out. The VFD’s were ordered at 480V when the building really has 208V at the units. Triton quick shipped 2 new units that week and the 340 unit was running the end of the week. Air noise and unit noise has been drastically reduced with the new unit but the motor was making a hum that was fixed by changing the tap voltage and reprogramming the VFD. The County asked for new isolation valves in back of the old ones and Triton has provided a change order cost for this. TTC forwarded to County and is awaiting a decision on allowance usage. The CHWS/R piping was connected backwards to the 340 units but Triton flipped it during the 360 changeout work. The 360 changeout work went much better but MB reported that the new CHW piping interferes with filter rack removal. Grady said that Triton will take care of this. MB reported that the motor noise was back and that the judges both complained. Grady had the VFD rep make some changes and thought it was fine. TTC will contact MB to check. TEI is waiting until the controls are finished before punchlisting the job which should be next week. We have coordinated this with MB. TTC told Triton that the dwgs do NOT specifically call out for the steam trap, isolation valve and strainer to be replaced. It was a lineweight issue with Queen City printing that showed them dark.

07/31/07 – TTC recommended the 340 AHU piping be changed during the 360 shutdown weekend work. We can coordinate it so that we do 360 AHU first and if there is time left we can do the 340 AHU. QCE will also change the 30A fuses to 25A during the shutdown. Carrier will investigate the harmonics of the motor noise to get that removed from the unit. AHU seems to be cooling the courtroom just fine.

07/30/07 – NR verified the fuse and wire sizing for QCE. There only modification from the original design is that the disconnect switch frame is larger than intended. The fuse and wire size match the design so the frame size is acceptable.

07/27/07 – After a week of tribulations, the 340 AHU is up and running. There are several issues that have happened on this project that have this changeout tough. TEI spec’d the wrong voltage on the VFD and had the CHWS and CHWR labeled backwards on the unit. Triton did not have all the material they needed. The unit is whisper quiet but the motor makes a funny noise that I think is not right. We are working on resolving these issues and will get this unit checked and balanced before we changeout the 360 unit.

07/23/07 – Triton informed everyone today that TEI spec’d 480V VFD’s while the correct supply voltage to the units is 230V. TTC checked the drawings and we do call for 480V. We don’t know how this happened because we checked this several times during design and coordinated amongst ourselves. Regardless, we have to find 230V VFD’s from Carrier immediately to keep the project moving forward. The old 340 air handler has been removed and Triton was to remove it from site yesterday. The new unit is installed but not powered up yet. Triton is working on the new drive as we speak.

07/18/07 – Triton submitted a $2300 allowance charge for replacing all pneumatic actuators with electric actuators. AKM approved and TTC sent out the paperwork and approved the change. We are scheduled to have a coordination meeting on Friday with Triton for last minute changeout details for the new unit.

06/25/07 – We had our first construction mtg yesterday with our new project manager Grady Reed. John Sanders is going to be our field foreman on this job. We walked the site and Triton is now putting together a schedule and employee list. Mike is checking on best weekends for the judges. TTC is going to distribute the meeting notes for the project.

05/30/07 – Equipment is scheduled to be delivered in mid-June for immediate installation into project.

05/17/05 – TEI finished the control submittals and sent them back to Grady Reed of Triton Services.

03/15/07 – Submittals have been returned to Triton for ordering on this project. Controls asked about 120V power and NR answered via email that 4th floor is fine.

02/21/07 – Pre Const Mtg yesterday. Meeting notes today. PDF

09/01/06 – TTC met Tim Centers on site today to discuss fitting the unit into the space. The only door from the hallway has a maximum width of 35”. The unit must be completely broke down and re-assembled in place. This is possible with the Carrier 39M series although the entire unit must be taken apart. Habegger reported that the unit will be quieter especially since the fan and all associated components are located inside the unit. We are buying a VFD with the unit for speed control. Also we discovered today that the unit when constructed is tall enough so the fan discharge will match up with the existing supply ductwork. There is no need to put a supply elbow on top of the unit. We are also going to keep the existing return and OA as is and butt our unit up against the existing filter rack. There is just enough room for the transition to fit. TTC went over the controls with GD today. The chilled water valve and actuator are new. The steam valve and its actuator may be replaced. The OA damper will remain in use as is.

by TEI TEI

CH – MARBLE STAIR TREAD REPLACEMENT

12/03007 – CRA Punch work completed.

12/05/07 – CRA final pay app and c/o sent to Hummel.

11/14/07 – Awaiting Hummel completion of punch work this week.

10/01/07 – Awaiting Hummel completion of punch work. Awaiting closeout documents. Need As-Built drawings.

09/19/07 – Work completed 9/12. Punchlist issued 09/19.

by TEI TEI

COURTHOUSE Generator Replacement Project

07/29/09 – We have not received final closeout documents from DK on this project yet. A couple of weeks ago, Bert reported that DK was working on closeout and had all the permit requirements to close this project out.

06/22/09 – Triton has sent everything for closeout on this project. DK got all their final permits and approvals, now we just need the remaining closeout documents.

06/16/09 – DK sent in O&M’s and George approved them. We sent them to Bert. TTC called Pam today and DK and said all final payments will be sent to Bert this week. We have not received the Triton closeout docs yet but TTC inquired with Triton about them.

05/22/09 – We are still working on closeout and permits. Jeff Jones approved the configuration and gave it a UL stamp. DK can now get final inspection and complete the project.

02/19/09 – Non Conformance letter from IBI is top issue. GM is working on. Bert wants a separate SOSS to upgrade the building later.

02/02/09 – Final permit sent to Bert Watts last Friday.

01/05/09 – TTC contact DK and Triton about closeout on this project today. DK needs to fix their invoice #3. Triton needs to send in final payapp even though all closeout paperwork is received. TTC sent allowance usage forms to both contractors for $2000 and $5000 use on this project. DK emailed lots of questions about how to fix billing. That is worked out. Bert asked for final inspections to be completed soon and for Triton to get theirs done.

12/20/08 – Bert reports that most of the work is complete. The XFER Switch component is being ordered and special made at the factory.

11/25/08 – DK said work will be starting next week and should wrap up soon.

11/24/08 – TTC contact DK about closeout.

10/16/08 – Debra sent in change order numbers that equal $5000 (the allowance). These have been approved by Bert and TEI. TEI will process the paperwork for usage and send to DK to facilitate the work occurring at the generator project. We will get insulation on the exhaust piping, a refill on the fuel oil and a part replacement on the transfer switch in the room that was not originally part of the scope. TTC dropped of Kohler O&M’s for this project to Bert by hand. Bert called and said he got them.

09/18/08 – O&M’s received and sent to Bert for Triton. Triton is 100% complete with Permit & As-Builts drawings that are being hand delivered to Bert. DK is missing as-builts and permit and need to provide a quote to insulated the exhaust piping.

09/10/08 – County reported issues from load test include loosened bolts, too hot, heat detector tripped and room overheated. Bert wants inspection certificates and we want o insulated the exhaust piping. Triton cannot provide insulation within project budget. BW advised to proceed with DK. DK sent submitted payapp 2. Allowances usage added to pay app page.

08/29/08 – Load bank test is scheduled for 09/06 at 10am. TTC requested all closeout and inspection forms from contractors.

08/25/08 – Vic at Buckeye responded today that they will set up the load bank test for this project directly through Bert.

08/22/08 – Bert wants TEI to call for load bank test. We contact DK via email.

08/08/08 – Both Mechanical and Electrical punchlists have been given to contractors. TEI has finished mechanical as-built dwg.

08/01/08 – TTC sent out the Mechanical Punchlist to Triton and requested closeout paperwork and allowance usage numbers for the upsized generator exhaust. GM sent out the electrical punchlist for this project.

07/29/08 – Generator mini test and training went well. TEI to do punchlist for project. Automatic Transfer switch acts funny but that was not in the scope of this project

07/04/08 – ThermalTech visited the site as Triton & Debra Kuempel installed the equipment. The entire weekend went smoothly.-

by TEI TEI

COURTHOUSE Clerk of Courts Room 315 Renovations

12/31/09 – Archived on 12/31/09 by TTC

04/15/09 – CR Architecture + Design awaiting Trades to complete work so that punchlist/ walk through can occur.

05/05/08 – Issued for permit.

04/21/08 – Preliminary plans sent to Bert Watts for review.

04/07/08 – Approved SOSS received.

03/27/08 – CR Architecture + Design submitted SOSS.

03/10/08 – SOSS requested.

by TEI TEI

CH – Plaza Lighting Study

02/22/10 – Joe Schamer is setting up the const work schedule and will get back to us soon. He is checking with his sub on work.

07/23/10 – TJ finish field work and sent results to TTE

07/28/10 – TTE send report to Bert recommending fix for lighting fixture 6. The only viable option is to replace the conduit between poles 5 & 6.

08/13/10 – Bert asked for paver cost. TEI to contact JB Schmitt for an estimate.

09/29/10 – New dwgs are updated. THP has a set of markups for their work. Should bid out next week.

10/28/10 – Second prebid schedule 10/29 at 11am.

01/24/11 – Pre Const Mtg was held two weeks ago. As soon as the weather warms up, we will begin our work on the plaza lights.

03/30/11- All work has been completed. TTC will check the wiring covering at the Partner meeting.

04/27/11 – TEI is holding final payapps in the hopes of getting the brass covers installed but hasn’t contacted the restorer after passing along the information. We can make this call directly for HamCo to close this out. We need to discuss with Bert extra charges on this project vs. const cost.

06/14/11 – TTC asked SECO if they have anything in warehouse that would match the light post. Have not gotten a response yet. TTC is checking some stores to see if he can find anything. TTC also came up with a new idea, but needs to check something first to see if it will work.

06/29/11 – Seco has not responded from our last communication. We hope to bring our brass cover down on Thursday to review.

07/21/11 – Spoke to local wiremold vendor. Wiremold offers brass raceway, but is a custom product with 250 ft minimum order. Local rep suggested buying standard product and field painting to match.

08/28/11 – We have purchased the materials, now we just need to install the new wiremold.

12/23/11 – Project will be archived for 2012.

by TEI TEI

CH – Room 315 Data Center Unit

10/06/11 – Glen Wilhoit has asked TTC if there is a scope for this project yet. TTC asked Tony M and Bert W what the status of the SOSS is so he can reply to Glen’s question.

12/23/11 – Geiler won the bid. They are working on submitting hard copy bid documents to the county including the performance bond which they state will be over to the County the last week of the year. The contract is waiting to go over to the prosecutors office after final bid documents are received. ThermalTech will accept either horizontal or vertical unit replacement on this project and Geiler is to submit either at the pre-const meeting. There is no allowance on this project at this time because we took the PO out for the bid amount late in the year.

03/27/12 – ThermalTech received submittals from Geiler for their proposed solution. ThermalTech to review the submittals, and issue a Change Order for modifications to the electrical requirements per the submitted equipment.

04/06/12 – Permit was submitted and TTC asked Geiler to call for any inspections needed after the City approves the permit.

04/12/12 – TTC asks Geiler to have their field personnel move the humidifier back to a wall or install it in a better service orientation. Glen at the County said the plenum location their field man installed it behind the unit and he will not be able to service ti in the future. We show it on the wall in a remote mount installation – please get with your man and see if it can be fixed for Hamilton County. Mike Folck can be present to meet with in the field to help coordinate. Geiler sent a suggestion for the problem to ThermalTech. TTC responded that it was acceptable provided – 1.This location does not void the Carrier warranty in any way (mounting equipment on their case). 2. This location does not interfere with inner workings of CRAC unit. 3. This installation meets NEC requirements for electrical. 4. Greg and Glenn give their blessing too.

04/20/12 – The system is up and running and ready for an inspection. ThermalTech will punch this job out on Tuesday (4/24) after the Ameresco meeting if okay with Greg and Tony.

06/20/12 – Close out documents were sent for this project. The remaining items to take care of are: Only 1 O&M was sent, specs require 3 (TE will do the other 2), warranty letter, contact sheet for warranty or service calls, signed punchlist from Geiler, As-built markup.

Payapp #2 is approved and sent to Tony Matre. Payapp #3 is on hold pending the receipt of the above mentioned items.

06/25/12 – All closeout documents received, approved and hand delivered to Tony at Monthly Partner Mtg. ThermalTech completed O&M manuals for Geiler and inserted As-Built, warranty, contact information along with standard equipment care.

09/28/12 – TTC archived this project after the PM meeting.

by TEI TEI

CH – AHU 566 & 580 Replacements

02/22/10 – TTC has to meet with Greg to discuss additional AHU replacements in the Courthouse.

03/30/11 – As soon as the new contract is ready, we will be submitting an SOSS for this work.

05/06/11 – SOSS submitted for approval.

06/29/11 – Kickoff meeting

07/07/11 – We met with Spencer to gather some intel about the existing soffits in the Courtrooms this morning. GBBN knew that these soffits are left over from the old plumbing chases on the 6th floor jail and have always been part of the building. That makes a lot of sense. TEI will believes they have to be incorporated into the new design and has sent diffuser suggestions to Tony for approval. Habegger has been contacted for vendor sizing and budget quotes for the wor

07/22/11 – TTC and MAF completed fieldwork for demo of existing AHU’s and associated ductwork. Also completed fieldwork for layout of new AHU, VAV’s to serve two courtrooms / judge’s chambers on the 5th floor, and routing of hot water piping thru main hallway. Habegger has submitted two options for replacement AHU. Unit will have to be disassembled and rebuilt to fit thru existing doorway on the 4th floor.

08/09/11 – MCB and MAF completed fieldwork for demo of existing electric feeds to AHU’s and fan coil units. Also completed fieldwork for layout of new electric feeds from 5th floor south panel to new AHU, VFD, and VAVs.

08/22/11 – Mech/Elec drawings and specs in the process of TEI internal review.

09/30/11 – Review set given to TTC today for approval. TTC will have ready for bidding and request ITB for public bidding in about 2 weeks.

12/23/11 – Contracts have been approved by HamCo. The project will start in January with a pre-construction meeting.

02/21/12 – TTC has stamped 6 sets of mechanical drawings that will be delivered to Debra-Kuempel for permitting. NV is stamping 3 sets of IBI drawings that will be delivered to QCE. Proceed with permitting this week.

02/23/12 – Lead paint was found in MER. SOV approved for both contractors. Unit will be 80/80 for electric and QCE is okay with this size and feeder wires. No charge.

02/24/12 – DeBra-Kuempel will not disturb during their demo, but SCS construction will have to saw cut the new opening in for the door where the lead paint is located and the electrical and controls contractor may have some conduits, boxes, clamps etc. mounted on the same wall. They were not sure what the industry standard is for removal or remediation. Tony Matre may be able to answer some of the questions as to how the county will proceed. DeBra-Kuempel has everything going through with the background checks, AHU ordered and the permit has been submitted.

03/27/12 – During demo, Queen City found unknown circuits fed from the existing 566 fan coil unit location. These may be feeding added outlets or lights within the judge’s chamber. Depending on the use of these unknown circuits, we have the option to demo, leave in place, or re-route to the appropriate location. Mike B will be investigating further with QCE and update on this issue.

03/28/12 – Mike B met with Tim Ziebold on site. They traced the conduit and it appears to feed an outlet (with nothing plugged in) on the wall in the judges chambers. They were not able to determine if there were any downstream devices connected to this circuit. Thermaltech spoke with Greg Jennings and the plan is as follows: At a time when the courtroom is not in use, and the Judge has left for the day, we shall disconnect the power to the circuit feeding the branch circuit that runs through the disconnect. At that time, we shall go into the chambers and confirm if the circuit feeds the judges computer (via a light on the power strip). If no lights or equipment in the chambers or courtroom are affected, the per Greg Jennings it is OK to leave off at this time. Feed is to re-routed as follows. Intercept feed at corridor wall (north wall of mezz), provide a junction box above door. Route feed east on the north wall to the east wall, turn south and then down to pick up feed where it feeds up, reconnect and turn back on. Any calls or questions should be directed to Mike Barth.04/10/12 – We call for the sprinkler line to remain in the mechanical room on your 566/580 project. After the new ductwork is in and as a matter of fact even with the old ductwork configuration, that one sprinkler might not give you optimal coverage in that room. I spoke to Debra about note #7 on M200 which tells them to work around or modify that sprinkler line under their contract which will leave it in the same place. After they get the ductwork layout done, we may want to address “better” coverage of that space by extending the head up and running a horizontal line over and changing that to two heads that spray down. That was not in their contract to do but I thought we should at least discuss it now. It may come up on final inspection by the building dept if they pay close attention to this one detail. TTC asked Greg who was doing most of the sprinkler repairs. TTC also tells Greg to feel free to direct all engineering type questions to Folck, Barth or himself if he does not know the answer. TTC stopped in on the contractors and answered a few more questions and went over project progress with Clark and Greg (HVAC & Piping).

04/11/12 – The following is a summary of questions from Debra-Kuempel and clarifications regarding this Project:

– New supply/return ductwork to allow for clear access to 5th floor platform and VAV’s through existing doorway. VAV’s to be installed at working height.

– Existing transfer grille about 5th floor driveway is not necessary for current operation. Damper should be closed off and left in place.

– Existing 580 Judge’s Chamber flex return combines with 580 Courtroom return ductwork. No new ductwork needed for 580 Judge’s Chamber return, leave as is.

– New diffuser locations in 580 Courtroom were not spaced evenly in ceiling bulkhead due to existing structural supports, piping and wiring. New diffuser locations are acceptable and were approved in ceiling bulkhead.

– Existing sprinkler line within mechanical room to be worked around or relocated if necessary per note #7, Drawing M200. Existing coverage within mechanical room may not be adequate. Thermal Tech will address final sprinkler coverage with Debra-Kuempel and County when ductwork layout is complete.

– Location of existing chilled water isolation valves are unknown. Debra-Kuempel is tracing chilled water piping to find nearest location valves. Mike Folck attached pdf’s of 5th floor chilled water piping layout. Existing valves are not indicated.

– New hot water supply/return piping to be routed above ceiling south side of hallway. Proposed routing by Debra-Kuempel is acceptable and was approved.

– New air handling unit to be placed along west wall of mechanical room, allowing space for existing/new condensate drains.

– Total external static of air handling unit is 2.96 in wg. Clearance beneath unit is expected to be 4″. This should allow sufficient space for condensate trap without raising unit or removing part of existing floor.

– Thermal Tech/THP will be providing Debra-Kuempel with clarification drawing regarding new 5th floor platform. Debra-Kuempel will provide new railing along back wall openings for fall protection.

– Building permit not approved by historical conservation. Thermal Tech has contacted historical conservation for status update.

– Direct all engineering questions or comments to Thermal Tech. Contact Greg Jennings for all shutdowns and building access.

– Thermal Tech will be setting up a construction meeting for next week (4/16-4/20) once all availability is known.

04/16/12 – Terry spoke to Tony about Const Mtg setup for Tuesday. TTC attend meeting and take notes.

04/25/12 – Permit to relocate the sprinkler in the mechanical room has been approved. S.A. Comunale is the sprinkler company for the courthouse and has been quoted not to exceed $1,400 to relocate it. TTC questioned if the old head could be grandfathered in or if it needed to be changed. TTC offered to call and discuss if necessary.

05/14/12 – TTC sent out the Construction Meeting notes #1, #2, and #3.

05/30/12 – QCE on site to connect external VFD bypass switch, VFD and AHU this week.

06/25/12 – Initial punchlist complete. Inspections underway for final. Sprinkler work scheduled for this Thursday. Smoke detectors and other punchlist items to be addressed soon.

06/28/12 – TEI did walk through and the items that need to be addressed are:

1. Unit is not started. Start unit and provide all startup reports in O&M.

2. Unit is not programmed. Programmed unit to TEI and Hamco Sequences issued in the submittals. Commission unit to meet all published sequences. Demonstrate to owner unit operates as scheduled. In particular that these units operate on local occupancy sensors then by override when needed. Outdoor air is used for economizer only and VAV boxes work to representative zones. VFD on unit programmed and operating according to sequences.

3. Pain door, install hardware, lock door to master key system given by owner.

4. Touch up the condensate drain work. The hub has been cut too low. Instead only use 45 deg ell at outlet or have cond drain PVC pipe 45 into hub. Put standard hub back so water level does not crest over top of lip. Maintain air gap but don’t cut the hub drain down to achieve it.

5. Install fire sprinkler back into room where removed.

6. Conduct all inspections and test. Get final approval permits from IBI and City. Is the smoke detector missing a base or is that the full installed device?

7. Seal the wall above the existing door with fire stop where conduits and piping penetrate.

8. Clean mechanical room.

9. Clean staging area. Remove masonite, clean floors, tough up walls. Restore room to prior to project conditions.

10. Remove walk off mats and door signage.

11. Begin closeout paperwork process.

12. Provide As-Built markups for work and changes.

06/29/12 – Pretest did not happen, technician spent time repairing shorts on the system from demolition of the old systems. He has everything repaired and new parts installed and programmed. Monday the 9th is scheduled to finish checkout of fire alarm system and DDC controls on the air handler. After the test is completed the Inspectors can be scheduled.

07/03/12 – DeBra Kuempel has asked when the final for the smoke alarm test can be set up. Randy Waddell with Honeywell has verified the pretest. Joe has informed them they would like to use the courtroom on Monday 7-9. Randy is completing the controls, and our air balancers will be in on Friday 7-6 to balance. The Fire alarm permit hasn’t been submitted yet for the AHU project with the City. This has always been an easy miss because it is one little blank on the Bldg and Inspections submit App. How can this be rectified prior to Monday? Did Larry say anything about the final test because typically the fire marshal is usually here for those tests and I see on the city permit website we are still in the “not approved” section for that little bit. Reasons for items not approved: The following table lists items not approved along with the most current action taken on the Item. Items may be linked to a history of prior comments. The list of comments can be viewed by clicking on the item like if present (920-Final HVAC Inspection6/19/2012 Partial approval install ok, need test of duct detector, tied to FA system. Queen City spoke with the city building inspector a couple of weeks ago and asked if we would need to have a fire alarm permit. They explained the specific scope of work. He said that all he wanted to see was unit shutdown in a smoke condition. A separate fire alarm permit would not be required. We already have our IBI final. Now that the pre-test has been confirmed, we should be good to go.

07/18/12 – Queen City is working on closeout documents. Punchlist item 1.03 regarding the flexible conduit in question is the control’s contractor’s responsibility. It will be removed by the end of the week. We will have everything over as soon as we can.

07/25/12 – QCE emailed electronic O&M and AS-Builts to ThermalTech on 07/24/12. QCE to forward closeout package ASAP.

09/19/12 – DeBra Kuempel will drop off O&M’s to TEI today. TTC already has the As-Builts and will hold them until Honeywell’s work at the CH has been approved.

09/27/12 – TTC dropped off two 50ft extension cords for Rich to give to Courts to replaced the ones that ended up missing from the construction work.

09/28/12 – TTC discussed closeout with Rich and Tony. The County seems fine with the revised sequences and temperatures. The VFD is running at 20% minimum when occupied and will ramp up when occupancy increases. TTC delivered the O&M documents to Tony Matre by hand on 9/27. This project can be paid out in full.

09/30/12- TTC archived this project on 9/30/12.

by TEI TEI

Courthouse Window Replacement – Consultant to Ameresco

04/25/12 – Project to be archived. THP completed window wind load code study and provided recommendations to GBBN.

3/02/12- Window Removal schedule with HGC to remove existing window in Mechanical Room on Second Floor. Only partially removed but evidence of materials and installation gathered.

03/12/12- Meeting with Ameresco, Facilities and GBBN to discuss status of project.

03/15/12- Meeting with Tim Davis of Aluminum Resources and Beth Sullebarger of The Preservation Group, LLC to discuss historic preservation requirements and scheduling.

03/28/12- Window Mock-up meeting with County to discuss design features of windows and mock-up for presentation to Judges and Building Committee.

by TEI TEI

Coroners – HVAC Replacement – Project Status

05/24/05 – TE emailed out all the field data collected when helping trying to solve an airflow issue with AHU-3 to BS, AKM and MJ. PDF

05/20/05 – TTC and County discussed flow and pressure issues. TTC stated that trouble could be most likely related to a flex duct being blown off a termination as this is what we have found to cause this type of flow issues in the past.

04/15/05 – TTC stopped at the building on Friday to follow up with overheating of the 3rd floor space. TE tested the airflow into the general spaces and found that almost all areas in questions had too little airflow into the spaces. Some area had as much as 50% less or 100 cfm low airflow. The chilled water temperature and the cold deck air temperature are both right on design numbers – this appears to be an airflow issue at the actual boxes. TTC called A&M and found that they would be willing to give the County HVAC Pro for calibrating the boxes and troubleshooting the system. The serial cable to make the connection cost extra. MJ and TTC talked and MJ will follow up with A&M about getting the software and receiving training in it for this troubleshooting type work. We will start by checking the t-stat operation. The Mitsubishi units will help control this space nicely and the county trades are in the midst of installing one in the CODIS room right now. The electrical permit was approved on 03/21/05 for this work.

03/07/05 – TTC sent an email to A&M asking them to adjust the gain on the boilers for quicker turn down from high fire and to label the DX-9100.

03/03/05 – The County has let ThermalTech balance the pump volume to flows that TE believe will balance the system such that the primary water flow will always be greater than the secondary water flow or in such close proximity that no secondary bypass of water will occur. Overnight, Boiler #1 shutdown on a local controller (either water or high limit) and the plant kept the building warm with the backup boiler all night long. The setpoint varied from its reset value by as much as 15 degrees but the secondary boiler would come on and heat the building up, hence all the floor temperatures were fine this morning. TTC found that Boiler #1 was not running at 8:30 a.m. even though the control system had the boiler commanded on and targeted for 100% firing rate. The problem is that this nuisance trip alarm on the boiler is not reporting as an actual boiler alarm where the horn goes off and this makes it impossible for the DDC system to know that the boiler is in alarm so it must rely on the lag boiler to keep the building warm (this control worked perfectly as demonstrated by the history reports on the DDC system). Changing the flows seems to have worked perfectly as the leaving temperature of the boiler and the hot water secondary temperature have been within 1 or 2 degrees of each other consistently throughout the test. There are still some issues that have to be addressed and I am on it: 1.) The firing rate on the boiler controls seems to overshoot the setpoint to far and to long before backing off the firing rate and this could easily contribute to the boilers shutting down on a local high limit. 2.) The operating limit on the boilers doesn’t seem to shut Boiler #1 down before the High Limit – this forces a manual reset of the boiler as we saw last night. 3.) The actual pump flow for balancing is not know because the new balance points are too far toward the end of the curve and TTC will verify with actual manufacturer what the new flow are – we know the pressure drop at this point. 4.) The pressure gauge ball valves were all left open on the pumps – this causes water to bypass the pumps and decreases he effectiveness of the pumping system. TTC closed these valves and demonstrated that this bypass represented about 10 ft of hd on the secondary pumps and about 3 ft of hd on the primary pumps – these may have been left open when I flow tested the system the first time. 5.) TTC called the boiler manufacturer to discuss alarming points and what and how the low water alarm and the high limit alarm should operate, in conjunction with that, I need to confirm that A&M takes modulation control of the boiler and either the operating limit is in series with the high limit and they still act as safeties or only the high limit is acting as a safety for boiler water temperature. 6.) I believe we discovered that these boilers have a low limit controller that keeps the boiler on low fire when the vessel temperature is below X setpoint (i have to verify with Lynn from Bryan). When the temperature in the system was allowed to go down to 80 degrees last night during testing, we showed that the boiler when started locked into low fire until we took manual control and heated up the system (by the way the local and manual control worked perfectly last night for us and that was a change we had A&M make a couple of month ago when we discovered a power issue between Metasys and DDC). I will work these issues out but at first glance, the changes are working and the results are predictable. I may still have the check valve put into the system once I am convinced the system is operating and working properly.

03/01/05 – TTC has reviewed and calculated various system flows at the Coroner’s Office and presented to Gary Davies. I have targeted and shown reverse flow characteristics as shown by Mike’s and Roger’s field testing. I have found several fixes that can be implemented that require mechanical work in the form of adding a check valve in the bypass location or adding a blind flange in the bypass to prevent reverse flow. Neither of these address the unmatched flow characteristics between the primary and secondary pumps that is necessary to address since the current three way control valve position has all waters diverted into the primary loop under all conditions. TTC will present this possible solution to MJ at the Coroners Office in the next couple of days. With the cold weather we may be able to immediately test the solution under full load conditions.

03/01/05 – MJ contacted TE to discuss boiler rotation appears to automatically rotate. BLR-1, HWP-1 and SHWP-4 are on as of 03/01. We will wait to see what happens on 04/01. A low water alarm does not register on the front end when it happens at this time. We cannot tell from the as-builts how the DDC is connected to the boiler alarm sequence or contact point inside the boiler control panel. Humidifier #3 Burner #3 is locked out and TE contact RCME about this being right or if we were waiting on a couple of parts to fix that. TE stated that as of 09/07/04 we thought the humidifiers were 100% up and running and logged no missing parts features. Steve Garr of RCME confirmed later in the day that indeed everything was 100% working when they left and no additional parts were on order. RCME will assist in discovering why burner #3 is locked out.

02/22/05 – In an ongoing effort to never completely finish a project, we are meeting at the site today to prove that a check valve must be installed in the hot water decoupler between the primary loop and the secondary loop. TTC has already reviewed the P&ID and flows with Gary Davies and there does indeed appear to be a need for a check valve.

11/03/04 – Final Boiler Sequence submitted by A&M. 11×17 dwgs are to follow for O&M Manuals. PDF

10/14/04 – A&M visited the site today and rewired the boiler control panels to operate exactly as TE and Mike Jackson proposed. The Metasys/Local switch will allow the owner to switch between DDC control and local control. Under local control the boiler will either: A) if the boiler panel toggle switch is in “auto”, operate from the local boiler controls (penn controllers) mounted on the boiler shell or B) if the boiler panel toggle switch is in the “hand or local” position, operate directly from the potentiometer mounted on the boiler. The electrical 120V is a result of the combustion air dampers being interlocked to the boiler firing circuit ( TE is proposing to either put a placard indicated where else the power must be shut down or adding a relay that when the boiler main power is off a relay also shuts off this power supply too). A&M remapped the HWS value so that it is correctly displayed on the DDC graphics package.

10/12/04 – A&M has requested to meet Mike Jackson on site to go over his issues with the Boiler plant programming, 120 Volt inside boiler control panel and being able to run the boilers in local control.

09/07/04 – All parties can meet on 09/09/04 at 2:00 pm for startup.

09/01/04 – Geiler onsite reworking humidifier piping. TE to schedule RCME startup.

09/01/04 – Tony has approved state contract. TE working with JCI to provide all documents.

JCI has submitted a revised cost estimate to remove and rework the chiller plant. The new cost is $45,235.00. PDF

Geiler must rework the humidifier installation to separate the intake manifolds.

Automation & Mechanical is supposed to have new boiler programming installed by August 27, 2004.

County must decide to accept/reject JCI new price quote for the chillers.

by TEI TEI

Coroners – Chiller Plant DDC Upgrade

12/15/05 – Tony reported today that certified payroll has been received. TTC inquired about final payment being sent in that case.

12/14/05 – TTC emailed Tammy back to inform here that Certified Payroll still has to be received by the County before payment can be made.

11/15/05 – Tammy emailed us today to ask about final payment. Tony confirmed that we have not received certified payroll reports on this job. He needs those to process the final payapp. TTC contacted Tammy and asked for the payroll reports.

10/18/05 – TTC copied and will hand deliver tomorrow the final payapp and paperwork for this project.

05/27/05 – TTC received approval from Owner on completion of punchlist items at Coroner’s Office. TE sent 4 sets of as-built drawings and 2 sets of D size as-built dwgs to Tony Matre today via UPS. 4 sets of Owner’s Manuals were also sent with this mailing. TE will keep one copy of the O&M for its purposes at TE. TE will approve final pay application when it arrives from JCI. JCI submitted their Warranty letter to include as well PDF.

05/25/05 – TE received 2 “D” size set of control drawings per the County’s request. TE will deliver to the County. JCI provided a signed punchlist with the package. TTC approved some of the items. There is a remaining allowance of $5000 that needs to be used or returned before closeout on this project. The County needs to tell TE if the punchlist items are complete. Final paperwork should be sent in very shortly.

05/19/05 – TTC, JCI and the County met at the building and TE released the punchlist for the project. The commissioning phase of this project went extremely well and the chiller plant reacted and passed every test given to it. The County asked that the isolation valve be closed when the chiller is in alarm and JCI did that work on the fly. TTC asked for an OA setpoint where the County could tell the pump what speed to index to 100% at and JCI will add this trough graphics. Other than that the job was pretty much completely accepted. All timers, temperatures, start/stops and alarms functioned perfectly. Here is the punchlist PDF

04/18/05 – JCI was onsite finalizing programming on the DX9100 and wiring in the glycol pump and the chillers. TE arranged for Trane to be onsite at no additional cost to the County, JCI or TE and they worked out with JCI termination points to the chiller plant from the Metasys system. JCI pointed out that the graphics are not working yet but the text points are in place on the DDC computer in Ed’s office. JCI also showed Dave Maier how to enable the chiller and not have to bypass the relay in the penthouse. JCI will be back on Thursday April 21 to finalize the design.

03/24/05 – JCI has new DX9100 installed, powered up and programmed. JCI has verified all points to drive and is finishing point mapping so you can manually start pumps & valves from DDC if needed. JCI is completing hard wire commissioning check out of terminations for as-built sheets. Chiller would still have to be started manually also if cooling was needed at this time. JCI still awaiting wiring diagrams from Trane to wire directly into the Chiller. A Trane technician may be needed to coordinate the new input signals from JCI in the RTAA Chillers and Ed is calling Don Webber of Trane to assist with this work.

03/08/05 – Mike called today to say the Ed was looking for some Trane Control drawings. Mike Denton helped out with telling Ed that they are on the far unit and also with a follow up fax. Mike prefers these questions go through Trane or TE. I directed Ed to go to Trane with further similar questions. Chillers on manual control for next 2 weeks was also agreed upon. MJ stated that Trane Tracer hardware could be disposed of. TTC contacted Trane for pickup in the hopes that we can use this “nice” gesture on a future project with the Trane office.

03/07/05 – TTC visited the site on Monday and held the second construction meeting. EZ, TTC and Ed Braemer were present for the meeting. Much of the control wiring is pulled into place. TTC authorized the Trane Trace Panel to be removed and stored for turn over to Mike Jackson. I believe the power supply on the existing computer box is burnt up so when the computer is also turned over, remember that someone will have to replace the power supply before the OWS is completely functional if you ever plan on using the Trane Tracer system and computer. The DDC boxes should work fine. The automatic control of the chiller plant will now be down for approximately two weeks

03/01/05 – Mike Jackson reported the Automated Power is on site running cable. TTC contacted Ed Braemer and setup a construction mtg for March 7 at 8:30 a.m. for TTC, MJ, EB and EZ to kickoff project locally and discuss removing the Trane Controllers next week. The Chiller Plant will be down during this upgrade except possibly on manual control if needed.

02/10/05 – TTC sent the revised sequences to Ed Braemer for us on this project.

02/01/05 – After reviewing with EB on 01/25, TTC marked the prints up some more and revised the sequence. They will be delivered on 02/02.

01/25/05 – TTC has marked up the DDC control print for JCI. They are to be hand delivered today to Ed Braemer at 2:00 p.m. at the Admin Bldg.

01/21/05 – TTC inquiry regarding status. Still on hold and unable to speak to a person at the telephone listed on Johnson Controls Invoice. TTC gives direct telephone numbers to Alan Krupa and Tammy Blaeske. Left voice message for Tammy explaining that the invoice information should be on AIA documents G702 and G703 with a Affidavit of Original or Subcontractor form accompanying it. Called Alan Krupa and spoke to him in regards to the forms required. Tammy calls back and says that she has the AIA forms G702 and G703 completed but did not know what the Affidavit of Original or Subcontractor form is. TEI to mail the Affidavit for to Tammy, Tammy to hold the AIA forms for the Affidavit and send all three documents to ThermalTech Engineering.

01/03/05 – Called Johnson Controls and left voice message for Alan Krupa explaining that a notarized AIA G702 is required with a AIA G703 (Continuation Sheet) and a Form 44 Affidavit of Original or Subcontractor form instead of the Invoice form supplied.

11/29/04 – Off of the Budget Dept. County Engineer stamped on invoice. (Invoice mailed and received at the wrong place).

11/05/04 – Pre Construction Meeting

09/21/04 – SOSS scheduled to be delivered.

09/20/04 – Tony and Diana requested an SOSS for this project today. TE completed the state term contract for JCI and has already sent it down.

by TEI TEI

Comm Center – Fire Alarm Upgrade

10/05/04 – TEI, Greg and Dan from HCDF meet with Security Fence to review proposal. It was agreed upon that we where going to reuse the existing closer that HCDF installed and provide new hinges. This is less that work that originally quoted. Security fence is to revise quote and get to TEI for approval from Diana.

09/30/04 – Final test approved by fire marshal and county building inspector. Waiting for final close out documents from T.J. Williams to close project out. Simplex to provide training to Greg and other personnel. TEI to send as-built drawings with final closeout documents to Diana. We are nearing completion. Contractors preparing for pre-testing with simplex. Final testing approximately in two weeks around.

09/28/04 – Pretest complete and everything tested correctly. As Built drawings complete and delivered to site. Final test with the County Inspectors and Fire Marshal scheduled for 9/30/04 at 9:00am.

09/3004 – Fire Alarm Test passed.

10/19/04 – TEI left a message with TJ Williams for Rick and spoke to the person doing billing to request final billing documents. She was familiar with HC AIA documents and I explained the ones needed. She indicated she would get with Rick and send the correct documents. I spoke with them over a week ago, but I have not received anything yet.

The issue with the man gate, when we meet Charles from Security Fence he was going to get you a new quote that is corrected to reflect the work they are proposing to do, which is less than original. I have not received anything from him yet. I plan to follow up this week.

10/21/04 – TEI to follow up with T.J. Williams for training. As builts were forwarded from TEI to HCFD for archive room. TEI spoke with T.J. Williams and assisting in final documents.

by TEI TEI

Communication Center – Electric Building and Boiler Plant Upgrade

09/02/06 – TEI has paid Triton over $6000 in boiler cost and now another $4000 is going to them to cover the boiler snafu at the Comm Center. TTC advised Grady to finish up the paperwork and they will provide the Change Order for the project for completion. TTC asked for completion this week.

04/26/06 – TTC and Kevin approved the $850 allowance usage for Triton for stronger panels in the old louver spaces. TTC requested all O&M’s and final paperwork back from the contractors via email. Triton said the painting was complete for both the prime coat and supposedly the final coat two as Grady said is should be done last week.

04/04/06 – TTC emailed Kevin back with this information: gary and I already sent an FINAL email to AO Smith Monday morning and gave them a deadline of tomorrow at 4pm (wed) to reverse the cost on this project or else be out of spec’s. we will know something on thursday. I emailed grady today because, he, majid and I were working on a solution that could possibly remove all extra cost to the county in regards to the boiler work – don’t know if he made it work yet. finally, I have asked triton to paint the gas line and brace the aluminum pans in the boiler room to complete their punchlist. dave meier had a control question about a specific wire which I asked spectrum to give him a call directly. talk to you soon. ttc

03/06/06 – TTC and Mike talked about the building being in cooling today even though the OA temperature was 39F. Nevertheless the personnel had the building in cooling. We will keep an eye on the building. Kevin requested the price breakdown on the boilers. TTC will fax to Kevin today.

03/01/06 – TTC & Matt reprogrammed the boilers to run on the new differentials given to me by Lochinvar. The way the control algorithms are written we are using a OA setpoint with 10 degree offset above SP a 10 degree differential back down to SP and then the boiler should restart. To prevent nuisance trips I have reset the OA reset schedule to 0=160 and 60=110 with the 10 degree F.

02/24/06 – Mike reported that the boiler timing circuit was set up to 2 minutes because the valves take 1:40 seconds to stroke completely. On the reverse side the chiller is going out on freezestat because the minimal water flow on the slow opening diverter valves is creating too cold of inlet water temperature. JB Schmitt called today to report that their change order was $100 off.

02/22/06 – The boilers are going out on high limit safety so TTC visited the site and sent this email: Grady, I met Matt and Mike Jackson out at the site yesterday and am actively trying to resolve these issues with the boiler plant.

* after we added the common HWS temperature probe the boilers can run in lead/lag as we wanted but the probe can be as much at 18 degrees off what the water temp really is. this forces the boilers to high fire too long and they go out on a manual reset for high limit. I believe we will need the common HWS temperature put into a thermowell into the water stream to get a more accurate supply water temperature quicker to the boilers. the strap on location is okay it just needs to be immersed. can this be done on this 3″ pipe? * the boilers are locking out at 190 degrees hot water temp and this seems low even though it is their default. I have a two fold question about this and they are, is the factory setting 180 degree with a 10 degree deadband or is the autoreset actually 190 degrees with 0 degree setpoint? I am asking that this be reprogrammed up to 200 degrees F as this still provides plenty of safety before we steam at 212 and will allow us to reach our max system temp of 180. * how come the boilers can’t operate in lead/lag fashion when the common HWS temperature is not fed to the microprocessor? as soon as we disconnect the common the boiler both come on and fire at the same firing rate – is this field convertible? if so, we may be able to forego the above two steps and still meet my sequence requirements on this job but I think matt said that the factory said that we have to have a common T sensor point into the boiler to correctly do the lead/lag. * the outdoor sensor is still 10 degrees off so I instructed matt to move the new sensor over the old existing OA location in the area well and re-use existing wiring that the owner had in place to refeed the OA signal to the boielr panel. it is now near the bsmt ahu inlet and should be very well shaded all the time. please check with denny and the factory and let me know.

for now we are keeping both boilers running on the local outlet temp to prevent nuisance tripping but they are running in tandem all the time. no real lead/lag features even though they back each other up.

02/17/06 – since matt moved the hot water supply temperature we are seeing some funky conditions at the comm center as we discussed previously today by phone. mike jackson did a lot of troubleshooting for me today at the site so I want to relate everything he saw there and I am sure that another trip out there is warranted by matt. the boilers were off on flow when mike first checked them. he reset them manually as is req’d and set the delay timers up to 1 minute to allow the 3-way valves to stroke. in about 10 minutes the building temp was rising up to a reasonable rate but the boilers went out on high limit with an error message that says “Auto Reset – High Limit” our first thought is that this is certainly related to relocated the common HWS temperature and that the boilers were overfiring and going out on local high limit while the loop temp remained below its SP. we also checked actual pressure at the pumps in case the flow at the boiler is lower than is necessary. the outlet psig is 30 so that looks fine. boiler #1 appears to still be in the lead position and during the day mike continued to check on the boilers and reported this condition: at 40 degree OA the HWS setpoint should be 140 degrees. the boiler plant brought the boilers on at 139 degree F and the boilers again went out on high limit with the outlet temperatures at 190 degree F on the inlet at 158 degree F. at this condition it is understandable that the high should engage but why are the boilers still even firing at 18 degrees above setpoint? the LED panel has an point reporting as System Temp and it was 146 degrees. is this the reading at the strap-on sensor or is it the setpoint we are shooting for based on the OA temp? if it is the setpoint it is off by 6 or 8 degrees and I would question whether we are using 0-180 to 60-120 or the old reset of 0-180 to 60-140? can matt verify this for us. if having the remote sensor on the common supply pipe is going to cause this much trouble, I may just go back to using the outlet of each boiler to control its local firing rate and get rid of all these nusicance trips we saw today, but I tend to think that maybe a parameter is currently misprogrammed. as you suggested we should certainly check the flow through each boiler to make sure we are removing the correct amount of heat from the boiler while it is firing and the water supply is sufficient. without any real water sitting in the heat exchange these boilers will always tend to heat up fast on the outlet while the system loop lags behind. my final request was to isolate boiler #2 and turn off boiler #2 so that #1 handles the load all weekend. with all the flow going through one boiler there should be no mixing and the outlet of the boiler should be the same as the strap-on sensor and we shouldn’t have any problems this weekend. can you schedule matt to meet me out there with jackson on tuesday to go over all of this? I have to take my father in law to the hospital on monday and my wife to the dentist for dental surgery so that is going to be sucky day for me all the way around. please set up tuesday at 9am at the comm center

02/15/06 – TEI, ICS, Triton, Facilities and PW met on site with Bldg Dept and Fire Dept for final inspection and testing. After an initial roadblock, Buckeye Power to come over and simulated a electric failure in which the building passed. The Colerain Fire Dept contact, Jim witnessed and will report to the County Bldg Dept. The generators came on in under 2 seconds and restored power to the building in under 10 seconds. TEI also verified the operation of fuel oil pumps on both the supply and return system for the generators. The boilers were tested and they backed each other up properly. The electric boiler provided manual backup of the gas boilers. The pumps provided backup for each other within 20 seconds of failure. TEI has instructed Triton to move three temperature sensors. The OA is going to be moved to the area well and they can reuse existing wiring for that connection. The HWS temperature will be moved inside the new boiler room to after the electric boiler location in the common HWS header. The chilled water temperature strap-on should be moved to the common outlet of the 3-way valve connection (currently it is about 10 degrees off so we set the start temperature of the chiller to 80 instead of 80). We will allow facilities 2 weeks of working with the boilers before we have final training on operating the Lochinvar’s. The boilers are scheduled to rotate daily for lead at this time. The master controller is located in side Boiler #1 and if for any reason the power to this boiler is turned off or failed then Boiler #2 would have to be removed from cascading control to operate in Boiler #1’s place. The gas line still needs to be painted. The fuel oil pumps always switch lead after about 10 seconds and Triton will fix this. The County would like to have Triton put reinforcing bars put behind the louvered openings for security and stability. Facilities will have the smoke head remounted in the new boiler room. TEI want like JC Schmitt to do a final cleanup around the building as TTC noted a lot of debris still laying in the field (metal rods, sheet metal pieces, wood, studs, canvas, plastic containers, etc). Triton should update the sequence to reflect backup and alarm sequences and submit to County along with final boiler programming settings and include this information in the O&M. Krebs came out after the testing was complete but did not have to do any work as TTC met them at the gate and told them we were complete. We may have to pay a service call for that arrival.

02/10/06 – IBI gave final approval to the building electrical and now we are ready for final mechanical and building inspections.

01/23/06 – Lochinvar provided the final programming instructions for the boilers. Triton has every intention of completing the punchlist and programming this week. Jeff Jones is supposed to provide the UL listing on the fuel oil pumps. We are shooting to have all final work and inspections completed this week on this project. TTC will try to coordinate final inspections this week. JB Schmitt entered a allowance usage request of: cut out scuppers: $42 + $191.84 remove concrete piers: $50 + $40 + $130 +$191.84 lock conversion: $227 + $335.72, total: $1329.24. TTC forwarded to KM, TA and TS for approval.

01/20/06 – John Bates emailed TTC today to report that Sean has looked at the boilers and AO Smith should have an answer early next week.

01/17/06 – TTC emailed ICS to inquire about electrical status on this project. ICS reported that the fire alarm test passed. Vic Spang will email everyone the status of the UL listing. Vic also said that O&M manuals should arrive today. ICS should be done with our punchlist items about 2 days after the control panels are UL listed.

01/16/06 – TTC spoke to John Bates last Friday and emailed our these notes today. TTC talked to Grady and got this update on the punchlist status. Triton has gotten the collars, valve tags and miscellaneous punchlist items into their shop and will continue to finish the punchlist this week. They will clean up the generator room this week as well and fix the two leaks that Greg called TTC about. The control programming codes are expected in this week from Lochinvar and Triton will finalize all control work this week and next at the Comm Center. Final completion date should be no later than 01/27/06. TTC will contact ICS to coordinate the final building inspection on this project.

Kevin and Grady,

John Bates and I finally connected via the telephone late last Friday afternoon. It was a good conversation and he apologized to everyone for it taking so long to get back to us on this issue. I recapped everything that has happened on this project beginning in March of 2005 when we were finishing the construction documents up to the partial installation. Suffice to say that he said he believes what I have told him is the order of events and the phone calls I made to the local rep and the factory and that although it is not the AO Smith normal policy to take back opened boilers, he will begin the return authorization process by sending an AO Smith representative to Hughes supply to confirm the condition of the existing boilers and take some digital photographs. The restocking fee will have to be paid but I believe he believes he will be able to get the factory to take the boilers back.

I have called Gary Tedesco of Hughes and advised him that Shawn from AO Smith will be stopping by to look at the boilers. As I get further information I will pass it along to you guys.

Right now the boilers are being stored at Hughes Supply.

01/13/06 – TTC called and emailed John Bates again about returning the boilers.

John,

I have called a couple of more times yesterday and am on the phone right now trying to get a hold of you and leaving you another message. This issue has been lingering since the first week of December. I understand that you must be very very busy but I have scheduled meetings with the contractor to discuss change orders and closeout for this whole project – the job is complete and we still have you boilers on site, so you can see I have a dilemma – we can’t closeout this project and I believe that Mr. Maue is desperately trying to get this project wrapped up.

This is only my second time working with Mr. Maue of Hamilton County Public Works and I have no doubt that he has serious concerns about my ability to handle the construction issues on a project like this so I desperately need your help. I don’t know that I have ever had something stall for so long or been so worried that there may not be an equitable resolution on the horizon – I am still confident that if you call me back we can work something out.

As I have specified AO Smith boilers for so long – 13 years now (I personally enjoy using the cyclone water heaters as they are a fantastic product – I spec’d them as part of an energy savings project, we use them in all the hotels we design and that’s about 5 a year and the County also used them at a recent renovation at their sheriff patrol hdqtrs), I am finding it difficult to understand our current dilemma. Your former local rep built a great relationship with ThermalTech and your company has always been foremost in our specs and schedules – I am sure that neither of us wants that to change, nor does it have to, please call me back or email me the return authorization and procedure.

If this is a decision that you cannot make, please send me the contact information of you supervisor or someone in customer service who can make the authorization.

Thanks again and hope you have a great weekend and please let me know by the end of business today?

01/12/06 – Kevin emailed Grady directly to inquire about the progress of this project. TTC informed him that the email server may be down.

01/11/06 – Greg Chambers called TTC confirmed today’s 1:00 p.m. test is only a Fire Alarm test. TTC called John Bates two more times this morning about returning the Legend 2000 boilers. TTC called back at 3:00 and left another message. TTC got John’s email address and emailed him the same request that he has been calling about:

Dear Mr. Bates,

We have missed each other in phone tag and I know you are on the road a lot for your job. I am hoping it will be easier to give me a RMA process and number via this email.

In March 2005, I spoke with your local rep Disney-McClain and called your Corporate Sales department in search of boiler that we could use as base bid in our Dual Temperature Water system at the Hamilton County Communication Center, my first inclination was to use the intelli-fin boiler but you don’t manufacture those in a 500,000 btuh size so after speaking to sales staff at your companies and vendors, we spec’d the Legend 2000 with the bypass pump option to keep the incoming water temperature at 140 inlet.

We bid the project, accepted the submittals and received two boilers on site. They were uncrated, placed on the housekeeping pads and one boiler was actually piped in. Electrically, a junction box was mounted on the side of the boiler for controls. At that point I was called down to discuss the bypass pump. In the arrangement and in your documentation the pump was piped in series with the water flow and a bypass pipe around the boiler was on the outside. This is not what I had envisioned as I was lead to believe that he bypass pump would have self contained controls that would modulate in the bypass to return enough water to keep the inlet at 140. In actuality that is not the case and after speaking with Kevin Binkley at length, I could see these boilers do not meet the spec or requirement of this job. The alternate boilers on this project are all condensing boilers and would not have an issue with incoming water temperature. This dual temperature water system can be switched from chiller to boiler several times a week so you can see the Legend 2000 boilers will not work well without a storage tank and recirc pump which is exactly what I was trying to stay away from on this project.

At this point the boilers HAVE NEVER BEEN TURNED ON or used in any way. What I would like is a return authorization so that the local contractor (Triton Services) can return the boilers to local warehouse (Hughes Supply) and they can send them back to AO Smith.

There may very well be a restocking fee that I am willing to work with you on that but in my humble opinion, I believe I talked to everyone that I needed too during the design phase of this project to let AO Smith base bid on this project – I just got some bad information on application data for this boiler. Maybe you guys thought the dual temperature water system meant it would only be switched to heating once a year and that the boiler heat exchanger could handle that??? I never said that, but there was definitely a communication error.

I am sure that these boilers can be resold to someone – they are in like near perfect shape and never started even though they have been uncrated. They have been stored indoors even after we removed them from the boiler room. Please give me an return authorization approval, an RMA number, an address and a procedure to return these boilers.

The County has used many AO Smith products in the past and when they build their new jail complex in the next few years, many more AO Smith could be used in that project so I am sure we are all striving for the same goal here.

Please call me or email as needed.

Thanks again, ttc

01/10/06 – TA distributed the electrical punchlist for the job. TTC distributed a revised mechanical punchlist for the job. Triton sent payapp #3 in for pencil copy and question payment of payapp #2. Kevin submitted payapp #2 for payment per Grady’s inquiry.

1/09/06 – Kevin called today because he is concerned about Jeff Jones UL listing of the generator. TA talked to ICS to tell them to get something resolved or the County will hire Jeff Jones directly and back charge the contractor. TTC talked to Triton about controls and emailed everyone that the controls are semi-automated. The new actuator is installed and the summer/winter switch now strokes the valves just like they use to work. The system is currently at a state where it was before. The Comm Center can manually change from heating/cooling and the system automatically adjust and powers up the correct equipment. The new lead/lag controls are not installed yet nor is the OA reset on the boilers as they are set to 160 F until next week where the controls will be finalized. Triton still has to change the wiring on the fuel oil pump package to get the UL label.

01/04/06 – Triton called to report that the boiler circuit boards appear to be out not because of water issues but voltage issues. the transformers inside the boilers are bulged and burned, the DC power inverter on the circuit board was blown. Buckeye reported that there have been a couple of brown out issues and once occurred for one second on 12/26 and the generators were tested on 12/30 but they don’t see how this could be related to the boiler problem. No other components have failed in the building on any computer circuits if there was some sort of massive power spike. Matt from Triton talked about putting the new circuit boards on the boiler and monitoring the power into the boilers to ensure they are getting clean power. The IBI inspector, Earl Moon, requested to see UL labels on the fuel oil pumps, tank controllers and the electric boiler. TE investigated and determined that the electric boiler is UL listed and Triton needs to show the inspector where that label is, the fuel oil pumps will require a 3rd party certification from Jeff Jones.

12/28/05 – The mechanical punchlist was distributed. PDF. The electrical punchlist was distributed last week via a verbal conversation from Terry Allgeyer. The final building inspection and testing with the Fire Department is scheduled for Wednesday, January 4th at the site.

12/08/05 – The electric changeover went very well. TA attended for ThermalTech. The building is on new power. Training was completed on the same day for the people that stayed throughout the work and training day.

12/06/05 – TTC sent out CM notes #11. A one week meeting is scheduled for 12/07 to coordinate final changeover of building power to new system. Training is scheduled for 10am on Thursday per TA’s email. Blr-2 is on site. Triton has installed the louvers for the new generators but a final plenum piece is required to finish up. TTC discussed progress with Grady and questioned how the hot water control would be handled. TTC wants a common HWS point to feed the loop temperature to the boilers.

11/28/05 – ICS and Buckeye are onsite preparing final connections to the generators. Triton has finished the muffler install and are working on the final louver install. Kohler is scheduled for next Monday. TTC sent out const mtg #10 meeting minutes. Const Mtg #11 is set for Wed at 2:00 p.m. TTC and TE are working on boiler scheme at 237 WHT. The 3-way valve has been installed and is up and running. Triton provided a cost to the project of Material $162.40 Labor $1,564.00 New valve $1,131.00. This may be slightly higher if charged to the project.

11/23/05 – TTC & TA coordinated with ICS, Triton and Kevin about generator start schedule. It appears that Kohler will be onsite on 11/28 & 11/29 to finalize installation of generators and pretest before Kohler come out to the site. TTC will try to reset the meeting for next weekend.

11/17/05 – TTC called Grady today to confirm this evenings work with Triton and the 3-way valve changeout this evening. Pat and Matt from Triton will be onsite this evening about 11:30 p.m. for this work. Mike Jackson met with Pat from Triton today and went over which breakers will be shutdown tonight to turn the AHU’s off for this work. Matt will take care of the controls for this work this evening and if anything doesn’t work the system will fail into the heating mode for now through manual settings. TTC called the AO Smith Regional Manager, John Bates at 770-841-1305 (off) 770-521-4244 (cell) to discuss the return of these boilers to the factory, I had to leave a message on his voicemail. TTC may stop in tonight to review the work in progress.

11/08/05 – TTC called Hughes supply today trying to get authorization to return the unused water heaters. We still have not been able to talk to the regional representative.

11/03/05 – TTC has been talking to Triton and Hughes supply locally to get authorization to return the boilers to the factory and work on the restocking fee. Gary Tedesco at Hughes has been putting in calls to the regional sales person and the new rep in town LIBB. To date we have had no success in talking to anyone of consequence to help up return these boilers. We haven’t even received authorization to return the boilers to the factory yet, let alone handled the restocking fee and this is mostly due to the fact that A.O. Smith has severed representations with Disney-McClane locally. We probably won’t talk to AO Smith until Friday when they are out of conferences.

11/01/05 – The generators were delivered to the site on Monday. They are in place. Cable is pulled but not terminated. Cinergy will provide final inspection later this week. Power may be up and running by 11/14/05. TTC to send out mtg notes today or tomorrow.

10/27/05 – After the construction meeting yesterday, TTC and Triton went over the boiler workings, final hookups, controls and flow settings. We decided to start the boiler today and as of 2:00 p.m. the boiler is running perfectly and producing 140.5 degree hot water at 51 degrees F. TTC will try to stop out on Friday to check boiler over before weekend. Triton is going to verify that boiler turns off when the switch is changed from summer to winter position automatically. We want to verify this to ensure the boiler turns off whenever there is not flow into the boiler heat exchanger.

10/25/05 – The new boiler has been shipped. It has arrived in Fairfield at the FedEx terminal. It should be delivered this afternoon the Comm Center. Triton will be there to take delivery of the boiler and begin the install of the new boiler. TTC will help coordinate the install as needed.

10/24/05 – TTC emailed the whole team today reporting the following about the revised boiler design.

I want to document all that has happened on the boiler plant since Friday in this email.

I have decided that there is no “best” way to use the AO Smith non-condensing boilers at the Comm Center. My discussions with the factory, who have been very responsive, still have not offered a solution that is easy to implement or easy to maintain. The controls would be cumbersome and the energy usage would go way up since the boiler plant would have to run all year long once the heating season started. Short cycling could not be avoided without the introduction of a storage tank which was never part of the project in the first place. I am not sure what the factory was thinking when they told me this boiler could work with its bypass pump back during the design phase of this project, but the final word from Kevin Binkley at AO Smith is that it cannot do what I am asking it to do in its current design.

It is my opinion that the BEST and the CORRECT way to design this dual temperature system is to use a fully condensing boiler just like Alternates #1 and #2 in the original bid and my original design. This fully condensing boiler is not subject THERMAL SHOCK like other boiler and the STAINLESS STEEL HTX is not subject to corrosion like a standard heat exchanger causing the premature failure. This is accomplished by providing a STAINLESS STEEL heat exchanger with the boiler as with the KBN-500 boiler by Lochinvar. The boiler that I am now having Triton provide for this job is a Lochinvar KBN-500 boiler, Knight Model at 90.3 efficiency with a fully modulating burner with 5:1 turn down ratio, Energy Star qualified, ASME stainless steel heat exchanger (12 year limited warranty) , direct vent with sealed combustion (meaning PVC vented), direct spark, 0-10 VDC input for burner control, alarm contacts on any failure, 4″ air inlet and exhaust vent and low water cutout.

Other design notes: These boilers are rated for 50 gpm max and our pumping system is sized for 80 gpm so as it was originally intended the boilers can work in unison to meet supply temperature or one boiler can handle the whole load if needed but the water flow should go through both boilers all the time unless you have to isolate one for maintenance and at that time, the bypass should be opened if the 2nd boiler is going to be offline for a considerable period to keep the internal flow from getting too high in any one boiler. We will manually open the bypass partially while we are in one boiler mode on this project – your heating loop will still get plenty of hot water to keep the building warm in this temporary arrangement.

For the time being, I am going to have the hot water loop maintain 180 degree F until all the controls are installed and ready for operation with both gas boilers and the electric boiler in the plant (the second gas boiler should arrive on or around November 1st), the electric boiler is already here but there is no 208 Volt power for it yet. Each of the new gas boiler are going to have a 3/4 tygon tube that will discharge a little condensate into the existing floor drain the boiler room whenever they are running. The condensate discharge from these boiler is safe to the plumbing system and should have no “ill” effects on the drainage of the building or the floor drain in the room.

The new control system will monitor the dual temperature water loop temperature so that when you switch from winter mode to summer mode (140 degree supply to cooling 45 degree supply) instead of waiting three days for the loop to dissipate heat, the boilers will be turned off but the heating hot water flow will continue through the boiler system and out the building through the air handlers discharging all remaining heat from the DTW loop into the space until the loop reaches 90 degrees. At that time, the 3-way valves will changeover and go to cooling. This is a strategy that should be used with 2-pipe systems that often isn’t because it will allow you to get to the cooling cycle must faster than usual. The boilers should never be “on” without flow going through the heat exchanges – that is requirement of the factory. Inlet water temperature and water PH is not an issue with the stainless steel HTX. These boilers have intelligent controls that do everything the original sequencer was supposed to do and they communicate directly with each other in case of failure in a master/slave arrangement.

And finally – ONE KBN-500 BOILER will be on site on Tuesday. It is being shipped directly to the Comm Center from Nashville, TN today for the project.

That’s it for now. I plan on being out there tomorrow after the boiler arrives to work with Triton on installation, but I do not plan on being there today as originally intended because we are not setting the electric boiler in place at this time. I am not calling AO Smith about the 3rd party UL licensing at this time either.

The temperatures are supposed to stay in the mid 50’s all week with low’s in the mid to high 30’s so I am working to get this boiler online tomorrow or Wednesday.

I will give you another status update later today or tomorrow as more information arrives at my desk.

Kevin, Mike Jackson would probably be very interested in having all this information, please forward it on to him if you deem it is good info to have.

10/18/05 – TTC has contracted Lochinvar for more information on the “Knight” series condensing boiler to take care of problems at the Comm Center Project. TTC will call Kevin Maue to discuss results of speaking with AO Smith about DTW issue and will explore possibility of installing electric boiler now in Blr #1 equipment pad to alleviate any boiler issues due to cold weather. This may be the only viable option if we switch boiler types.

10/14/05 – TTC is working closely with AO Smith to get the control of the boilers correct before they are started at the site. Condensation and control is a huge issue.

10/06/05 – Grady reported today that the gas meter pad is being poured on Friday, the electricians have power to boiler #01, the gas line should be hooked up and meter placed by Monday, October 10. There will most likely not be boiler operation until that time. Triton is meeting their controls company on site Thursday afternoon for temporary hookup of boiler #01 to existing summer/winter control switch. It will remain hooked up like this until the new controls are put in place the 3rd week of November. This basically follows the original scheme except that we wanted boiler #01 operational by 10/01 with 10/15 as the drop dead boiler startup date. ICS contacted blacktopping company to resolve the roadway hump from the horizontal boring. This will hopefully mitigate the problem we are having with the roadway bump.

10/05/05 – Const Mtg 07 has been distributed. TA discussed shaving the roadway down to reduce the hum but that is not an option. TTC and TS are still working on louver details. TTC and Grady are trying to get boiler #01 online by Friday or Monday at the latest.

10/04/05 – Motorola confirmed that we just need to provide a ground rod from each generator to grade. No additional grounding ring or grounding methods are required for this building. ThermalTech will advise the contractors and send out the meeting notes later today including this new information.

10/03/05 – Motorola gave us an answer about the building ground. The new building requires a grounding ring or a separate ground rod. TA is currently investigation this item for the County.

09/28/05 – Const Mtg 07 was held on site. Several major issues surfaced that need to be addressed by ThermalTech and the Owner. Motorola is going to require additional grounding on the building metal (TA to investigate), Boiler #1 must be online by October 15th and the sooner the better (Triton to provide a 2 week schedule to accomplish this), gas meter pad has to be poured and set in place and tested before boiler can start, backfill must be completed, boilers are seismically braced yet.

09/23/05 – HamCo was required to sign an IBI temporary electric meter release stating that electric work is ongoing at the site. Greg Wenz stated he would take care of this for IBI. ICS sent TE the letter and we passed it along to the Comm Center.

09/22/05 – This week we have been working on scheduling since the generators are very late. We have also been working on louver details with Tom Speer and Triton as they are changing heights and design. TTC faxed over information to TS, GR and KM for review and approval. Slabs have been poured on site and work is ongoing.

09/14/05 – Triton reported that the pumps will be arriving later in the job rather than earlier. The pump inertia bases will sit a new housekeeping pad that will encompass both of the pumps and may surround the existing pump pad. All new anchors will be into the new housekeeping pad. JB Schmitt turned in payapp #2, TTC will deliver to Kevin at Wednesday’s meeting.

09/01/05 – TTC, Mike Cavanaugh and the County are trying to coordinate sending the demolished chiller to New Orleans for the relief effort. TTC sent out Const Mtg 05 notes to all the team members and TA must talk to Buckeye Power about a late delivery schedule. Buckeye Power reported that generator will be here mid-October but Vic will talk to factory and get back to us mid next week.

08/29/05 – TTC stopped at the site. The foundations are poured. A box was left for fuel oil and electric. Apparently, the electric had an issue with IBI and they will need one more underground inspection before acceptance. JC Schmitt worked diligently to get foundation in before Tuesdays rainstorms. TTC approved payapp #01 for JC Schmitt and #01 for ICS. TE has not received a payapp from Triton services yet.

08/17/05 – Const Mtg 04 today. We discussed and update the schedule quite a bit. JB Schmitt can still meet the anticipated schedule based on today’s schedule modifications. The CM # 04 notes were sent out for review. Triton brought Ketchum & Walton out to the site to resolve the gas line mounting issues that ThermalTech noted on the submittals. It was agreed during the meeting to allow the gas piping to be routed on the backside of the building wall at the parapet level of the rooftop. Triton went on to say that won’t change the permit submissions and it will still met seismic requirements. The County preferred the gas line to stay routed above the rooftop as it was designed on the drawings. Kevin and Terry are to check with County Examiners on releasing part of the permit set next Tuesday. TA has to answer an RFI noted in the meeting minutes for conduit inside switchgear on Citadon. TTC will send out an updated schedule this week.

08/16/05 – Construction Meeting # 03 is finished. TE check on permit progress. TTC talked to Disney McLane about requirements for electric boiler and told them it is only for heating hot water and its working pressure is 60 psig and it should have an output reset schedule of 0-180 degrees F/60-140 degrees F. Triton turned in two RFI’s to Citadon that TE is trying to resolve. We are trying to get more information from Buckeye power for answers. TTC will send out meeting notes and remind contractors of the Wednesday meeting.

07/21/05 – Construction Mtg 02 held. Triton to make pump tie-ins overnights and allow Geiler to do chiller tie-ins as well. Seismic calcs are not ready. ICS provided a checklist for consultants to fill out for this job. TE said they would try to get it back on Thursday. TE will answer all the questions they can. TTC will type of meeting notes on Friday for distribution. Tom Speer reported that the county is going to required 6 stamped stets. TS will have Ohio Blue make sets and send to TE for stamping. TE will return to Tom for permit. Schedule was discussed and a mid August start date on the new buildings was agreed to without a schedule slip. Triton should begin excavating next week.

07/12/05 – Construction Mtg 01 distributed. Work has begun. Seismic Calcs are behind. Triton schedule is slipping. TTC to supervise and get Triton on board and up to speed. Electrical equipment has been ordered and should arrive in late August. TTC to coordinate valve installation at site soon. TTC emailed JB and Speer about missing GC submittals. TTC emailed Triton about getting back on schedule and tried to setup a field trip for Thursday or Friday of this week. TTC emailed Grady to setup a field trip for controls and site work.

06/0605 – Pre Construction Mtg notes distributed.

06/03/06 – Electrical submittals are in the works. General submittals are uploaded to Citadon.

05/26/05 – Pre Construction Meeting at Site.

04/22/05 – All contracts arrived by 3:00 p.m. TTC will review and have to Public Works by 5:00 p.m. hand delivered.

04/21/05 – Pre Award Meeting is complete. Bid Tabs are accepted. Contractors have contracts and they are due back to TE by Friday at Noon. TE to deliver to KM by EOB on Friday.

04/19/05 – TE setup for Pre-Award meeting on Wednesday. Allgeyer is pushing Citadon for the whole project.

04/06/05 – Cinergy doesn’t have bid numbers to the vendors yet hence causing a bid extension of one week to the project. Addendum 02 was distributed. PDF

04/01/05 – Addendum 01 was distributed for Project. PDF

03/23/05 – TTC and HamCo Facilities visited Glen-Este Middle school to inspect a Fulton Boiler application installation. Basically this boiler has been in service for 7 years with little or no maintenance at this site other than the changing the spark plug. Turton showed the boiler, explained that it was solid, that it pulsed, had very few moving parts, and could keep a facility at a solid temperature no problem. Facilities has heard that Fulton are maintenance headaches and that parts are not kept locally. Jim Turton of Ballenger told us that all those parts are kept locally in his office. TTC added that the statement that everything was proprietary is certainly not true because the local controllers are standard Honeywell and the gas modulation valve is standard Honeywell and the only things that TTC saw as special could be the fireye which is a Fulton model of a Honeywell controller and the gas pressure switches but basically everyone’s are proprietary anyways. Facilities called back to inquire about the Weil-McClain Ultra series boiler but TE stated that max btuh is 310,000 so that boiler is too small and more residential grade boiler. Lochinvar Intellifin has built in recirculation loop but their smallest boiler is 1,500,000 which is 3x too big but they have a 14:1 turndown ratio but using them would mean the boiler is going to spend most of its life in the 20%-60% firing rate. TTC called the AO Smith factory and the local rep Disney McClain and discussed other options. Apparently the Legend 2000 boiler series IF it is ordered with the bypass option can handle cold water inlet conditions.

03/22/05 – TTC authorized Ohio Blue to use AIA documents to insert into spec book and then send to all personnel, later TE delivered the electronic forms for xeroxing to Ohio Blue for future use. SBO delivered a CD for purchasing to upload the entire project to Demandstar. This was completed.

03/21/05 – Bid Set is at Ohio Blue. TE delivered in morning.

03/10/05 – 90% Review Set and Review Meeting at Comm Center. TTC needs to find 2 more Pulse type boilers for bidding. KM requested updated Cost Estimates, Comm Center agreed to one or two 8-hours shutdowns for new 3-way flanges and Comm Center agreed to a 4-hour switchover on rented generator to go to new power from new source and new generators.

01/26/05 – TE sent out DD dwgs for review in anticipation of this Friday’s Meeting with the client.

12/27/04 – TTC and Kevin Maue discussed in detail the savings associated with a non fuel oil boiler and chimney. Kevin is going to take the scheme to Facilities for approval and give TE direction by end of week. Kevin is wondering where the Schematic Design documents are and would like to see them soon.

12/24/04 – Schematic Design Due. TTC turned over cost estimate for revised boiler design using 2 gas boilers and 1 electric boiler.